### v1

# SPEED: Scalable, Precise, and Efficient Concept Erasure for Diffusion Models

Abstract
Erasing concepts from large-scale text-to-image (T2I) diffusion models has become increasingly crucial due to the growing concerns over copyright infringement, offensive content, and privacy violations. However, existing methods either require costly fine-tuning or degrade image quality for non-target concepts (i.e., prior) due to inherent optimization limitations. In this paper, we introduce SPEED, a model editing-based concept erasure approach that leverages null-space constraints for scalable, precise, and efficient erasure. Specifically, SPEED incorporates Influence-based Prior Filtering (IPF) to retain the most affected non-target concepts during erasing, Directed Prior Augmentation (DPA) to expand prior coverage while maintaining semantic consistency, and Invariant Equality Constraints (IEC) to regularize model editing by explicitly preserving key invariants during the T2I generation process. Extensive evaluations across multiple concept erasure tasks demonstrate that SPEED consistently outperforms existing methods in prior preservation while achieving efficient and high-fidelity concept erasure, successfully removing 100 concepts within just 5 seconds. Our code and models are available at: https://github.com/Ouxiang-Li/SPEED.
Figure 1: Three characteristics of our proposed method, SPEED. (a) Scalable: SPEED seamlessly scales from single-concept to large-scale multi-concept erasure (e.g., 100 celebrities) without additional design. (b) Precise: SPEED precisely removes the target concept (e.g., Snoopy) while preserving the semantic integrity for non-target concepts (e.g., Hello Kitty and SpongeBob). (c) Efficient: SPEED can immediately erase 100 concepts within 5 seconds, achieving a ×$\times 350$ speedup over the state-of-the-art (SOTA) method.
11footnotetext: Equal Contributions.22footnotetext: Corresponding author.
1 Introduction
Large-scale text-to-image (T2I) diffusion models [24, 56, 57, 39, 49, 23] have facilitated significant breakthroughs in generating highly realistic and contextually consistent images simply from textual descriptions [11, 46, 16, 7, 50, 44, 13]. Alongside these advancements, concerns have also been raised regarding copyright violations [10, 54, 30], offensive content [51, 65, 67], and privacy concerns [8, 64, 31]. To mitigate ethical and legal risks in generation, it is often necessary to prevent the model from generating certain concepts, a process termed concept erasure [29, 17, 66]. However, removing target concepts without carefully preserving the semantics of non-target concepts can introduce unintended artifacts, distortions, and degraded image quality  [17, 42, 51, 66], compromising the model’s reliability and usability. Therefore, beyond ensuring the effective removal of target concepts (i.e., erasure efficacy), concept erasure should also maintain the original semantics of non-target concepts (i.e., prior preservation).
In this context, recent methods seek to balance between erasure efficacy and prior preservation, broadly categorized into two paradigms: training-based [29, 37, 35] and editing-based [18, 19]. The training-based paradigm fine-tunes T2I diffusion models to achieve concept erasure, incorporating an additional regularization term in training for prior preservation. In contrast, the editing-based paradigm is training-free. It jointly optimizes erasure and preservation objectives for target and non-target concepts in a closed-form solution, obtaining direct modifications over model parameters (e.g., projection weights in cross-attention layers [49]). This efficiency also allows editing-based methods to seamlessly extend to multi-concept erasure without additional designs.
However, as the number of target concepts increases, current editing-based methods struggle to balance between erasure efficacy and prior preservation. It can be attributed to the growing conflicts among erasure and preservation objectives, making such trade-offs increasingly difficult. Moreover, these methods rely on weighted least squares optimization [18, 19], inherently imposing a non-zero lower bound on preservation error (Appx. B.2). In multi-concept settings, this accumulation of preservation errors distorts non-target knowledge, further degrading generative quality.
To address the above limitations, we propose Scalable, Precise, and Efficient Concept Erasure for Diffusion Models (SPEED) (see Fig. 1), an editing-based method incorporating null-space constraints. Here, we define “null-space” as the subspace that does not alter the feature representations of non-target concepts. By projecting the model parameter updates for erasing concepts onto the null space of non-target concepts, SPEED can minimize the preservation error close to zero without compromising erasure efficacy. A key challenge for SPEED lies in constructing an effective null space based on a set of non-target concepts (i.e., retain set). We observe that the existing baseline with null-space constraints [14] confronts a fundamental dilemma during concept erasure: While a narrow retain set limits the prior coverage, increasing the diversity of the retain set restricts the determination of the null space due to rank saturation, thereby compromising prior preservation (see Fig. 2). In this light, we introduce Prior Knowledge Refinement, a suite of techniques that alleviates optimization difficulty and broadens prior coverage.
Particularly, to mitigate optimization difficulty, we propose Influence-based Prior Filtering (IPF), which selectively retains non-target concepts from the retain set that are most affected by the erasure process. These concepts serve as anchors for defining the null space, providing meaningful constraints for preservation while maintaining sufficient flexibility for model editing. To further prevent excessive prior knowledge reduction, we introduce Directed Prior Augmentation (DPA) to enhance prior coverage, which expands the selected retain set after IPF with augmented samples. Instead of simply introducing random noise, we generate semantically consistent variations of preserved concepts, mitigating excessive preservation over meaningless concepts. Furthermore, we incorporate Invariants Equality Constraint (IEC), which identifies critical invariants in the generation process and enforces equality regularization to explicitly preserve them. We evaluate SPEED on three representative concept erasure tasks, i.e., few-concept, multi-concept, and implicit concept erasure, where it consistently exhibits superior prior preservation across all erasure tasks. Overall, our contributions are summarized as follows:
  • We propose SPEED, a scalable, precise, and efficient concept erasure method with null-space constrained model editing, capable of erasing 100 concepts in 5 seconds.
  • We introduce Prior Knowledge Refinement to construct a well-defined null space that remains meaningful and tractable for editing. By integrating Influence-based Prior Filtering, Directed Prior Augmentation, and Invariants Equality Constraint, it balances rank saturation and prior coverage effectively with precise concept erasure.
  • Our extensive experiments show that SPEED consistently outperforms existing methods in prior preservation across various erasure tasks with minimal computational costs.
Figure 2: Impact of the increasing number of non-target concepts in retain set. The baseline [14] maintains prior preservation with a small retain set (e.g., Refer to caption), but as the retain set grows, rank saturation occurs, leading to gradually degraded semantics in non-target generations (e.g., Refer to caption and Refer to caption). In contrast, our method achieves consistent prior preservation by optimizing the retain set.
2 Related Works
Concept erasure. Current T2I diffusion models inevitably involve unauthorized and NSFW (Not Safe For Work) generations due to the noisy training data from web [52, 53]. Apart from applying additional filters or safety checkers [47, 41, 48], prevailing methods modify diffusion model parameters to erase specific target concepts, mainly categorized into two paradigms. The training-based paradigm fine-tunes model parameters with specific erasure objectives [29, 17, 66] and additional regularization terms [29, 37, 35]. In contrast, the editing-based paradigm edits model parameters using a closed-form solution to facilitate efficiency in concept erasure. For example, UCE [18] modifies model weights by balancing both erasure and preservation error through a weighted least squares objective and RECE [19] iteratively derives new target concept embeddings. These methods can erase numerous concepts within seconds, demonstrating superior efficiency in practice.
Null-space constraints. The null space of a matrix, a fundamental concept in linear algebra, refers to the set of all vectors that the matrix maps to the zero vector. The null-space constraints are first applied to continual learning (CL) by projecting gradients onto the null space of uncentered covariances from previous tasks [60]. Subsequent studies [36, 61, 63, 28, 32] further explore and extend the application of null space in CL. In model editing, AlphaEdit [14] restricts model weight updates onto the null space of preserved knowledge, effectively mitigating trade-offs between editing and preservation. The null-space constraints also apply to various other tasks, e.g., machine unlearning [9] and MRI reconstruction [15], offering great promise for editing-based concept erasure.
3 Problem Formulation
3.1 Concept Erasure in Closed-Form Solution
In T2I diffusion models, each concept is encoded by a set of text tokens via CLIP [45], which are then aggregated into a single concept embedding $\boldsymbol{c}\in\mathbb{R}^{d_{0}}$. For concept erasure, there are two sets of concepts: the erasure set $\mathbf{E}$ and the retain set $\mathbf{R}$. The erasure set consists of $N_{E}$ target concepts to be removed, denoted as $\mathbf{E}=\{\boldsymbol{c}_{1}^{(i)}\}_{i=1}^{N_{E}}$. The retain set includes $N_{R}$ non-target concepts that should be preserved during editing, denoted as $\mathbf{R}=\{\boldsymbol{c}_{0}^{(i)}\}_{i=1}^{N_{R}}$. To effectively erase each target concept $\boldsymbol{c}_{1}^{(i)}\in\mathbf{E}$ (e.g., Snoopy), it is mapped onto an anchor concept $\boldsymbol{c}_{*}^{(i)}$ that retains general semantics while avoiding the original concept-specific information (e.g., Dog). The paired erasure set is denoted as: $\mathbf{\tilde{E}}=\{(\boldsymbol{c}_{1}^{(i)},\boldsymbol{c}_{*}^{(i)})\}_{i=1}^{N_{E}}$.
For editing-based methods [42, 18, 19], concept embeddings from the erasure set $\mathbf{\tilde{E}}$ and the retain set $\mathbf{R}$ are first organized into three structured matrices: $\mathbf{C}_{1},\mathbf{C}_{*}\in\mathbb{R}^{d_{0}\times N_{E}}$ and $\mathbf{C}_{0}\in\mathbb{R}^{d_{0}\times N_{R}}$, which represents the stacked embeddings of the target, anchor, and non-target concepts, respectively. To derive a closed-form solution for concept erasure, existing methods typically optimize a perturbation $\boldsymbol{\Delta}$ to model parameters $\mathbf{W}$, balancing between erasure efficacy and prior preservation. For example, UCE [18] formulates concept erasure as a weighted least squares problem:
$\boldsymbol{\Delta}_{\text{UCE}}=\mathop{\text{arg\ min}}\limits_{\boldsymbol{\Delta}}\underbrace{\|(\mathbf{W}+\boldsymbol{\Delta})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}}_{\text{$e_{1}$}}+\underbrace{\|\boldsymbol{\Delta}\mathbf{C}_{0}\|^{2}}_{\text{$e_{0}$}},$ (1)
where the erasure error $e_{1}$ ensures that each target concept is mapped onto its corresponding anchor concept and the preservation error $e_{0}$ minimizes the impact on non-target priors. This formulation provides a closed-form solution to obtain $\boldsymbol{\Delta}_{\text{UCE}}$ (see Appx. B.1) for parameter updates, achieving computationally efficient optimization. However, as the number of target concepts increases, the accumulated preservation errors $e_{0}$ across multiple erased concepts would amplify the distortion on non-target knowledge, further degrading prior preservation.
3.2 Null-Space Constraints
To mitigate the limitation of weighted optimization in prior preservation, SPEED integrates null-space constraints to achieve model editing [14]. These constraints restrict the solution space of $\boldsymbol{\Delta}$ within the null space of $\mathbf{C}_{0}$ to ensure that parameter updates do not interfere with non-target concepts, enforcing $e_{0}$ close to zero without a tradeoff between $e_{0}$ and $e_{1}$. To obtain $\boldsymbol{\Delta}$, we first perform singular value decomposition (SVD) on $\mathbf{C}_{0}\mathbf{C}_{0}^{\top}$***Considering $\mathbf{C}_{0}\in\mathbb{R}^{d_{0}\times N_{R}}$ may have high dimensionality depending on $N_{R}$, we compute the null space by performing SVD on $\mathbf{C}_{0}\mathbf{C}_{0}^{\top}\in\mathbb{R}^{d_{0}\times d_{0}}$, as both $\mathbf{C}_{0}$ and $\mathbf{C}_{0}\mathbf{C}_{0}^{\top}$ share the same null space. and construct a projection matrix $\mathbf{P}$:
$\left\{\mathbf{U},\mathbf{\Lambda},\mathbf{U}^{\top}\right\}=\text{SVD}\left(\mathbf{C}_{0}\mathbf{C}_{0}^{\top}\right),\quad\mathbf{P}=\hat{\mathbf{U}}\hat{\mathbf{U}}^{\top},$ (2)
where $\hat{\mathbf{U}}$ is the matrix composed of eigenvectors w.r.t. the zero singular values in $\mathbf{\Lambda}$. This projection matrix $\mathbf{P}$ can map $\boldsymbol{\Delta}$ onto the null space of $\mathbf{C}_{0}$, i.e., $\|\left(\boldsymbol{\Delta}\mathbf{P}\right)\mathbf{C}_{0}\|^{2}=0$. Thus, the objective becomes:
$\left(\boldsymbol{\Delta}\mathbf{P}\right)_{\text{Null}}=$ $\mathop{\text{arg\ min}}\limits_{\boldsymbol{\Delta}}\underbrace{\|(\mathbf{W}+\boldsymbol{\Delta}\mathbf{P})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}}_{\text{$e_{1}$}}$
[…]
where $\|\boldsymbol{\Delta}\mathbf{P}\|^{2}$ is an additional regularization term to guarantee convergence. This objective facilitates model parameter editing in the null space of the retain set $\mathbf{R}$, showing great promise for prior-preserved concept erasure.
Figure 3: Prior shift comparison with erasing different target concepts. The prior shift is measured by $\|\boldsymbol{\Delta}_{\text{erase}}\boldsymbol{c}\|^{2}$ introduced in Sec. 4.1, where $\boldsymbol{c}$ denotes the concept embedding. A larger value indicates a greater impact on the concept after erasure.
4 Prior Knowledge Refinement
The core of null-space constraints in SPEED is to determine a projection matrix $\mathbf{P}$ that constrains the solution space of $\boldsymbol{\Delta}$ within the null space of the retain set $\mathbf{R}$, ensuring that non-target concepts remain unaffected after erasure. The rank of the matrix $\mathbf{C}_{0}\mathbf{C}_{0}^{\top}$ reflects the diversity of the retain set, directly shaping the structure of its null space, which is determined by $\mathbf{R}$. As $\mathbf{R}$ increases, $\mathbf{C}_{0}\mathbf{C}_{0}^{\top}$ gradually reaches full rank, its null space narrows and reduces to the trivial null space $\{\mathbf{0}\}$. However, shrinking $\mathbf{R}$ would eliminate prior coverage by allowing more unintended shifts in non-target concepts not included. To balance this dilemma, we introduce three complementary techniques: Influence-Based Prior Filtering (IPF) to retain the most affected non-target concepts while avoiding excessive rank saturation, Directed Prior Augmentation (DPA) to expand prior coverage with semantically consistent variations, and Invariant Equality Constraints (IEC) to enforce additional constraints to preserve key invariants in generation.
4.1 Influence-Based Prior Filtering (IPF)
Given a pre-defined retain set, existing editing-based methods [18, 19] generally neglect to account for the varying influence of different target concepts on non-target elements, where closer semantic relationships (e.g., Mickey and Dog for Snoopy) lead to greater prior shifts, as shown in Fig. 3. This differing influence suggests that not all non-target concepts contribute equally to preserving prior knowledge, and weakly influenced concepts provide marginal benefit while introducing additional ranks to constrain the null space. To improve, we propose to filter the original retain set by selectively retaining those non-target concepts most influenced by the erased target concepts.
To evaluate this influence, we first obtain the closed-form solution $\boldsymbol{\Delta}_{\text{erase}}$, which only considers the erasure efficacy during editing. Specifically, we remove the preservation error $e_{0}$ from Eq. 1 and introduce an additional regularization term $\|\boldsymbol{\Delta}\|^{2}$ to guarantee convergence:
$\boldsymbol{\Delta}_{\text{erase}}$ $=\mathop{\text{arg\ min}}\limits_{\boldsymbol{\Delta}}\underbrace{\|(\mathbf{W}+\boldsymbol{\Delta})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}}_{e_{1}}+\underbrace{\|\boldsymbol{\Delta}\|^{2}}_{\text{regularization}}$
$=\mathbf{W}\left(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\right)\left(\mathbf{I}+\mathbf{C}_{1}\mathbf{C}_{1}\right)^{-1}.$ (4)
Then we can quantify the influence on each concept embedding $\boldsymbol{c}$ by measuring the prior shift $=\|\boldsymbol{\Delta}_{\text{erase}}\boldsymbol{c}\|^{2}$. This metric stems from the intuition that after erasing the target concept, a larger prior shift in the non-target concept indicates a greater influence on the corresponding prior knowledge. Building upon this, we can filter the original retain set $\mathbf{R}$ with the prior shift by the following process:
$\mathbf{R}_{f}:\mathbf{R}\mapsto\{\boldsymbol{c}_{0}\in\mathbf{R}\mid\|\boldsymbol{\Delta}_{\text{erase}}\boldsymbol{c}_{0}\|^{2}>\mu\},$ (5)
where the mean value $\mu=\mathbb{E}_{\boldsymbol{c}_{0}\sim\mathbf{R}}\left[\|\boldsymbol{\Delta}_{\text{erase}}\boldsymbol{c}_{0}\|^{2}\right]$ serves as the filtering threshold.
4.2 Directed Prior Augmentation (DPA)
To prevent IPF from generating an overly small retain set, which would compromise prior coverage, an intuitive strategy is to augment the prior knowledge by perturbing the non-target embedding $\boldsymbol{c}_{0}$ with random noise [37]. However, this strategy would introduce meaningless embeddings that fail to generate semantically coherent images (e.g., noise image), resulting in excessive preservation with increasing ranks. To search for more semantically consistent concepts, we introduce directed noise by projecting the random noise $\boldsymbol{\epsilon}$ onto the direction in which the model parameters $\mathbf{W}$ exhibit minimal variation. This operation ensures that the perturbed embeddings express closer semantics to the original concept after being mapped by $\mathbf{W}$, as shown in Fig. 4. Specifically, we first derive a projection matrix $\mathbf{P}_{\text{min}}$:
$\left\{\mathbf{U}_{\mathbf{W}},\mathbf{\Lambda}_{\mathbf{W}},\mathbf{U}_{\mathbf{W}}^{\top}\right\}=\text{SVD}\left(\mathbf{W}\right),\quad\mathbf{P}_{\text{min}}=\mathbf{U}_{\text{min}}\mathbf{U}_{\text{min}}^{\top},$ (6)
where $\mathbf{U}_{\text{min}}={\mathbf{U}}_{\mathbf{W}}[:,-r:]$ denotes the eigenvectors w.r.t. the smallest $r$ singular values, which represent the $r$ least-changing directions of $\mathbf{W}$. Then the directed noise $\boldsymbol{\epsilon}\cdot\mathbf{P}_{\text{min}}$ is used to perturb the original embedding via:
$\boldsymbol{c}^{\prime}_{0}=\boldsymbol{c}_{0}+\boldsymbol{\epsilon}\cdot\mathbf{P}_{\text{min}},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$ (7)
Given a retain set $\mathbf{R}$, the augmentation process can be formulated as follows:
$\mathbf{R}^{\text{aug}}:\mathbf{R}\mapsto\bigcup_{\boldsymbol{c}_{0}\in\mathbf{R}}\left\{\boldsymbol{c}_{0,k}^{\prime}\mid k=1,\dots,N_{A}\right\},$ (8)
where $N_{A}$ denotes the augmentation times and $\boldsymbol{c}_{0,k}^{\prime}$ represents the $k$-th augmented embedding given $\boldsymbol{c}_{0}\in\mathbf{R}$ using Eq. 7. In implementation, we first filter the original retain set $\mathbf{R}$ to obtain $\mathbf{R}_{f}$ using IPF. Subsequently, further augmentation and filtering are applied to $\mathbf{R}_{f}$ using DPA and IPF to obtain $\left(\mathbf{R}_{f}\right)^{\text{aug}}_{f}$, and the two filtered retain sets are then combined together to serve as the final refined retain set $\mathbf{R}_{\text{refine}}=\mathbf{R}_{f}\cup\left(\mathbf{R}_{f}\right)^{\text{aug}}_{f}$.
Figure 4: t-SNE distribution of perturbing the original concept with random noise and our directed noise. (a) Similar to random noise, our method can span a broad concept embedding space. (b) Our directed noise preserves semantic similarity to the original concept with closer distances in the space mapped by $\mathbf{W}$.
Figure 5: Qualitative comparison of the few-concept erasure in erasing single and multiple instances. The erased and preserved generations are highlighted with red and green boxes, respectively. Our method demonstrates superior prior preservation compared to the SOTA method, UCE [18], in single-concept erasure, and this superiority is amplified as the target concepts increase. For example, in the Hello Kitty case, our generation exhibits less alteration after erasing three concepts, whereas UCE’s generations are gradually deteriorating.
4.3 Invariant Equality Constraints (IEC)
In parallel, we also identify certain invariants during the T2I generation process, i.e., intermediate variables that remain unchanged with sampling prompts. One such invariant is the CLIP-encoded [SOT] token. Since the encoding process is masked by causal attention and all prompts are prefixed with the fixed [SOT] token during tokenization, the corresponding embedding remains unchanged. Another invariant is the null-text embedding, as it corresponds to the unconditional generation under the classifier-free guidance [23], which also remains unchanged despite prompt variations. Given the invariance of these embeddings, we consider additional protection measures for these invariant embeddings to ensure their outputs remain unchanged during concept erasure. Specifically, we introduce explicit equality constraints for invariants based on Eq. 3.2:
$\begin{gathered}\mathop{\text{min}}\limits_{\boldsymbol{\Delta}}\underbrace{\|(\mathbf{W}+\boldsymbol{\Delta}\mathbf{P})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}}_{e_{1}}+\underbrace{\|\boldsymbol{\Delta}\mathbf{P}\|^{2}}_{\text{regularization}},\\ \quad\text{s.t.}\,\underbrace{(\boldsymbol{\Delta}\mathbf{P})\mathbf{C}_{2}=\mathbf{0}}_{\text{equality constraints}},\end{gathered}$ (9)
where $\mathbf{C}_{2}$ denotes the stacked invariant embedding matrix of [SOT] and null-textSince the null-text embeddings are only composed of [EOT] tokens (excluding [SOT]), we use the k-means algorithm [38] to select k centroids to reduce redundancy.. With the projection matrix $\mathbf{P}$ derived from $\mathbf{R}_{\text{refine}}$, we can compute the closed-form solution of Eq. 9 using Lagrange Multipliers from Appx. B.3:
$\left(\boldsymbol{\Delta}\mathbf{P}\right)_{\text{SPEED}}=\mathbf{W}\left(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\right)\mathbf{P}\mathbf{Q}\mathbf{M},$ (10)
where
$\begin{gathered}\mathbf{M}=\left(\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\mathbf{P}+\mathbf{I}\right)^{-1},\\ \mathbf{Q}=\mathbf{I}-\mathbf{M}\mathbf{C}_{2}\left(\mathbf{C}_{2}^{\top}\mathbf{P}\mathbf{M}\mathbf{C}_{2}\right)^{-1}\mathbf{C}_{2}^{\top}\mathbf{P}.\end{gathered}$ (11)
This closed-form solution enforces the equality constraints by projecting the parameter update onto the subspace orthogonal to the invariant embeddings. Since image generation inevitably depends on these invariant embeddings, such constraints inherently preserve prior knowledge.
5 Experiments
In this section, we conduct extensive experiments on three representative erasure tasks, including few-concept erasure, multi-concept erasure, and implicit concept erasure, validating our superior prior preservation. The compared baselines include ConAbl [29], MACE [35], RECE [19], and UCE [18], which have achieved SOTA performance across various concept erasure tasks. In implementation, we conduct all experiments on SDv1.4 [1] and generate each image using DPM-solver sampler [34] over 20 sampling steps with classifier-free guidance [23] of 7.5. More implementation details and compared baselines (e.g., SPM [37]) can be found in Appx. C and Appx. D.2.
5.1 On Few-Concept Erasure
Evaluation setup. To compare the few-concept erasure performance with baseline methods, we conduct experiments on instance erasure and artistic style erasure following [37], where all methods are evaluated based on 80 instance templates and 30 artistic style templates, generating 10 images per template per concept. We use two metrics for evaluation: CLIP Score (CS) [45] measuring the text-image similarity and Fréchet Inception Distance (FID) [22] assessing the distributional distance before and after erasure. Following [37], we select non-target concepts with similar semantics to the target concept for comparison and report CS for targets and FID for non-targets in the main paper. Full comparisons are presented in Appx. D.1. We further compare the generations on MS-COCO captions [33], where we generate images with the first 1,000 captions, and report the CS and FID as the measure of general knowledge preservation.
Concept Snoopy Mickey Spongebob Pikachu Hello Kitty MS-COCO
CS CS CS CS CS CS FID
SD v1.4 28.51 26.62 27.30 27.44 27.77 26.53 -
Erase Snoopy
CS ↓ FID ↓ FID ↓ FID ↓ FID ↓ CS ↑ FID ↓
ConAbl 25.44 37.08 38.92 26.14 36.52 26.40 21.20
MACE 20.90 105.97 102.77 65.71 75.42 26.09 42.62
RECE 18.38 26.63 34.42 21.99 32.35 26.39 25.61
UCE 23.19 24.87 29.86 19.06 27.86 26.46 22.18
Ours 23.50 23.41 24.64 16.81 21.74 26.48 19.95
Erase Snoopy and Mickey
CS ↓ CS ↓ FID ↓ FID ↓ FID ↓ CS ↑ FID ↓
ConAbl 25.26 26.58 45.08 35.57 41.48 26.42 24.34
MACE 20.53 20.63 112.01 91.72 106.88 25.50 55.15
RECE 18.57 19.14 35.85 26.05 40.77 26.31 30.30
UCE 23.60 24.79 30.58 23.51 31.76 26.38 26.06
Ours 23.58 23.62 29.67 22.51 28.23 26.47 23.66
Erase Snoopy and Mickey and Spongebob
CS ↓ CS ↓ CS ↓ FID ↓ FID ↓ CS ↑ FID ↓
ConAbl 24.92 26.46 25.12 46.47 48.24 26.37 26.71
MACE 19.86 19.35 20.12 110.12 128.56 23.39 66.39
RECE 18.17 18.87 16.23 40.52 52.06 26.32 32.51
UCE 23.29 24.63 19.08 29.20 38.15 26.30 28.71
Ours 23.69 23.93 21.39 21.40 26.22 26.51 24.99
Table 1: Quantitative comparison of the few-concept erasure in erasing instances, where the erased target concept(s) are shaded in pink. Arrows on the headers indicate the preferred direction for each metric, and the best results are highlighted in bold. Our method, which does not achieve the lowest CS but has been proven sufficient in Fig. 5, consistently demonstrates significant improvements in prior preservation for non-target concepts.
Analysis and discussion. Tables 1 and 2 compare the results of erasing various instance concepts and artistic styles. Our method consistently achieves the lowest FIDs across all non-target concepts, demonstrating superior prior preservation with minimal alteration to the original content. Moreover, we emphasize that our erasure is sufficiently effective, even without achieving the lowest CS, as shown in Fig. 5. In contrast, lower CS values typically indicate over-erasure, which results in excessive degradation of prior knowledge. Notably, with the number of target concepts increasing from 1 to 3, our FID in Pikachu rises from 16.81 to 21.40 (4.59 ↑), while UCE increases from 19.06 to 29.20 (10.14 ↑). A similar pattern is observed in Hello Kitty (Our 4.48 ↑ v.s. UCE’s 10.29 ↑), showing the robustness of our method in erasing increasing multiple concepts.
5.2 On Multi-Concept Erasure
Evaluation setup. Another more realistic erasure scenario is multi-concept erasure, where massive concepts are required to be erased at once. Herein, we follow the experiment setup in [35] for erasing multiple celebrities, where we experiment with erasing 10, 50, and 100 celebrities and collect another 100 celebrities as non-target concepts. We prepare 5 prompt templates for each celebrity concept. For non-target concepts, we generate 1 image per template for each of the 100 concepts, totaling 500 images. For target concepts, we adjust the per-concept quantity to maintain a total of 500 images (e.g., erasing 10 celebrities involves generating 10 images with 5 templates per concept). In evaluation, we adopt GIPHY Celebrity Detector (GCD) [20] and measure the top-1 GCD accuracy, indicated by $\text{Acc}_{e}$ for erased target concepts and $\text{Acc}_{r}$ for retained non-target concepts. Meanwhile, the harmonic mean $H_{o}$ is adopted to assess the overall erasure performance:
$H_{o}=\frac{2}{\left(1-\text{Acc}_{e}\right)^{-1}+\left(\text{Acc}_{r}\right)^{-1}}.$ (12)
Additionally, we report the results on MS-COCO to demonstrate the prior preservation of general concepts.
Concept Van Gogh Picasso Monet Paul Gauguin Caravaggio MS-COCO
CS CS CS CS CS CS FID
SD v1.4 28.75 27.98 28.91 29.80 26.27 26.53 -
Erase Van Gogh
CS ↓ FID ↓ FID ↓ FID ↓ FID ↓ CS ↑ FID ↓
ConAbl 28.16 77.01 63.80 63.20 79.25 26.46 18.36
MACE 26.66 69.92 60.88 56.18 69.04 26.50 23.15
RECE 26.39 60.57 61.09 47.07 72.85 26.52 23.54
UCE 28.10 43.02 40.49 32.62 61.72 26.54 19.63
Ours 26.29 35.86 16.85 24.94 39.75 26.55 20.36
Erase Picasso
FID ↓ CS ↓ FID ↓ FID ↓ FID ↓ CS ↑ FID ↓
ConAbl 60.44 26.97 36.23 65.23 79.12 26.43 20.02
MACE 59.58 26.48 37.02 46.35 66.20 26.47 22.86
RECE 51.09 26.66 25.39 46.08 75.61 26.48 23.03
UCE 37.58 26.99 16.72 32.48 59.27 26.50 20.33
Ours 19.18 26.22 19.87 24.73 43.63 26.51 19.98
Erase Monet
FID ↓ FID ↓ CS ↓ FID ↓ FID ↓ CS ↑ FID ↓
ConAbl 68.77 64.25 27.05 57.33 71.88 26.45 21.03
MACE 61.50 48.41 25.98 49.66 65.87 26.47 22.76
RECE 56.26 45.97 25.87 46.38 64.19 26.49 24.94
UCE 42.25 38.73 27.12 33.00 56.49 26.51 21.58
Ours 28.78 41.21 25.06 27.85 55.20 26.48 20.87
Table 2: Quantitative comparison of the few-concept erasure in erasing single artistic style.
Erase 10 Celebrities MS-COCO Erase 50 Celebrities MS-COCO Erase 100 Celebrities MS-COCO
$\text{Acc}_{e}$ $\text{Acc}_{r}$ $H_{o}$ CS ↑ FID ↓ $\text{Acc}_{e}$ $\text{Acc}_{r}$ $H_{o}$ CS ↑ FID ↓ $\text{Acc}_{e}$ $\text{Acc}_{r}$ $H_{o}$ CS ↑ FID ↓
SD v1.4 91.99 89.66 14.70 26.53 - 93.08 89.66 12.85 26.53 - 90.18 89.66 17.70 26.53 -
UCE 0.20 71.19 83.10 24.07 83.81 0.00 31.94 48.41 13.45 209.93 0.00 20.92 34.60 13.49 185.46
RECE 0.34 67.43 80.44 16.75 170.65 1.03 19.77 32.95 13.49 213.39 2.43 23.71 38.16 12.09 177.57
MACE 1.62 87.73 92.75 26.36 37.25 3.41 84.31 90.03 25.45 45.31 4.80 80.20 87.06 24.80 50.41
Ours 1.81 89.09 93.42 26.47 30.02 3.46 88.48 92.34 26.46 39.23 5.87 85.54 89.63 26.22 44.97
Table 3: Quantitative comparison of the multi-concept erasure in erasing 10, 50, and 100 celebrities. The best results are highlighted in bold. Our method is capable of erasing up to 100 celebrities at once with low $\text{Acc}_{e}$ (%) and preserving other non-target celebrities with less appearance alteration with high $\text{Acc}_{r}$ (%), resulting in the best overall erasure performance $H_{o}$ (shaded in pink).
NudeNet Detection Results on I2P MS-COCO
Armpits Belly Buttocks Feet Breasts (F) Genitalia (F) Breasts (M) Genitalia (M) Total ↓ CS ↑ FID ↓
SD v1.4 123 134 19 14 258 9 16 3 576 26.53 -
ConAbl 24 43 5 6 68 1 6 4 157 26.14 39.26
MACE 28 19 1 20 37 3 6 5 119 24.06 52.78
RECE 17 29 3 7 14 1 8 1 80 25.98 40.37
UCE 29 42 2 11 36 3 9 7 139 26.24 38.60
Ours 20 42 7 3 29 2 5 5 113 26.29 37.82
Table 4: Evaluation of implicit concept erasure on I2P benchmark. We report the number of nude body parts (F: Female, M: Male) detected by the NudeNet with threshold $=$ 0.6. The best and second-best results are marked in bold and underlined. (Left) Our method effectively removes nude content, even though nudity is not explicitly mentioned in prompts from I2P, achieving the second-best total count. (Right) Our method also consistently achieves superior prior preservation for non-target concepts to other methods on MS-COCO.
Analysis and discussion. Table 3 showcases a notable improvement of our method on multi-concept erasure, particularly in prior preservation with the highest $\text{Acc}_{r}$. In comparison with the SOTA method, MACE [35], our method achieves superior prior preservation with better $\text{Acc}_{r}$, while maintaining comparable erasure efficacy, as reflected in similar $\text{Acc}_{e}$, resulting in the best overall erasure performance indicated by the highest $H_{o}$. Meanwhile, our method attains the lowest FID across all methods on MS-COCO. The other methods, UCE [18] and RECE [19], although achieving considerable balance in few-concept erasure, fail to maintain this balance as the number of target concepts increases as shown in Fig. 6, with catastrophic prior damage evidenced by MS-COCO as well. Notably, our method can erase up to 100 celebrities in 5 seconds, whereas MACE requires around 30 minutes (×$\times 350$ time). In real-world scenarios, this efficiency underscores our potential for the instant erasure of massive concepts.
5.3 On Implicit Concept Erasure
Evaluation setup. We further evaluate the erasure efficacy on implicit concepts, where the target concept does not explicitly appear in the text prompt. We conduct experiments on the Inappropriate Image Prompt (I2P) benchmark [51], which consists of various implicit inappropriate prompts involving violence, sexual content, and nudity. We follow the same setting in [19] to erase nudity $\to$ ‘ ’. Specifically, we generate images using all 4,703 text prompts in I2P and use NudeNet [6] to identify if the nude content is successfully erased with the threshold of 0.6. Additionally, we report the results on MS-COCO to demonstrate the prior preservation of general concepts.
Analysis and discussion. As shown in Table 4, our method can effectively erase the implicit concept, i.e., nudity, with the second-best number of detected nude body parts. The SOTA method, RECE [19], achieves the best total number by extending the erasure set with more target concepts, but this comes at the cost of sacrificing prior preservation on MS-COCO. In contrast, our method achieves the best prior preservation, demonstrating effective erasure while maintaining strong prior preservation, striking a favorable balance between erasure and preservation.
Figure 6: Quantitative comparison of multi-concept erasure in erasing different numbers of celebrities (celeb). The erased and preserved generations are highlighted with red and green boxes, respectively. Our method can precisely erase 100 celebrities while maintaining the proper generation of other non-target concepts.
5.4 Further Analysis
Duration comparison. Table 5 presents the duration comparison in erasing 1, 10, and 100 concepts across different methods. It is obvious that training-based methods necessitate significantly higher computational costs than editing-based ones. In contrast, our method achieves precise multi-concept erasure in a remarkably short time, demonstrating superior efficiency while maintaining erasure performance, as evidenced in Table 3.
Component ablation. In Table 6, we compare the individual impact of our components on prior preservation and draw the following conclusions: (1) Impact of IEC (Ablation 1 v.s. 2): IEC reduces the non-target FID and the MS-COCO FID, demonstrating its effectiveness by preserving invariant embeddings with equality constraints. (2) Impact of IPF (Ablation 2 v.s. 3): Incorporating IPF results in a significant improvement in both FIDs, underscoring its critical role in filtering out less-influenced concepts in the retain set under rank saturation. (3) Impact of DPA (Ablation 4 v.s. Ours): DPA improves RPA with directed noise and leads to a substantial improvement in non-target and MS-COCO FIDs, highlighting its advantage by introducing semantically similar concepts into the refined retain set. To conclude, the proposed three components (i.e., IEC, IPF, and DPA) improve the prior preservation from different perspectives and contribute to our method with the best prior preservation under null space constraints.
Training-based Editing-based
ConAbl MACE UCE RECE Ours
Data Prepare $n\times 1000t_{1}$ $n\times(8t_{1}+8t_{2})$ 0 0 0
1 concept $1\times 90$ 55.1 1.2 1.5 3.6
10 concepts $10\times 90$ 207.0 1.5 2.5 3.8
100 concepts $100\times 90$ 1750.9 2.1 11.0 5.0
Table 5: Duration comparison (s) in erasing multiple celebrities on one A100 GPU, where $n$ is the number of target concepts. During data preparation, ConAbl requires pre-sampling 1,000 images ($t_{1}$ per image) for each target concept while MACE needs 8 pre-sampled images along with 8 corresponding segmentation masks ($t_{2}$ per mask) using SAM [27].
Ablation Components Target Non-Target MS-COCO
IEC IPF RPA DPA CS ↓ FID ↓ CS ↑ FID ↓
1 $\times$ $\times$ $\times$ $\times$ 27.20 50.43 26.42 26.33
2 $\times$ $\times$ $\times$ 27.20 48.17 26.44 24.95
3 $\times$ $\times$ 26.68 38.02 26.54 20.57
4 $\times$ 26.30 32.62 26.52 20.99
Ours $\times$ 26.29 29.35 26.55 20.36
SD v1.4 - - - - 28.75 - 26.53 -
Table 6: Ablation study on the proposed components in erasing Van Gogh, with the non-target FID averaged over other four styles (i.e., Picasso, Monet, Paul Gauguin, Caravaggio) as in Table 2. Ablation 1 corresponds to the original objective from [14] in Eq. 3.2. The ablated components include: IEC (Invariant Equality Constraints), IPF (Influence-based Prior Filtering), RPA (Random Prior Augmentation), and DPA (Directed Prior Augmentation).
Figure 7: More applications across various T2I diffusion models. (a) We conduct both single and composite concept erasure for “Snoopy + Van Gogh” on two community T2I models [3, 4]. (b) Our method also enables model knowledge editing by specifying the anchor concept, such as “Wonder WomanWoman” and “SupermanBatman” on SDXL [44]. (c) Our method can seamlessly transfer to novel DiT-based T2I models (e.g., SDv3 [13]).
More applications on other T2I models. To validate the transferability of our method across versatile applications, we conduct further experiments on various T2I models with different weights and architectures, including: (1) Composite concept erasure on DreamShaper [3] and RealisticVision [4] from Fig 7 (a): Our method can precisely erase the target concept(s) while preserving other non-target elements within the prompt, such as the Van Gogh-style background (2nd column) and the Snoopy character (3rd column). (2) Knowledge editing on SDXL [44] from Fig 7 (b): The arbitrary nature of anchor concepts allows us to edit the pre-trained model knowledge. Herein, our method effectively edits the model knowledge while maintaining the overall layout and semantics of the generated images. (3) Instance erasure on SDv3 [13] from Fig 7 (c): To accommodate the diffusion transformer (DiT) [43] architecture in T2I models, we adapt our method to a DiT-based model, demonstrating a well-balanced trade-off between erasure (1st row) and preservation (2nd row) as well.
6 Conclusion
This paper introduced SPEED, a scalable, precise, and efficient concept erasure method for T2I diffusion models. It formulates concept erasure as a null-space constrained optimization problem, facilitating effective prior preservation along with precise erasure efficacy. Critically, SPEED overcomes the inefficacy of editing-based methods in multi-concept erasure while circumventing the prohibitive computational costs associated with training-based approaches. With our proposed Prior Knowledge Refinement involving three complementary techniques, SPEED not only ensures superior prior preservation but also achieves a 350×350$\times$350 × acceleration in multi-concept erasure, establishing itself as a scalable and practical solution for real-world applications.
Appendix A Preliminaries
T2I diffusion models. T2I generation has seen significant advancements with diffusion models, particularly Latent Diffusion Models (LDMs) [49]. Unlike pixel-space diffusion, LDMs operate in the latent space of a pretrained autoencoder, reducing computational costs while maintaining high-quality synthesis. LDMs consist of a vector-quantized autoencoder [59, 12] and a diffusion model [11, 24, 55, 26, 57]. The autoencoder encodes an image $\boldsymbol{x}$ into a latent representation $\boldsymbol{z}=\mathcal{E}(\boldsymbol{x})$ and reconstructs it via $\boldsymbol{x}\approx\mathcal{D}(\boldsymbol{z})$. The diffusion model learns to generate latent codes through a denoising process. The training objective is given by [24, 49]:
$\mathcal{L}_{\text{LDM}}=\mathbb{E}_{\boldsymbol{z}\sim\mathcal{E}(\boldsymbol{x}),\boldsymbol{c},\boldsymbol{\epsilon}\sim\mathcal{N}(0,1),t}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\boldsymbol{z}_{t},t,\boldsymbol{c})\right\|^{2}_{2}\right],$ (13)
where $\boldsymbol{z}_{t}$ is the noisy latent at timestep $t$, $\boldsymbol{\epsilon}$ is Gaussian noise, $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}$ is the denoising network, and $\boldsymbol{c}$ is conditioning information from text, class labels, or segmentation masks [49]. During inference, a latent $\boldsymbol{z}_{T}$ is sampled from a Gaussian prior and progressively denoised to obtain $\boldsymbol{z}_{0}$, which is then decoded into an image via $\boldsymbol{x}_{0}\approx\mathcal{D}(\boldsymbol{z}_{0})$.
Cross-attention mechanisms. Current T2I diffusion models usually leverage a generative framework to synthesize images conditioned on textual descriptions in latent space [49]. The conditioning mechanism is implemented through cross-attention (CA) layers. Specifically, textual descriptions are first tokenized into $n$ tokens and embedded into a sequence of vectors $\boldsymbol{e}\in\mathbb{R}^{d_{0}\times n}$ via a pre-trained CLIP model [45]. These text embeddings serve as the key $\mathbf{K}\in\mathbb{R}^{n\times d_{k}}$ and value $\mathbf{V}\in\mathbb{R}^{n\times d_{v}}$ inputs using parametric projection matrices $\mathbf{W_{K}}\in\mathbb{R}^{d_{k}\times d_{0}}$ and $\mathbf{W_{V}}\in\mathbb{R}^{d_{v}\times d_{0}}$, while the intermediate image representations act as the query $\mathbf{Q}\in\mathbb{R}^{m\times d_{k}}$. The cross-attention mechanism is defined as:
$\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d_{k}}}\right)\mathbf{V}.$ (14)
This alignment enables the model to capture semantic correlations between the textual input and the visual features, ensuring that the generated images are semantically consistent with the provided text prompts.
Appendix B Proof and Derivation
B.1 Deriving the Closed-Form Solution for UCE
From Eq. 1, we are tasked with minimizing the following editing objective, where the hyperparameters $\alpha$ and $\beta$ correspond to the weights of the erasure error $e_{1}$ and the preservation error $e_{0}$, respectively:
$\mathop{\text{min}}\limits_{\boldsymbol{\Delta}}\left[\alpha\|(\mathbf{W}+\boldsymbol{\Delta})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}+\beta\|\boldsymbol{\Delta}\mathbf{C}_{0}\|^{2}\right].$ (15)
To derive the closed-form solution, we begin by computing the gradient of the objective function with respect to $\boldsymbol{\Delta}$. The gradient is given by:
$\alpha\left(\mathbf{W}\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}+\boldsymbol{\Delta}\mathbf{C}_{1}\right)\mathbf{C}_{1}^{\top}+\beta\boldsymbol{\Delta}\mathbf{C}_{0}\mathbf{C}_{0}^{\top}=0.$ (16)
Solving the resulting equation yields the closed-form solution for $\boldsymbol{\Delta}_{\text{UCE}}$:
$\boldsymbol{\Delta}_{\text{UCE}}=\alpha\mathbf{W}\left(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\right)\left(\alpha\mathbf{C}_{1}\mathbf{C}_{1}^{\top}+\beta\mathbf{C}_{0}\mathbf{C}_{0}^{\top}\right)^{-1}.$ (17)
In practice, an additional identity matrix $\mathbf{I}$ with hyperparameter $\lambda$ is added to $\left(\alpha\mathbf{C}_{1}\mathbf{C}_{1}^{\top}+\beta\mathbf{C}_{0}\mathbf{C}_{0}^{\top}\right)^{-1}$ to ensure its invertibility. This modification results in the following closed-form solution for UCE:
$\boldsymbol{\Delta}_{\text{UCE}}=\alpha\mathbf{W}\left(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\right)\left(\alpha\mathbf{C}_{1}\mathbf{C}_{1}^{\top}+\beta\mathbf{C}_{0}\mathbf{C}_{0}^{\top}+\lambda\mathbf{I}\right)^{-1}.$ (18)
B.2 Proof of the Lower Bound of $e_{0}$ for UCE
Herein, we aim to establish the existence of a strictly positive constant $c>0$ such that
$e_{0}=\|\boldsymbol{\Delta}_{\text{UCE}}\mathbf{C}_{0}\|^{2}=\|\alpha\mathbf{W}\left(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\right)(\alpha\mathbf{C}_{1}\mathbf{C}_{1}^{\top}+\beta\mathbf{C}_{0}\mathbf{C}_{0}^{\top}+\lambda\mathbf{I})^{-1}\mathbf{C}_{0}\|^{2}\geq c>0.$ (19)
Assumption B.1.
We assume that $\alpha,\beta,\lambda\neq 0$, that $\mathbf{W}$ is a full-rank matrix, and that $\mathbf{C}_{0}\mathbf{C}_{0}^{\top}$ is rank-deficient. Furthermore, we assume that
$\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\neq\mathbf{0}.$
Proof.
Define the matrix $\mathbf{M}$ as
$\mathbf{M}=\alpha\mathbf{C}_{1}\mathbf{C}_{1}^{\top}+\beta\mathbf{C}_{0}\mathbf{C}_{0}^{\top}+\lambda\mathbf{I}.$ (20)
Since $\lambda>0$ and $\mathbf{I}$ is positive definite, it follows that $\mathbf{M}$ is strictly positive definite and therefore invertible.
Rewriting $e_{0}$ by defining $\mathbf{B}=\mathbf{M}^{-1}\mathbf{C}_{0}$, we obtain
$e_{0}=\|\alpha\mathbf{W}(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top})\mathbf{B}\|^{2}.$ (21)
Applying the singular value bound for matrix products, we have
$\|\mathbf{X}\mathbf{Y}\|\geq\sigma_{\min}(\mathbf{X})\|\mathbf{Y}\|,$ (22)
where $\sigma_{\min}(\mathbf{X})$ is the smallest singular value of $\mathbf{X}$. Applying this inequality, we obtain
$\|\mathbf{W}(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top})\mathbf{B}\|\geq\sigma_{\min}(\mathbf{W})\|(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top})\mathbf{B}\|.$ (23)
We start with the singular value decomposition (SVD) of the matrix $\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}$, given by
$\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\top}.$ (24)
Here, $\mathbf{U}$ and $\mathbf{V}$ are orthogonal matrices, and
$\mathbf{\Sigma}=\operatorname{diag}(\sigma_{1},\sigma_{2},\dots,\sigma_{r},0,\dots,0)$ (25)
is a diagonal matrix containing the singular values $\sigma_{1}\geq\sigma_{2}\geq\dots\geq\sigma_{r}>0$, followed by zeros.
Multiplying both sides by $\mathbf{B}$, we obtain
$(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top})\mathbf{B}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\top}\mathbf{B}.$ (26)
Define the projection of $\mathbf{B}$ onto the subspace spanned by the first $r$ right singular vectors (corresponding to the nonzero singular values) as
$\mathbf{B}_{\text{proj}}=\mathbf{V}_{r}^{\top}\mathbf{B},$ (27)
where $\mathbf{V}_{r}$ consists of the first $r$ columns of $\mathbf{V}$. Then, we can rewrite the expression as
$(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top})\mathbf{B}=\mathbf{U}\mathbf{\Sigma}\mathbf{B}_{\text{proj}}.$ (28)
Taking norms on both sides and using the fact that orthogonal transformations preserve norms, we get
$\|(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top})\mathbf{B}\|=\|\mathbf{\Sigma}\mathbf{B}_{\text{proj}}\|.$ (29)
Since $\mathbf{\Sigma}$ is a diagonal matrix, its smallest nonzero singular value $\sigma_{r}$ provides a lower bound:
$\|\mathbf{\Sigma}\mathbf{B}_{\text{proj}}\|\geq\sigma_{r}\|\mathbf{B}_{\text{proj}}\|.$ (30)
Next, we establish a lower bound for $\|\mathbf{B}_{\text{proj}}\|$. Since $\mathbf{B}$ is not entirely in the null space of $\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}$ (by assumption that the matrix is nonzero), there exists a constant $c_{1}>0$ such that
$\|\mathbf{B}_{\text{proj}}\|\geq c_{1}\|\mathbf{B}\|.$ (31)
Combining these inequalities, we obtain
$\|(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top})\mathbf{B}\|\geq\sigma_{r}\|\mathbf{B}_{\text{proj}}\|\geq\sigma_{r}c_{1}\|\mathbf{B}\|.$ (32)
Since $\mathbf{M}$ is positive definite, we use the standard norm inequality for an invertible matrix $\mathbf{M}$, which states that for any matrix $\mathbf{X}$,
$\|\mathbf{M}\mathbf{X}\|\leq\|\mathbf{M}\|\|\mathbf{X}\|.$ (33)
Setting $\mathbf{X}=\mathbf{M}^{-1}\mathbf{C}_{0}$, we obtain
$\|\mathbf{M}\mathbf{M}^{-1}\mathbf{C}_{0}\|\leq\|\mathbf{M}\|\|\mathbf{M}^{-1}\mathbf{C}_{0}\|.$ (34)
Since $\mathbf{M}\mathbf{M}^{-1}=\mathbf{I}$, the left-hand side simplifies to $\|\mathbf{C}_{0}\|$, yielding
$\|\mathbf{C}_{0}\|\leq\|\mathbf{M}\|\|\mathbf{M}^{-1}\mathbf{C}_{0}\|.$ (35)
Dividing both sides by $\|\mathbf{M}\|$, we obtain
$\|\mathbf{M}^{-1}\mathbf{C}_{0}\|\geq\frac{1}{\|\mathbf{M}\|}\|\mathbf{C}_{0}\|.$ (36)
Thus, it follows that
$\|\mathbf{B}\|=\|\mathbf{M}^{-1}\mathbf{C}_{0}\|\geq\frac{1}{\|\mathbf{M}\|}\|\mathbf{C}_{0}\|.$ (37)
Combining the above results, we obtain
$\|\mathbf{W}(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top})\mathbf{B}\|\geq\sigma_{\min}(\mathbf{W})\sigma_{r}c_{1}\frac{1}{\|\mathbf{M}\|}\|\mathbf{C}_{0}\|.$ (38)
Squaring both sides, we conclude that
$e_{0}=\|\alpha\mathbf{W}(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top})\mathbf{B}\|^{2}\geq\alpha^{2}\sigma_{\min}^{2}(\mathbf{W})\sigma_{r}^{2}c_{1}^{2}\frac{1}{\|\mathbf{M}\|^{2}}\|\mathbf{C}_{0}\|^{2}.$ (39)
Since all terms on the right-hand side are strictly positive by assumption, we establish the existence of a positive lower bound $c>0$ such that
$e_{0}\geq c>0.$ (40)
This completes the proof. ∎
B.3 Deriving the Closed-Form Solution for Ours
From Eq. 9, we are tasked with minimizing the following editing objective:
$\mathop{\text{min}}\limits_{\boldsymbol{\Delta}}\|(\mathbf{W}+\boldsymbol{\Delta}\mathbf{P})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}+\|\boldsymbol{\Delta}\mathbf{P}\|^{2},\quad\text{s.t.}\,(\boldsymbol{\Delta}\mathbf{P})\mathbf{C}_{2}=\mathbf{0}.$ (41)
This is a weighted least squares problem subject to an equality constraint. To solve it, we first formulate the Lagrangian function, where $\boldsymbol{\Lambda}$ is the Lagrange multiplier:
$\mathcal{L}(\boldsymbol{\Delta},\boldsymbol{\Lambda})=\|(\mathbf{W}+\boldsymbol{\Delta}\mathbf{P})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}+\|\boldsymbol{\Delta}\mathbf{P}\|^{2}+\boldsymbol{\Lambda}^{\top}\left((\boldsymbol{\Delta}\mathbf{P})\mathbf{C}_{2}\right).$ (42)
We compute the gradient of the Lagrangian function in Eq. 42 with respect to $\boldsymbol{\Delta}$ and set it to zero, yielding the following equation for $\boldsymbol{\Delta}$:
$\frac{\partial\mathcal{L}(\boldsymbol{\Delta},\boldsymbol{\Lambda})}{\partial\boldsymbol{\Delta}}=2\left((\mathbf{W}+\boldsymbol{\Delta}\mathbf{P})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\right)\mathbf{C}_{1}^{\top}\mathbf{P}^{\top}+2\boldsymbol{\Delta}\mathbf{P}\mathbf{P}^{\top}+\boldsymbol{\Lambda}\mathbf{C}_{2}^{\top}\mathbf{P}^{\top}=\mathbf{0}.$ (43)
Given that the projection matrix $\mathbf{P}$ is derived from $R_{\text{refine}}$ using Eq. 2, $\mathbf{P}$ is a symmetric matrix (i.e., $\mathbf{P}=\mathbf{P}^{\top}$) and an idempotent matrix (i.e., $\mathbf{P}^{2}=\mathbf{P}$), the above formulation can be simplified to:
$\frac{\partial\mathcal{L}(\boldsymbol{\Delta},\boldsymbol{\Lambda})}{\partial\boldsymbol{\Delta}}=2\left((\mathbf{W}+\boldsymbol{\Delta}\mathbf{P})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\right)\mathbf{C}_{1}^{\top}\mathbf{P}+2\boldsymbol{\Delta}\mathbf{P}+\boldsymbol{\Lambda}\mathbf{C}_{2}^{\top}\mathbf{P}=\mathbf{0}.$ (44)
Therefore, we can obtain the closed-form solution for $\boldsymbol{\Delta}\mathbf{P}$ from this equation:
$\boldsymbol{\Delta}\mathbf{P}=(\mathbf{W}\mathbf{C}_{*}\mathbf{C}_{1}^{\top}\mathbf{P}-\mathbf{W}\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\mathbf{P}-\frac{1}{2}\boldsymbol{\Lambda}\mathbf{C}_{2}^{\top}\mathbf{P})(\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\mathbf{P}+\mathbf{I})^{-1}.$ (45)
Next, we differentiate the Lagrangian function in Eq. 42 with respect to $\boldsymbol{\Lambda}$ and set it to zero:
$\frac{\partial\mathcal{L}(\boldsymbol{\Delta},\boldsymbol{\Lambda})}{\partial\boldsymbol{\Lambda}}=(\boldsymbol{\Delta}\mathbf{P})\mathbf{C}_{2}=\mathbf{0}.$ (46)
For simplicity, we define $\mathbf{M}=(\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\mathbf{P}+\mathbf{I})^{-1}$. Then, we substitute the result of Eq.45 into Eq.46 and obtain:
$(\mathbf{W}\mathbf{C}_{*}\mathbf{C}_{1}^{\top}\mathbf{P}-\mathbf{W}\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\mathbf{P}-\frac{1}{2}\boldsymbol{\Lambda}\mathbf{C}_{2}^{\top}\mathbf{P})\mathbf{M}\mathbf{C}_{2}=\mathbf{0}.$ (47)
Solving this equation leads to:
$\frac{1}{2}\boldsymbol{\Lambda}=\mathbf{W}(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top})\mathbf{P}\mathbf{M}\mathbf{C}_{2}(\mathbf{C}_{2}^{\top}\mathbf{P}\mathbf{M}\mathbf{C}_{2})^{-1}.$ (48)
Substituting Eq.48 back into Eq.45, we have the closed-form solution of our objective:
$(\boldsymbol{\Delta}\mathbf{P})_{\text{SPEED}}=\mathbf{W}(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top})\mathbf{P}\mathbf{Q}\mathbf{M},$ (49)
where $\mathbf{Q}=\mathbf{I}-\mathbf{M}\mathbf{C}_{2}(\mathbf{C}_{2}^{\top}\mathbf{P}\mathbf{M}\mathbf{C}_{2})^{-1}\mathbf{C}_{2}^{\top}\mathbf{P}$ and $\mathbf{M}=(\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\mathbf{P}+\mathbf{I})^{-1}$.
Appendix C Implementation Details
Group Number Anchor Concept Celebrity
Erasure Set 10 ‘person’ ‘Adam Driver’, ‘Adriana Lima’, ‘Amber Heard’, ‘Amy Adams’, ‘Andrew Garfield’, ‘Angelina Jolie’, ‘Anjelica Huston’, ‘Anna Faris’, ‘Anna Kendrick’, ‘Anne Hathaway’
50 ‘person’ ‘Adam Driver’, ‘Adriana Lima’, ‘Amber Heard’, ‘Amy Adams’, ‘Andrew Garfield’, ‘Angelina Jolie’, ‘Anjelica Huston’, ‘Anna Faris’, ‘Anna Kendrick’, ‘Anne Hathaway’, ‘Arnold Schwarzenegger’, ‘Barack Obama’, ‘Beth Behrs’, ‘Bill Clinton’, ‘Bob Dylan’, ‘Bob Marley’, ‘Bradley Cooper’, ‘Bruce Willis’, ‘Bryan Cranston’, ‘Cameron Diaz’, ‘Channing Tatum’, ‘Charlie Sheen’, ‘Charlize Theron’, ‘Chris Evans’, ‘Chris Hemsworth’, ‘Chris Pine’, ‘Chuck Norris’, ‘Courteney Cox’, ‘Demi Lovato’, ‘Drake’, ‘Drew Barrymore’, ‘Dwayne Johnson’, ‘Ed Sheeran’, ‘Elon Musk’, ‘Elvis Presley’, ‘Emma Stone’, ‘Frida Kahlo’, ‘George Clooney’, ‘Glenn Close’, ‘Gwyneth Paltrow’, ‘Harrison Ford’, ‘Hillary Clinton’, ‘Hugh Jackman’, ‘Idris Elba’, ‘Jake Gyllenhaal’, ‘James Franco’, ‘Jared Leto’, ‘Jason Momoa’, ‘Jennifer Aniston’, ‘Jennifer Lawrence’
100 ‘person’ ‘Adam Driver’, ‘Adriana Lima’, ‘Amber Heard’, ‘Amy Adams’, ‘Andrew Garfield’, ‘Angelina Jolie’, ‘Anjelica Huston’, ‘Anna Faris’, ‘Anna Kendrick’, ‘Anne Hathaway’, ‘Arnold Schwarzenegger’, ‘Barack Obama’, ‘Beth Behrs’, ‘Bill Clinton’, ‘Bob Dylan’, ‘Bob Marley’, ‘Bradley Cooper’, ‘Bruce Willis’, ‘Bryan Cranston’, ‘Cameron Diaz’, ‘Channing Tatum’, ‘Charlie Sheen’, ‘Charlize Theron’, ‘Chris Evans’, ‘Chris Hemsworth’, ‘Chris Pine’, ‘Chuck Norris’, ‘Courteney Cox’, ‘Demi Lovato’, ‘Drake’, ‘Drew Barrymore’, ‘Dwayne Johnson’, ‘Ed Sheeran’, ‘Elon Musk’, ‘Elvis Presley’, ‘Emma Stone’, ‘Frida Kahlo’, ‘George Clooney’, ‘Glenn Close’, ‘Gwyneth Paltrow’, ‘Harrison Ford’, ‘Hillary Clinton’, ‘Hugh Jackman’, ‘Idris Elba’, ‘Jake Gyllenhaal’, ‘James Franco’, ‘Jared Leto’, ‘Jason Momoa’, ‘Jennifer Aniston’, ‘Jennifer Lawrence’, ‘Jennifer Lopez’, ‘Jeremy Renner’, ‘Jessica Biel’, ‘Jessica Chastain’, ‘John Oliver’, ‘John Wayne’, ‘Johnny Depp’, ‘Julianne Hough’, ‘Justin Timberlake’, ‘Kate Bosworth’, ‘Kate Winslet’, ‘Leonardo Dicaprio’, ‘Margot Robbie’, ‘Mariah Carey’, ‘Melania Trump’, ‘Meryl Streep’, ‘Mick Jagger’, ‘Mila Kunis’, ‘Milla Jovovich’, ‘Morgan Freeman’, ‘Nick Jonas’, ‘Nicolas Cage’, ‘Nicole Kidman’, ‘Octavia Spencer’, ‘Olivia Wilde’, ‘Oprah Winfrey’, ‘Paul Mccartney’, ‘Paul Walker’, ‘Peter Dinklage’, ‘Philip Seymour Hoffman’, ‘Reese Witherspoon’, ‘Richard Gere’, ‘Ricky Gervais’, ‘Rihanna’, ‘Robin Williams’, ‘Ronald Reagan’, ‘Ryan Gosling’, ‘Ryan Reynolds’, ‘Shia Labeouf’, ‘Shirley Temple’, ‘Spike Lee’, ‘Stan Lee’, ‘Theresa May’, ‘Tom Cruise’, ‘Tom Hanks’, ‘Tom Hardy’, ‘Tom Hiddleston’, ‘Whoopi Goldberg’, ‘Zac Efron’, ‘Zayn Malik’
Retain Set 10, 50, and 100 - ‘Aaron Paul’, ‘Alec Baldwin’, ‘Amanda Seyfried’, ‘Amy Poehler’, ‘Amy Schumer’, ‘Amy Winehouse’, ‘Andy Samberg’, ‘Aretha Franklin’, ‘Avril Lavigne’, ‘Aziz Ansari’, ‘Barry Manilow’, ‘Ben Affleck’, ‘Ben Stiller’, ‘Benicio Del Toro’, ‘Bette Midler’, ‘Betty White’, ‘Bill Murray’, ‘Bill Nye’, ‘Britney Spears’, ‘Brittany Snow’, ‘Bruce Lee’, ‘Burt Reynolds’, ‘Charles Manson’, ‘Christie Brinkley’, ‘Christina Hendricks’, ‘Clint Eastwood’, ‘Countess Vaughn’, ‘Dakota Johnson’, ‘Dane Dehaan’, ‘David Bowie’, ‘David Tennant’, ‘Denise Richards’, ‘Doris Day’, ‘Dr Dre’, ‘Elizabeth Taylor’, ‘Emma Roberts’, ‘Fred Rogers’, ‘Gal Gadot’, ‘George Bush’, ‘George Takei’, ‘Gillian Anderson’, ‘Gordon Ramsey’, ‘Halle Berry’, ‘Harry Dean Stanton’, ‘Harry Styles’, ‘Hayley Atwell’, ‘Heath Ledger’, ‘Henry Cavill’, ‘Jackie Chan’, ‘Jada Pinkett Smith’, ‘James Garner’, ‘Jason Statham’, ‘Jeff Bridges’, ‘Jennifer Connelly’, ‘Jensen Ackles’, ‘Jim Morrison’, ‘Jimmy Carter’, ‘Joan Rivers’, ‘John Lennon’, ‘Johnny Cash’, ‘Jon Hamm’, ‘Judy Garland’, ‘Julianne Moore’, ‘Justin Bieber’, ‘Kaley Cuoco’, ‘Kate Upton’, ‘Keanu Reeves’, ‘Kim Jong Un’, ‘Kirsten Dunst’, ‘Kristen Stewart’, ‘Krysten Ritter’, ‘Lana Del Rey’, ‘Leslie Jones’, ‘Lily Collins’, ‘Lindsay Lohan’, ‘Liv Tyler’, ‘Lizzy Caplan’, ‘Maggie Gyllenhaal’, ‘Matt Damon’, ‘Matt Smith’, ‘Matthew Mcconaughey’, ‘Maya Angelou’, ‘Megan Fox’, ‘Mel Gibson’, ‘Melanie Griffith’, ‘Michael Cera’, ‘Michael Ealy’, ‘Natalie Portman’, ‘Neil Degrasse Tyson’, ‘Niall Horan’, ‘Patrick Stewart’, ‘Paul Rudd’, ‘Paul Wesley’, ‘Pierce Brosnan’, ‘Prince’, ‘Queen Elizabeth’, ‘Rachel Dratch’, ‘Rachel Mcadams’, ‘Reba Mcentire’, ‘Robert De Niro’
Table 7: The evaluation setup for multi-concept erasure. This celebrity dataset contains an erasure set with 100 celebrities and a retain set with another 100 celebrities. We experiment with erasing 10, 50, and 100 celebrities with the pre-defined target concepts and the entire retain set is utilized in all cases.
C.1 Experimental Setup Details
[…]
Multi-concept erasure. We then compare methods on multi-concept erasure, a more challenging and realistic concept erasure task. Following the experiment setup from [35], we introduce a dataset consisting of 200 celebrities, where their portraits generated by SDv1.4 [1] can be recognizable with exceptional accuracy by the GIPHY Celebrity Detector (GCD) [20]. This dataset is divided into two groups: an erasure set with 10, 50, and 100 celebrities and a retain set with 100 other celebrities. The full list for both sets is presented in Table 7. We experiment with erasing 10, 50, and 100 celebrities with the pre-defined target concepts and the entire retain set is utilized in all cases. In evaluation, we prepare five celebrity templates, (i.e., “a portrait of {Celebrity}”, “a sketch of {Celebrity}”, “an oil painting of {Celebrity}”, “{Celebrity} in an official photo”, and “an image capturing {Celebrity} at a public event”) and generate 500 images for both sets. For non-target concepts, we generate 1 image per template for each of the 100 concepts, totaling 500 images. For target concepts, we adjust the per-concept quantity to maintain a total of 500 images (e.g., erasing 10 celebrities involves generating 10 images with 5 templates).
Implicit concept erasure. We adopt the same setting in [19] to erase nudity $\to$ ‘ ’ as the erasure set and ‘ ’ as the retain set. In evaluation, we generate images using all 4,703 prompts in I2P and use NudeNet [6] to identify nude content with the threshold of 0.6.
Snoopy Mickey Spongebob Pikachu Hello Kitty MS-COCO
CS FID CS FID CS FID CS FID CS FID CS FID
SD v1.4 28.51 - 26.62 - 27.30 - 27.44 - 27.77 - 26.53 -
Erase Snoopy
CS ↓ FID ↑ CS ↑ FID ↓ CS ↑ FID ↓ CS ↑ FID ↓ CS ↑ FID ↓ CS ↑ FID ↓
ConAbl 25.44 98.38 26.63 37.08 26.95 38.92 27.47 26.14 27.65 36.52 26.40 21.20
MACE 20.90 165.74 23.46 105.97 23.35 102.77 26.05 65.71 26.05 75.42 26.09 42.62
RECE 18.38 151.46 26.62 26.63 27.23 34.42 27.47 21.99 27.78 32.35 26.39 25.61
UCE 23.19 102.86 26.64 24.87 27.29 29.86 27.47 19.06 27.75 27.86 26.46 22.18
Ours 23.50 108.51 26.67 23.41 27.31 24.64 27.48 16.81 27.82 21.74 26.48 19.95
Erase Snoopy and Mickey
CS ↓ FID ↑ CS ↓ FID ↑ CS ↑ FID ↓ CS ↑ FID ↓ CS ↑ FID ↓ CS ↑ FID ↓
ConAbl 25.26 106.78 26.58 57.05 26.81 45.08 27.34 35.57 27.74 41.48 26.42 24.34
MACE 20.53 170.01 20.63 142.98 22.03 112.01 24.98 91.72 23.64 106.88 25.50 55.15
RECE 18.57 150.84 19.14 145.59 27.29 35.85 27.37 26.05 27.71 40.77 26.31 30.30
UCE 23.60 99.30 24.79 86.32 27.32 30.58 27.38 23.51 27.74 31.76 26.38 26.06
Ours 23.58 103.62 23.62 83.70 27.34 29.67 27.39 22.51 27.78 28.23 26.47 23.66
Erase Snoopy and Mickey and Spongebob
CS ↓ FID ↑ CS ↓ FID ↑ CS ↓ FID ↑ CS ↑ FID ↓ CS ↑ FID ↓ CS ↑ FID ↓
ConAbl 24.92 112.66 26.46 63.95 25.12 102.68 27.36 46.47 27.72 48.24 26.37 26.71
MACE 19.86 175.43 19.35 140.13 20.12 143.17 19.76 110.12 21.03 128.56 23.39 66.39
RECE 18.17 155.26 18.87 149.77 16.23 178.55 27.34 40.52 27.71 52.06 26.32 32.51
UCE 23.29 101.40 24.63 88.11 19.08 140.40 27.45 29.20 27.82 38.15 26.30 28.71
Ours 23.69 103.33 23.93 86.55 21.39 109.28 27.47 21.40 27.76 26.22 26.51 24.99
Table 8: Full quantitative comparison of the few-concept erasure in erasing instances from Table 1. The erased target concept(s) are shaded in pink, the best results are highlighted in bold, and the grey columns are indirect indicators for measuring erasure efficacy on target concepts or prior preservation on non-target concepts.
Van Gogh Picasso Monet Paul Gauguin Caravaggio MS-COCO
CS FID CS FID CS FID CS FID CS FID CS FID
SD v1.4 28.75 - 27.98 - 28.91 - 29.80 - 26.27 - 26.53 -
Erase Van Gogh
CS ↓ FID ↑ CS ↑ FID ↓ CS ↑ FID ↓ CS ↑ FID ↓ CS ↑ FID ↓ CS ↑ FID ↓
ConAbl 28.16 129.57 27.07 77.01 28.44 63.80 29.49 63.20 26.15 79.25 26.46 18.36
MACE 26.66 169.60 27.39 69.92 28.84 60.88 29.39 56.18 26.19 69.04 26.50 23.15
RECE 26.39 171.70 27.58 60.57 28.83 61.09 29.58 47.07 26.21 72.85 26.52 23.54
UCE 28.10 133.87 27.70 43.02 28.92 40.49 29.62 32.62 26.23 61.72 26.54 19.63
Ours 26.29 131.02 27.96 35.86 28.94 16.85 29.71 24.94 26.24 39.75 26.55 20.36
Erase Picasso
CS ↑ FID ↓ CS ↓ FID ↑ CS ↑ FID ↓ CS ↑ FID ↓ CS ↑ FID ↓ CS ↑ FID ↓
ConAbl 28.66 60.44 26.97 131.45 28.72 36.23 29.68 65.23 26.20 79.12 26.43 20.02
MACE 28.68 59.58 26.48 137.09 28.73 37.02 29.71 46.35 26.23 66.20 26.47 22.86
RECE 28.71 51.09 26.66 126.40 28.87 25.39 29.69 46.08 26.22 75.61 26.48 23.03
UCE 28.72 37.58 26.99 102.21 28.92 16.72 29.71 32.48 26.22 59.27 26.50 20.33
Ours 28.76 19.18 26.22 117.71 28.88 19.87 29.75 24.73 26.24 43.63 26.51 19.98
Erase Monet
CS ↑ FID ↓ CS ↑ FID ↓ CS ↓ FID ↑ CS ↑ FID ↓ CS ↑ FID ↓ CS ↑ FID ↓
ConAbl 28.58 68.77 27.43 64.25 27.05 96.67 29.09 57.33 26.09 71.88 26.45 21.03
MACE 28.56 61.50 27.74 48.41 25.98 116.34 29.39 49.66 25.98 65.87 26.47 22.76
RECE 28.63 56.26 27.88 45.97 25.87 121.28 29.43 46.38 26.20 64.19 26.49 24.94
UCE 28.65 42.25 27.91 38.73 27.12 98.37 29.58 33.00 26.16 56.49 26.51 21.58
Ours 28.76 28.78 27.93 41.21 25.06 134.11 29.66 27.85 26.22 55.20 26.48 20.87
Table 9: Full quantitative comparison of the few-concept erasure in erasing styles from Table 2. The erased target concept(s) are shaded in pink, the best results are highlighted in bold, and the grey columns are indirect indicators for measuring erasure efficacy on target concepts or prior preservation on non-target concepts.
Concept Snoopy Mickey Spongebob Pikachu Hello Kitty
CS FID CS FID CS FID CS FID CS FID
SD v1.4 28.51 - 26.62 - 27.30 - 27.44 - 27.77 -
Erase Snoopy
CS ↓ FID ↑ CS ↑ FID ↓ CS ↑ FID ↓ CS ↑ FID ↓ CS ↑ FID ↓
SPM 23.72 116.26 26.62 31.21 27.21 31.96 27.41 19.82 27.80 30.95
SPM w/o FT 23.72 116.26 26.55 43.03 26.84 42.96 27.38 25.95 27.71 42.53
Ours 23.50 108.51 26.67 23.41 27.31 24.64 27.42 16.81 27.82 21.74
Erase Snoopy and Mickey
CS ↓ FID ↑ CS ↓ FID ↑ CS ↑ FID ↓ CS ↑ FID ↓ CS ↑ FID ↓
SPM 23.18 122.17 22.71 117.30 26.92 38.35 27.35 27.13 27.76 39.61
SPM w/o FT 26.66 83.11 25.64 75.46 26.73 42.30 27.50 28.19 27.75 41.19
Ours 23.58 103.62 23.62 83.70 27.24 29.67 27.39 22.51 27.78 28.23
Erase Snoopy and Mickey and Spongebob
CS ↓ FID ↑ CS ↓ FID ↑ CS ↓ FID ↑ CS ↑ FID ↓ CS ↑ FID ↓
SPM 22.86 125.66 22.08 123.20 20.92 153.36 27.50 37.51 27.63 46.63
SPM w/o FT 27.75 68.89 26.24 63.30 24.82 77.38 27.44 30.00 27.74 38.89
Ours 23.69 103.33 23.93 86.55 21.39 109.28 27.47 21.40 27.76 26.22
Table 10: Quantitative comparison with SPM and SPM w/o FT (Facilitated Transport). The erased target concept(s) are shaded in pink, the best results are highlighted in bold, and the grey columns are indirect indicators for measuring erasure efficacy on target concepts or prior preservation on non-target concepts.
C.2 Erasure Configurations
Implementation of previous works. In our series of three concept erasure tasks, we mainly compare against four methods: ConAblhttps://github.com/nupurkmr9/concept-ablation [29], MACE§§§https://github.com/Shilin-LU/MACE [35], RECEhttps://github.com/CharlesGong12/RECE [19], and UCEhttps://github.com/rohitgandikota/unified-concept-editing [18], as they achieve SOTA performance across different concept erasure tasks. All the compared methods are implemented using their default configurations from the corresponding official repositories. One exception is that for MACE when erasing 50 celebrities, since it doesn’t provide an official configuration and the preserve weight varies with the number of target celebrities, we set it to $1.2\times 10^{5}$ to ensure a consistent balance between erasure and preservation.
Implementation of SPEED. In line with previous methods [29, 35, 19, 18], we edit the cross-attention (CA) layers within the diffusion model due to their role in text-image alignment [21]. In contrast, we only edit the value matrices in the CA layers, as suggested by [62]. This choice is grounded in the observation that the keys in CA layers typically govern the layout and compositional structure of the attention map, while the values control the content and visual appearance of the images [58]. In the context of concept erasure, our goal is to effectively remove the semantics of the target concept, and we find that only editing the value matrices is sufficient as shown in Fig. 5 and 6 (further ablation comparison is provided in Appx. D.3). The augmentation times $N_{A}$ in Eq. 8 is set to 10 and the augmentation ranks $r$ in Eq. 6 is set to 1 as ablated in Appx. D.3. Meanwhile, given that eigenvalues are rarely strictly zero in practical applications when determining the null space, we select the eigenvectors corresponding to the eigenvalues below $10^{-1}$ on few-concept and implicit concept erasure and $10^{-4}$ on multi-concept erasure following [14]. Moreover, since the retain set only includes ‘ ’ in implicit concept erasure, we add an identity matrix $\mathbf{I}$ with weight $\lambda=0.5$ to the term $(\mathbf{C}_{2}^{\top}\mathbf{P}\mathbf{M}\mathbf{C}_{2})^{-1}$ in Eq. 11 to ensure invertibility following [18].
Appendix D Additional Experiments
D.1 Full Comparison on Few-Concept Erasure
We present full quantitative comparisons of few-concept erasure, including both CS and FID, in Table 8 and Table 9. Our results demonstrate that our method consistently achieves superior prior preservation, as indicated by higher CS and lower FID across the majority of non-target concepts.
D.2 More Baselines
In this section, we compare against more methods because of the page limit in our main paper. Since our method focuses on improving prior preservation and multi-concept erasure performance, we mainly compare it with similar methods, other methods like ESD [17], FMN [66], and SLD [51] are omitted, as they fail to achieve satisfactory prior preservation proved by previous comaprisons [37, 35, 62]. The remaining comparable method is SPM******https://github.com/Con6924/SPM [37], which is proposed to improve prior preservation and can scale to multi-concept erasure tasks. Notable, SPM not only fine-tunes the model weights using LoRA [25] but also intervenes in the image generation process through Facilitated Transport. Specifically, this module dynamically adjusts the LoRA scale based on the similarity between the sampling prompt and the target concept. In other words, if the prompt contains the target concept or is highly relevant, this scale is set to a large value, whereas if there is little to no relevance, it is set close to 0, functioning similarly to a text filter. We argue that such a comparison with SPM is not fair since we only focus on modifying the model parameters, and therefore, we compare both the original SPM and SPM without Facilitated Transport (SPM w/o FT) for a fair comparison. In the latter version, the LoRA scale is set to 1 by default.
The quantitative comparative results are shown in Table 10. It can be seen that our method consistently achieves the best prior preservation compared to both SPM and SPM w/o FT. Even equipped with Facilitated Transport, our method achieves the lowest non-target FID (e.g., on Pikachu and Hello Kitty). This superiority amplifies as the number of target concepts increases. For example, with the number of target concepts increasing from 1 to 3, our FID in Pikachu rises from 16.81 to 21.40 (4.59 ↑), while SPM increases from 19.82 to 37.51 (17.69 ↑), where a similar pattern is observed in Hello Kitty (Our 4.48 ↑ v.s. SPM’s 15.68 ↑). Once removing the Facilitated Transport module, SPM w/o FT shows poor erasure efficacy in multi-concept erasure (highlighted by red in Table 10 and illustrated in Fig. 8). This indicates that the success of SPM in multi-concept erasure relies on the Facilitated Transport module, which dynamically allocates the LoRA scales by calculating the similarity between the sampling prompt and each target concept. For example, when erasing Snoopy + Mickey + SpongeBob, if the sampling prompt is “a photo of Snoopy”, SPM will allocate a larger scale to Snoopy’s LoRA according to the text similarity. On the contrary, if the sampling prompt is “a photo of Pikachu” with the non-target concept, all three LoRA scales will be assigned lower values, thereby preserving the prior knowledge. We argue that this strategy of dynamically tuning the LoRA scales based on the sampling prompt similarity is vulnerable to attacks and easily bypassed, especially in white-box attack scenarios, where an attacker can reconstruct the erased concepts by simply modifying the code with extremely low attack costs, e.g., open-source T2I models like Stable Diffusion [1, 2].
Figure 8: Qualitative comparison with SPM and SPM w/o FT in erasing single and multiple instances. The erased and preserved generations are highlighted with red and green boxes, respectively. Our method demonstrates superior prior preservation compared to both SPM and SPM w/o FT. Meanwhile, without the Facilitated Transport module, SPM w/o FT shows poor erasure efficacy in multi-concept erasure (e.g., the 1st case in Snoopy, 2nd case in erasing Mickey, and 3rd case in SpongeBob) with relatively high target CS in Table 10.
D.3 Ablation Studies
Augmentation times. We ablate the augmentation times $N_{A}$ proposed in the Directed Prior Augmentation (DPA) module in Sec. 4.2, which controls the balance between rank saturation and prior coverage along with the Influence-based Prior Filtering (IPF) module. It can be observed from Fig. 9 (a) that: (1) As $N_{A}$ increases, the non-target FID exhibits a trend of first decreasing and then increasing. This suggests that when $N_{A}$ is small (i.e., $1\to 10$), augmenting existing non-target concepts with semantically similar concepts facilitates a more comprehensive prior coverage, thereby improving prior preservation. However, when $N_{A}$ exceeds a certain threshold (i.e., $10\to 20$), further augmentation of non-target concepts leads to narrowing the null-space derivation due to rank saturation, ultimately degrading prior preservation. (2) Target CS generally shows a declining trend, indicating that the proposed Prior Knowledge Refinement strategy not only improves prior preservation but also exerts a positive impact on erasure efficacy.
Augmentation ranks. Another hyperparameter to be ablated is the augmentation ranks $r$. From Eq. 6, we introduce the number of the smallest singular values, i.e., augmentation ranks $r$ in deriving $\mathbf{P}_{\text{min}}=\mathbf{U}_{\text{min}}\mathbf{U}_{\text{min}}^{\top}$ with $\mathbf{U}_{\text{min}}={\mathbf{U}}_{\mathbf{W}}[:,-r:]$. Mathematically, $r$ represents the directions in which the DPA module can augment in the concept embedding space and constrains the rank of the augmented embeddings to a maximum of $r$. As shown in Fig. 9 (b), as $r$ increases, the non-target FID exhibits an overall upward trend, indicating that introducing more ranks does not benefit prior preservation, as it narrows the null space. At the same time, as shown in Table 6, such augmentation by DPA also remains necessary, as it enables more comprehensive coverage of non-target knowledge with semantically similar concepts, leading to improved prior preservation.
Figure 9: Ablation study on two parameters, i.e., augmentation times $N_{A}$ and augmentation ranks $r$ of the DPA module. We report the target CS of erasing Van Gogh and the non-target FID averaged over other four styles (i.e., Picasso, Monet, Paul Gauguin, Caravaggio).
Ablation Parameters Van Gogh Picasso Monet Paul Gauguin Caravaggio MS-COCO
Key Value CS ↓ FID ↓ FID ↓ FID ↓ FID ↓ CS ↑ FID ↓
1 $\times$ 27.67 42.11 26.09 28.08 52.44 26.55 18.72
2 26.24 48.41 28.65 33.79 57.23 26.53 23.20
Ours $\times$ 26.29 35.86 16.85 24.94 39.75 26.55 20.36
Table 11: Ablation study on the edited parameters. Our scheme on only editing the value matrices achieves a superior balance between erasure efficacy (e.g., target CS of 26.29) and prior preservation (e.g., the lowest FIDs across all non-target concepts).
Metirc Van Gogh Picasso Monet Paul Gauguin Caravaggio MS-COCO
CS ↓ FID ↓ FID ↓ FID ↓ FID ↓ CS ↑ FID ↓
Text Similarity 26.35 36.87 19.69 25.18 41.44 26.52 20.78
Prior Shift (Ours) 26.29 35.86 16.85 24.94 39.75 26.55 20.36
Table 12: Ablation study on the importance metrics in IPF.
Edited parameters. We compare the impact on editing different CA parameters in Table 11 and draw the following conclusions: (1) Only editing the key matrices cannot achieve effective erasure, with the target CS being 27.67 (v.s. the original CS of 28.75). This is because they mainly arrange the layout information of the generation and cannot effectively erase the semantics of the target concept. (2) Simultaneously editing both the key and value matrices can achieve effective erasure, but it will also excessively damage prior knowledge. (3) Only editing the value matrices achieves a superior balance between erasure efficacy and prior preservation. Compared to Ablation 2, the editing of key matrices leads to excessive erasure, which is unnecessary in concept erasure.
Importance metrics in IPF. In Sec. 4.1, we propose Importance-based Prior Filtering (IPF) in Eq. 5 and evaluate this importance with the metric prior shift $=\|\boldsymbol{\Delta}_{\text{erase}}\boldsymbol{c}\|^{2}$. Another intuitive and plausible metric is based on text similarity, e.g., the cosine similarity between each non-target embedding $\boldsymbol{c}_{0}$ and each target concept embedding $\boldsymbol{c}_{1}$, i.e., $\cos(\boldsymbol{c}_{0},\boldsymbol{c}_{1})$. Herein, we conduct an additional ablation study in terms of the metric selection in Table 12. It can be seen that text similarity can also serve as an effective metric for evaluating importance with improved non-target FID while the prior shift provides better prior preservation. This may be because text similarity is implicitly related to importance, while prior shift explicitly reflects the impact of erasure on different concepts from the model updates $\mathbf{\Delta}$. Moreover, our method can be directly scaled up to multi-concept erasure scenarios, whereas text similarity calculates $n$ similarities for $n$ target concepts, requiring additional fusion or selection strategies, introducing accumulated errors during fusion or selection.
Appendix E Limitation
While SPEED demonstrates superior prior preservation, its erasure efficacy may not be as strong as some adversarial training/editing-based methods (e.g., RECE [19]), which explicitly optimize for robust concept removal. This trade-off arises from SPEED’s emphasis on maintaining non-target knowledge, potentially leading to residual traces of erased concepts in extreme cases. However, due to its efficiency and scalability in multi-concept erasure, an interesting direction for future work is to explore simultaneous erasure of adversarial examples. Given that null-space constraints inherently minimize the impact on prior knowledge, even with the addition of extra target concepts, SPEED is expected to achieve better prior preservation compared to existing methods while effectively handling adversarial concept erasure.

### v2

@@
 Erasing concepts from large-scale text-to-image (T2I) diffusion models has become increasingly crucial due to the growing concerns over copyright infringement, offensive content, and privacy violations.
-However, existing methods either require costly fine-tuning or degrade image quality for non-target concepts (i.e., prior) due to inherent optimization limitations.
-In this paper, we introduce SPEED, a model editing-based concept erasure approach that leverages null-space constraints for scalable, precise, and efficient erasure.
-Specifically, SPEED incorporates Influence-based Prior Filtering (IPF) to retain the most affected non-target concepts during erasing, Directed Prior Augmentation (DPA) to expand prior coverage while maintaining semantic consistency, and Invariant Equality Constraints (IEC) to regularize model editing by explicitly preserving key invariants during the T2I generation process.
-Extensive evaluations across multiple concept erasure tasks demonstrate that SPEED consistently outperforms existing methods in prior preservation while achieving efficient and high-fidelity concept erasure, successfully removing 100 concepts within just 5 seconds.
+In scalable applications, fine-tuning-based methods are time-consuming to precisely erase multiple target concepts, while real-time editing-based methods often degrade the generation quality of non-target concepts due to conflicting optimization objectives.
+To address this dilemma, we introduce SPEED, an efficient concept erasure approach that directly edits model parameters.
+SPEED searches for a null space, a model editing space where parameter updates do not affect non-target concepts, to achieve scalable and precise erasure.
+To facilitate accurate null space optimization, we incorporate three complementary strategies: Influence-based Prior Filtering (IPF) to selectively retain the most affected non-target concepts, Directed Prior Augmentation (DPA) to enrich the filtered retain set with semantically consistent variations, and Invariant Equality Constraints (IEC) to preserve key invariants during the T2I generation process.
+Extensive evaluations across multiple concept erasure tasks demonstrate that SPEED consistently outperforms existing methods in non-target preservation while achieving efficient and high-fidelity concept erasure, successfully erasing 100 concepts within only 5 seconds.
 Our code and models are available at: https://github.com/Ouxiang-Li/SPEED.
-Figure 1: Three characteristics of our proposed method, SPEED.
+Figure 1: Three characteristics of our proposed concept erasure method for diffusion models, SPEED.
 (a) Scalable: SPEED seamlessly scales from single-concept to large-scale multi-concept erasure (e.g., 100 celebrities) without additional design.
-(b) Precise: SPEED precisely removes the target concept (e.g., Snoopy) while preserving the semantic integrity for non-target concepts (e.g., Hello Kitty and SpongeBob).
-(c) Efficient: SPEED can immediately erase 100 concepts within 5 seconds, achieving a ×$\times 350$ speedup over the state-of-the-art (SOTA) method.
-11footnotetext: Equal Contributions.22footnotetext: Corresponding author.
+(b) Precise: SPEED precisely removes the target concept (e.g., Snoopy) while preserving the semantics for non-target concepts (e.g., Hello Kitty and SpongeBob).
+(c) Efficient: SPEED immediately erases 100 concepts within 5 seconds, achieving new state-of-the-art (SOTA) performance with a $350\times$ speedup over competitive methods.
 1 Introduction
-Large-scale text-to-image (T2I) diffusion models [24, 56, 57, 39, 49, 23] have facilitated significant breakthroughs in generating highly realistic and contextually consistent images simply from textual descriptions [11, 46, 16, 7, 50, 44, 13].
-Alongside these advancements, concerns have also been raised regarding copyright violations [10, 54, 30], offensive content [51, 65, 67], and privacy concerns [8, 64, 31].
-To mitigate ethical and legal risks in generation, it is often necessary to prevent the model from generating certain concepts, a process termed concept erasure [29, 17, 66].
-However, removing target concepts without carefully preserving the semantics of non-target concepts can introduce unintended artifacts, distortions, and degraded image quality  [17, 42, 51, 66], compromising the model’s reliability and usability.
+Text-to-image (T2I) diffusion models Ho et al.
+(2020); Song et al.
+(2020a; b); Nichol & Dhariwal (2021); Rombach et al.
+(2022); Ho & Salimans (2022) have facilitated significant breakthroughs in generating highly realistic and contextually consistent images simply from textual descriptions Dhariwal & Nichol (2021); Ramesh et al.
+(2021); Gal et al.
+(2022); Betker et al.
+(2023); Ruiz et al.
+(2023); Podell et al.
+(2023); Esser et al.
+(2024).
+Alongside these advancements, concerns have also been raised regarding copyright violations Cui et al.
+(2023); Shan et al.
+(2023), offensive content Schramowski et al.
+(2023); Yang et al.
+(2024b); Zhang et al.
+(2025), and privacy concerns Carlini et al.
+(2023); Yang et al.
+(2023).
+To mitigate ethical and legal risks in generation, it is often necessary to prevent the model from generating certain concepts, a process termed concept erasure Kumari et al.
+(2023); Gandikota et al.
+(2023); Zhang et al.
+(2024a).
+However, removing target concepts without carefully preserving the semantics of non-target concepts can introduce unintended artifacts, distortions, and degraded image quality Gandikota et al.
+(2023); Orgad et al.
+(2023); Schramowski et al.
+(2023); Zhang et al.
+(2024a), compromising the model’s usability.
 Therefore, beyond ensuring the effective removal of target concepts (i.e., erasure efficacy), concept erasure should also maintain the original semantics of non-target concepts (i.e., prior preservation).
-In this context, recent methods seek to balance between erasure efficacy and prior preservation, broadly categorized into two paradigms: training-based [29, 37, 35] and editing-based [18, 19].
-The training-based paradigm fine-tunes T2I diffusion models to achieve concept erasure, incorporating an additional regularization term in training for prior preservation.
-In contrast, the editing-based paradigm is training-free.
-It jointly optimizes erasure and preservation objectives for target and non-target concepts in a closed-form solution, obtaining direct modifications over model parameters (e.g., projection weights in cross-attention layers [49]).
-This efficiency also allows editing-based methods to seamlessly extend to multi-concept erasure without additional designs.
-However, as the number of target concepts increases, current editing-based methods struggle to balance between erasure efficacy and prior preservation.
-It can be attributed to the growing conflicts among erasure and preservation objectives, making such trade-offs increasingly difficult.
-Moreover, these methods rely on weighted least squares optimization [18, 19], inherently imposing a non-zero lower bound on preservation error (Appx.
+In this context, recent methods strive to seek a balance between erasure efficacy and prior preservation, broadly categorized into two paradigms: training-based Kumari et al.
+(2023); Lyu et al.
+(2024); Lu et al.
+(2024a) and editing-based Gandikota et al.
+(2024); Gong et al.
+(2025).
+The training-based paradigm fine-tunes diffusion models to achieve concept erasure, incorporating an additional regularization into the training objective for prior preservation.
+In contrast, the editing-based paradigm avoids additional fine-tuning by directly modifying model parameters (e.g., projection weights in cross-attention layers Rombach et al.
+(2022)), with such modifications derived from a closed-form objective that jointly accounts for erasure and preservation.
+This efficiency also facilitates editing-based methods to extend to multi-concept erasure without additional designs seamlessly.
+However, as the number of target concepts increases, current editing-based methods Gandikota et al.
+(2024); Gong et al.
+(2025) struggle to balance between erasure efficacy and prior preservation.
+This can be attributed to the growing conflicts between erasure and preservation objectives, making such trade-offs increasingly difficult.
+Moreover, these methods rely on weighted least squares optimization, inherently imposing a non-zero lower bound on preservation error (see Appx.
 B.2).
-In multi-concept settings, this accumulation of preservation errors distorts non-target knowledge, further degrading generative quality.
+In multi-concept settings, this accumulation of preservation errors gradually distorts non-target knowledge, thereby degrading prior preservation.
 To address the above limitations, we propose Scalable, Precise, and Efficient Concept Erasure for Diffusion Models (SPEED) (see Fig.
 1), an editing-based method incorporating null-space constraints.
-Here, we define “null-space” as the subspace that does not alter the feature representations of non-target concepts.
-By projecting the model parameter updates for erasing concepts onto the null space of non-target concepts, SPEED can minimize the preservation error close to zero without compromising erasure efficacy.
-A key challenge for SPEED lies in constructing an effective null space based on a set of non-target concepts (i.e., retain set).
-We observe that the existing baseline with null-space constraints [14] confronts a fundamental dilemma during concept erasure: While a narrow retain set limits the prior coverage, increasing the diversity of the retain set restricts the determination of the null space due to rank saturation, thereby compromising prior preservation (see Fig.
-2).
-In this light, we introduce Prior Knowledge Refinement, a suite of techniques that alleviates optimization difficulty and broadens prior coverage.
-Particularly, to mitigate optimization difficulty, we propose Influence-based Prior Filtering (IPF), which selectively retains non-target concepts from the retain set that are most affected by the erasure process.
-These concepts serve as anchors for defining the null space, providing meaningful constraints for preservation while maintaining sufficient flexibility for model editing.
-To further prevent excessive prior knowledge reduction, we introduce Directed Prior Augmentation (DPA) to enhance prior coverage, which expands the selected retain set after IPF with augmented samples.
-Instead of simply introducing random noise, we generate semantically consistent variations of preserved concepts, mitigating excessive preservation over meaningless concepts.
-Furthermore, we incorporate Invariants Equality Constraint (IEC), which identifies critical invariants in the generation process and enforces equality regularization to explicitly preserve them.
+Specifically, we search for the null space of prior knowledge, a model editing space where parameter updates do not affect the feature representations of non-target concepts.
+By projecting the model parameter updates for concept erasure onto such null space, SPEED can minimize the preservation error to zero without compromising erasure efficacy, thereby enabling scalable and precise concept erasure without affecting non-target concepts.
+Figure 2: Semantic degradation with increasing non-target concepts in retain set.
+The baseline null-space constrained method Fang et al.
+(2024) preserves non-target semantics given a small retain set Refer to caption.
+However, as the retain set expands, the corresponding matrix approaches higher rank, making null space estimation increasingly inaccurate (see Eq.
+4) with inevitable approximation errors, thereby degrading prior semantics Refer to caption Refer to caption.
+The key contribution of SPEED lies in defining an effective null space from a set of non-target concepts (i.e., retain set).
+We observe that the existing baseline with null-space constraints Fang et al.
+(2024) confronts a fundamental dilemma during concept erasure: While a small retain set limits the coverage of prior knowledge, enlarging the retain set makes it increasingly difficult to identify an accurate null space.
+This difficulty arises because a large retain set causes the corresponding feature matrix to approach full rank, necessitating the estimation of its null space to ensure sufficient degrees of freedom for optimization (i.e., concept erasure).
+However, this estimation inevitably introduces semantic degradation to the retain set and deteriorates prior preservation in Fig.
+2.
+In this light, we introduce Prior Knowledge Refinement, a suite of techniques that strategically and selectively refine the retain set to mitigate the semantic degradation in searching for the null space.
+Particularly, we propose Influence-based Prior Filtering (IPF), which first quantifies the influence of concept erasure on each non-target concept.
+It then prunes the retain set by removing minimally affected concepts, preventing the correlation matrix from approaching full rank and thus maintaining an accurate null space.
+Subsequently, to further enhance prior preservation over the resulting retain set, we propose Directed Prior Augmentation (DPA), which expands the retain set with directed, semantically consistent perturbations to improve retain coverage.
+In addition, we incorporate Invariant Equality Constraints (IEC) to preserve specific representations, such as the [SOT] token, that should remain unchanged during editing.
+IEC enforces equality constraints on such invariants to regularize the retaining of essential generation properties.
 We evaluate SPEED on three representative concept erasure tasks, i.e., few-concept, multi-concept, and implicit concept erasure, where it consistently exhibits superior prior preservation across all erasure tasks.
-Overall, our contributions are summarized as follows:
+Overall, our contributions can be summarized as follows:
   • We propose SPEED, a scalable, precise, and efficient concept erasure method with null-space constrained model editing, capable of erasing 100 concepts in 5 seconds.
-  • We introduce Prior Knowledge Refinement to construct a well-defined null space that remains meaningful and tractable for editing.
-By integrating Influence-based Prior Filtering, Directed Prior Augmentation, and Invariants Equality Constraint, it balances rank saturation and prior coverage effectively with precise concept erasure.
+  • We introduce Prior Knowledge Refinement to construct an accurate null space over the retain set for effective editing.
+Leveraging three complementary techniques, IPF, DPA, and IEC, our method balances semantic degradation and retain coverage, enabling precise and scalable concept erasure.
   • Our extensive experiments show that SPEED consistently outperforms existing methods in prior preservation across various erasure tasks with minimal computational costs.
-Figure 2: Impact of the increasing number of non-target concepts in retain set.
-The baseline [14] maintains prior preservation with a small retain set (e.g., Refer to caption), but as the retain set grows, rank saturation occurs, leading to gradually degraded semantics in non-target generations (e.g., Refer to caption and Refer to caption).
-In contrast, our method achieves consistent prior preservation by optimizing the retain set.
 2 Related Works
 Concept erasure.
-Current T2I diffusion models inevitably involve unauthorized and NSFW (Not Safe For Work) generations due to the noisy training data from web [52, 53].
-Apart from applying additional filters or safety checkers [47, 41, 48], prevailing methods modify diffusion model parameters to erase specific target concepts, mainly categorized into two paradigms.
-The training-based paradigm fine-tunes model parameters with specific erasure objectives [29, 17, 66] and additional regularization terms [29, 37, 35].
-In contrast, the editing-based paradigm edits model parameters using a closed-form solution to facilitate efficiency in concept erasure.
-For example, UCE [18] modifies model weights by balancing both erasure and preservation error through a weighted least squares objective and RECE [19] iteratively derives new target concept embeddings.
+Current T2I diffusion models inevitably involve unauthorized and offensive generations due to the noisy training data from web Schuhmann et al.
+(2021; 2022).
+Apart from applying additional filters or safety checkers Rando et al.
+(2022); Betker et al.
+(2023); Rao (2023), prevailing methods modify diffusion model parameters to erase specific target concepts, mainly categorized into two paradigms.
+The training-based paradigm fine-tunes model parameters with specific erasure objectives Kumari et al.
+(2023); Gandikota et al.
+(2023); Zhang et al.
+(2024a); Zhao et al.
+(2024b); Huang et al.
+(2024); Kim et al.
+(2024); Zhang et al.
+(2024b); Zhao et al.
+(2024a) and additional regularization terms Kumari et al.
+(2023); Lyu et al.
+(2024); Lu et al.
+(2024a).
+In contrast, the editing-based paradigm edits model parameters using a closed-form solution to facilitate efficiency in concept erasure Orgad et al.
+(2023); Gandikota et al.
+(2024); Gong et al.
+(2025).
 These methods can erase numerous concepts within seconds, demonstrating superior efficiency in practice.
+Beyond parameter modification, non-parametric methods (e.g., external modules and sampling interventions) have also been explored Schramowski et al.
+(2023); Wang et al.
+(2024b); Yoon et al.
+(2024); Jain et al.
+(2024); Lee et al.
+(2025b; a), but they are fragile in open-source settings.
 Null-space constraints.
 The null space of a matrix, a fundamental concept in linear algebra, refers to the set of all vectors that the matrix maps to the zero vector.
-The null-space constraints are first applied to continual learning (CL) by projecting gradients onto the null space of uncentered covariances from previous tasks [60].
-Subsequent studies [36, 61, 63, 28, 32] further explore and extend the application of null space in CL.
-In model editing, AlphaEdit [14] restricts model weight updates onto the null space of preserved knowledge, effectively mitigating trade-offs between editing and preservation.
-The null-space constraints also apply to various other tasks, e.g., machine unlearning [9] and MRI reconstruction [15], offering great promise for editing-based concept erasure.
+The null-space constraints are first applied to continual learning by projecting gradients onto the null space of uncentered covariances from previous tasks Wang et al.
+(2021).
+Subsequent studies Lu et al.
+(2024b); Wang et al.
+(2024a); Yang et al.
+(2024a); Kong et al.
+(2022); Lin et al.
+(2022) further explore and extend the application of null space in continual learning.
+In model editing, AlphaEdit Fang et al.
+(2024) restricts model weight updates onto the null space of preserved knowledge, effectively mitigating trade-offs between editing and preservation.
+Null-space constraints also apply to various tasks, e.g., machine unlearning Chen et al.
+(2024), MRI reconstruction Feng et al.
+(2023), and image restoration Wang et al.
+(2022), offering promise for editing-based concept erasure.
 3 Problem Formulation
+In T2I diffusion models, each concept is encoded by a set of text tokens via CLIP Radford et al.
+(2021), which are then aggregated into a single concept embedding $\bm{c}\in\mathbb{R}^{d_{0}}$.
+For concept erasure, there are two sets of concepts: the erasure set $\mathbf{E}$ and the retain set $\mathbf{R}$.
+The erasure set consists of $N_{E}$ target concepts to be removed, denoted as $\mathbf{E}=\{\bm{c}_{1}^{(i)}\}_{i=1}^{N_{E}}$.
+The retain set includes $N_{R}$ non-target concepts that should be preserved during editing, denoted as $\mathbf{R}=\{\bm{c}_{0}^{(j)}\}_{j=1}^{N_{R}}$.
+To enable efficient erasure efficacy for $\mathbf{E}$ and prior preservation for $\mathbf{R}$, we first formulate a closed-form editing objective in Sec.
+3.1, and enhance it with null-space constrained optimization in Sec.
+3.2.
 3.1 Concept Erasure in Closed-Form Solution
-In T2I diffusion models, each concept is encoded by a set of text tokens via CLIP [45], which are then aggregated into a single concept embedding $\boldsymbol{c}\in\mathbb{R}^{d_{0}}$.
-For concept erasure, there are two sets of concepts: the erasure set $\mathbf{E}$ and the retain set $\mathbf{R}$.
-The erasure set consists of $N_{E}$ target concepts to be removed, denoted as $\mathbf{E}=\{\boldsymbol{c}_{1}^{(i)}\}_{i=1}^{N_{E}}$.
-The retain set includes $N_{R}$ non-target concepts that should be preserved during editing, denoted as $\mathbf{R}=\{\boldsymbol{c}_{0}^{(i)}\}_{i=1}^{N_{R}}$.
-To effectively erase each target concept $\boldsymbol{c}_{1}^{(i)}\in\mathbf{E}$ (e.g., Snoopy), it is mapped onto an anchor concept $\boldsymbol{c}_{*}^{(i)}$ that retains general semantics while avoiding the original concept-specific information (e.g., Dog).
-The paired erasure set is denoted as: $\mathbf{\tilde{E}}=\{(\boldsymbol{c}_{1}^{(i)},\boldsymbol{c}_{*}^{(i)})\}_{i=1}^{N_{E}}$.
-For editing-based methods [42, 18, 19], concept embeddings from the erasure set $\mathbf{\tilde{E}}$ and the retain set $\mathbf{R}$ are first organized into three structured matrices: $\mathbf{C}_{1},\mathbf{C}_{*}\in\mathbb{R}^{d_{0}\times N_{E}}$ and $\mathbf{C}_{0}\in\mathbb{R}^{d_{0}\times N_{R}}$, which represents the stacked embeddings of the target, anchor, and non-target concepts, respectively.
-To derive a closed-form solution for concept erasure, existing methods typically optimize a perturbation $\boldsymbol{\Delta}$ to model parameters $\mathbf{W}$, balancing between erasure efficacy and prior preservation.
-For example, UCE [18] formulates concept erasure as a weighted least squares problem:
-$\boldsymbol{\Delta}_{\text{UCE}}=\mathop{\text{arg\ min}}\limits_{\boldsymbol{\Delta}}\underbrace{\|(\mathbf{W}+\boldsymbol{\Delta})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}}_{\text{$e_{1}$}}+\underbrace{\|\boldsymbol{\Delta}\mathbf{C}_{0}\|^{2}}_{\text{$e_{0}$}},$ (1)
-where the erasure error $e_{1}$ ensures that each target concept is mapped onto its corresponding anchor concept and the preservation error $e_{0}$ minimizes the impact on non-target priors.
-This formulation provides a closed-form solution to obtain $\boldsymbol{\Delta}_{\text{UCE}}$ (see Appx.
+To effectively erase each target concept $\bm{c}_{1}^{(i)}\in\mathbf{E}$ (e.g., Snoopy), it is specified to be mapped onto an anchor concept $\bm{c}_{*}^{(i)}$ that shares general semantics (e.g., Dog), termed as an anchor set $\mathbf{A}=\{\bm{c}_{*}^{(i)}\}_{i=1}^{N_{E}}$.
+For editing-based methods Orgad et al.
+(2023); Gandikota et al.
+(2024); Gong et al.
+(2025), concept embeddings from the erasure set $\mathbf{E}$, anchor set $\mathbf{A}$, and retain set $\mathbf{R}$ are first organized into three structured matrices: $\mathbf{C}_{1},\mathbf{C}_{*}\in\mathbb{R}^{d_{0}\times N_{E}}$ and $\mathbf{C}_{0}\in\mathbb{R}^{d_{0}\times N_{R}}$, representing the stacked embeddings of target, anchor, and non-target concepts, respectively.
+To derive a closed-form solution for concept erasure, existing methods typically optimize a perturbation $\bm{\Delta}$ to model parameters $\mathbf{W}$, balancing between erasure efficacy and prior preservation.
+For example, UCE Gandikota et al.
+(2024) formulates concept erasure as a weighted least squares problem:
+$\bm{\Delta}_{\text{UCE}}=\mathop{\text{arg\ min}}\limits_{\bm{\Delta}}\underbrace{\|(\mathbf{W}+\bm{\Delta})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}}_{\text{$e_{1}$}}+\underbrace{\|\bm{\Delta}\mathbf{C}_{0}\|^{2}}_{\text{$e_{0}$}},$ (1)
+where the erasure error $e_{1}$ ensures that each target concept is mapped onto its corresponding anchor concept and the preservation error $e_{0}$ minimizes the impact on non-target concepts, and $\|\cdot\|^{2}$ denotes the sum of the squared elements in the matrix (i.e., Frobenius norm).
+This formulation provides a closed-form solution $\bm{\Delta}_{\text{UCE}}$ (see Appx.
 B.1) for parameter updates, achieving computationally efficient optimization.
-However, as the number of target concepts increases, the accumulated preservation errors $e_{0}$ across multiple erased concepts would amplify the distortion on non-target knowledge, further degrading prior preservation.
-3.2 Null-Space Constraints
-To mitigate the limitation of weighted optimization in prior preservation, SPEED integrates null-space constraints to achieve model editing [14].
-These constraints restrict the solution space of $\boldsymbol{\Delta}$ within the null space of $\mathbf{C}_{0}$ to ensure that parameter updates do not interfere with non-target concepts, enforcing $e_{0}$ close to zero without a tradeoff between $e_{0}$ and $e_{1}$.
-To obtain $\boldsymbol{\Delta}$, we first perform singular value decomposition (SVD) on $\mathbf{C}_{0}\mathbf{C}_{0}^{\top}$***Considering $\mathbf{C}_{0}\in\mathbb{R}^{d_{0}\times N_{R}}$ may have high dimensionality depending on $N_{R}$, we compute the null space by performing SVD on $\mathbf{C}_{0}\mathbf{C}_{0}^{\top}\in\mathbb{R}^{d_{0}\times d_{0}}$, as both $\mathbf{C}_{0}$ and $\mathbf{C}_{0}\mathbf{C}_{0}^{\top}$ share the same null space.
-and construct a projection matrix $\mathbf{P}$:
-$\left\{\mathbf{U},\mathbf{\Lambda},\mathbf{U}^{\top}\right\}=\text{SVD}\left(\mathbf{C}_{0}\mathbf{C}_{0}^{\top}\right),\quad\mathbf{P}=\hat{\mathbf{U}}\hat{\mathbf{U}}^{\top},$ (2)
-where $\hat{\mathbf{U}}$ is the matrix composed of eigenvectors w.r.t.
-the zero singular values in $\mathbf{\Lambda}$.
-This projection matrix $\mathbf{P}$ can map $\boldsymbol{\Delta}$ onto the null space of $\mathbf{C}_{0}$, i.e., $\|\left(\boldsymbol{\Delta}\mathbf{P}\right)\mathbf{C}_{0}\|^{2}=0$.
-Thus, the objective becomes:
-$\left(\boldsymbol{\Delta}\mathbf{P}\right)_{\text{Null}}=$ $\mathop{\text{arg\ min}}\limits_{\boldsymbol{\Delta}}\underbrace{\|(\mathbf{W}+\boldsymbol{\Delta}\mathbf{P})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}}_{\text{$e_{1}$}}$
+However, as the number of target concepts increases, the accumulated preservation errors $e_{0}$, which prove to share a non-zero bound from Appx.
+B.2, across multiple target concepts would amplify the distortion on non-target knowledge and degrade prior preservation.
+3.2 Apply Null-Space Constraints
+To address the limitation of weighted optimization in prior preservation, SPEED incorporates null-space constraints Wang et al.
+(2021); Fang et al.
+(2024) to achieve prior-preserved model editing by forcing $e_{0}=0$.
+The null space of $\mathbf{C}_{0}$ consists of all vectors $\bm{v}$ such that $\bm{v}\mathbf{C}_{0}=\mathbf{0}$.
+Restricting the parameter update $\bm{\Delta}$ to this space ensures that such updates do not interfere with non-target concepts.
 […]
-where $\|\boldsymbol{\Delta}\mathbf{P}\|^{2}$ is an additional regularization term to guarantee convergence.
-This objective facilitates model parameter editing in the null space of the retain set $\mathbf{R}$, showing great promise for prior-preserved concept erasure.
-Figure 3: Prior shift comparison with erasing different target concepts.
-The prior shift is measured by $\|\boldsymbol{\Delta}_{\text{erase}}\boldsymbol{c}\|^{2}$ introduced in Sec.
-4.1, where $\boldsymbol{c}$ denotes the concept embedding.
-A larger value indicates a greater impact on the concept after erasure.
+$\left\{\mathbf{U},\mathbf{\Lambda},\mathbf{U}^{\top}\right\}=\text{SVD}\left(\mathbf{C}_{0}\mathbf{C}_{0}^{\top}\right),\ \ \mathbf{U}\in\mathbb{R}^{d_{0}\times d_{0}}\xrightarrow[\text{values}]{\text{zero singular}}\hat{\mathbf{U}}\Longrightarrow\mathbf{P}=\hat{\mathbf{U}}\hat{\mathbf{U}}^{\top}.$ (2)
+The final update applied to model parameters is $\bm{\Delta}\mathbf{P}$, which projects $\bm{\Delta}$ onto the null space of $\mathbf{C}_{0}$.
+This ensures that updates do not interfere with non-target concepts, satisfying $\|(\bm{\Delta}\mathbf{P})\mathbf{C}_{0}\|^{2}=0$.
+To solve for the updates, we minimize the following objective:
+$\bm{\Delta}_{\text{Null}}=\mathop{\text{arg\ min}}\limits_{\bm{\Delta}}\underbrace{\|(\mathbf{W}+\bm{\Delta}\mathbf{P})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}}_{\text{$e_{1}$}}\ +\ \cancel{{\color[rgb]{.5,.5,.5}\underbrace{\|\left(\bm{\Delta}\mathbf{P}\right)\mathbf{C_{0}}\|^{2}}_{e_{0}=0}}}\ +\underbrace{\|\bm{\Delta}\mathbf{P}\|^{2}}_{\text{regularization}},$ (3)
+where $\|\bm{\Delta}\mathbf{P}\|^{2}$ is a regularization term to ensure convergence.
+The preservation term $\|(\bm{\Delta}\mathbf{P})\mathbf{C}_{0}\|^{2}$ is omitted, as it is guaranteed to be zero by the null-space constraint.
+This objective enables us to update the model parameters such that target concepts are effectively erased while non-target representations remain unaffected, thereby achieving prior-preserved concept erasure.
 4 Prior Knowledge Refinement
-The core of null-space constraints in SPEED is to determine a projection matrix $\mathbf{P}$ that constrains the solution space of $\boldsymbol{\Delta}$ within the null space of the retain set $\mathbf{R}$, ensuring that non-target concepts remain unaffected after erasure.
-The rank of the matrix $\mathbf{C}_{0}\mathbf{C}_{0}^{\top}$ reflects the diversity of the retain set, directly shaping the structure of its null space, which is determined by $\mathbf{R}$.
-As $\mathbf{R}$ increases, $\mathbf{C}_{0}\mathbf{C}_{0}^{\top}$ gradually reaches full rank, its null space narrows and reduces to the trivial null space $\{\mathbf{0}\}$.
-However, shrinking $\mathbf{R}$ would eliminate prior coverage by allowing more unintended shifts in non-target concepts not included.
-To balance this dilemma, we introduce three complementary techniques: Influence-Based Prior Filtering (IPF) to retain the most affected non-target concepts while avoiding excessive rank saturation, Directed Prior Augmentation (DPA) to expand prior coverage with semantically consistent variations, and Invariant Equality Constraints (IEC) to enforce additional constraints to preserve key invariants in generation.
+However, as more diverse non-target concepts are included in the retain set, the rank of the correlation matrix $\mathbf{C}_{0}\mathbf{C}_{0}^{\top}$ increases [^2] We assume that the concepts are not exactly linearly dependent in the representation space, which is generally satisfied in practice due to the semantic diversity and high dimensionality of the embedding space..
+The null space, defined as the orthogonal complement of this span, correspondingly shrinks in dimension:
+$\dim(\mathrm{Null}(\mathbf{C}_{0}))=d_{0}-\mathrm{rank}(\mathbf{C}_{0}\mathbf{C}_{0}^{\top}).$ (4)
+Herein, the null space dimension characterizes the degrees of freedom available for editing without affecting the retained concepts.
+However, as this dimension shrinks, to ensure sufficient degrees of freedom for concept erasure, we are compelled to include singular vectors w.r.t.
+non-zero singular values in $\hat{\mathbf{U}}$ following Fang et al.
+(2024), which leads to an approximate null space and induces semantic degradation within the retain set (see Fig.
+2).
+To improve, we propose Prior Knowledge Refinement, a structured strategy for refining the retain set to enable accurate null-space construction, with three complementary techniques: Influence-Based Prior Filtering (Sec.
+4.1) to discard weakly affected non-target concepts to form a viable null space; Directed Prior Augmentation (Sec.
+4.2) to expand the retain set with targeted and semantically consistent variations; and Invariant Equality Constraints (Sec.
+4.3) to enforce equality constraints to preserve critical invariants during generation.
 4.1 Influence-Based Prior Filtering (IPF)
-Given a pre-defined retain set, existing editing-based methods [18, 19] generally neglect to account for the varying influence of different target concepts on non-target elements, where closer semantic relationships (e.g., Mickey and Dog for Snoopy) lead to greater prior shifts, as shown in Fig.
-3.
-This differing influence suggests that not all non-target concepts contribute equally to preserving prior knowledge, and weakly influenced concepts provide marginal benefit while introducing additional ranks to constrain the null space.
-To improve, we propose to filter the original retain set by selectively retaining those non-target concepts most influenced by the erased target concepts.
-To evaluate this influence, we first obtain the closed-form solution $\boldsymbol{\Delta}_{\text{erase}}$, which only considers the erasure efficacy during editing.
-Specifically, we remove the preservation error $e_{0}$ from Eq.
-1 and introduce an additional regularization term $\|\boldsymbol{\Delta}\|^{2}$ to guarantee convergence:
-$\boldsymbol{\Delta}_{\text{erase}}$ $=\mathop{\text{arg\ min}}\limits_{\boldsymbol{\Delta}}\underbrace{\|(\mathbf{W}+\boldsymbol{\Delta})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}}_{e_{1}}+\underbrace{\|\boldsymbol{\Delta}\|^{2}}_{\text{regularization}}$
-$=\mathbf{W}\left(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\right)\left(\mathbf{I}+\mathbf{C}_{1}\mathbf{C}_{1}\right)^{-1}.$ (4)
-Then we can quantify the influence on each concept embedding $\boldsymbol{c}$ by measuring the prior shift $=\|\boldsymbol{\Delta}_{\text{erase}}\boldsymbol{c}\|^{2}$.
-This metric stems from the intuition that after erasing the target concept, a larger prior shift in the non-target concept indicates a greater influence on the corresponding prior knowledge.
-Building upon this, we can filter the original retain set $\mathbf{R}$ with the prior shift by the following process:
-$\mathbf{R}_{f}:\mathbf{R}\mapsto\{\boldsymbol{c}_{0}\in\mathbf{R}\mid\|\boldsymbol{\Delta}_{\text{erase}}\boldsymbol{c}_{0}\|^{2}>\mu\},$ (5)
-where the mean value $\mu=\mathbb{E}_{\boldsymbol{c}_{0}\sim\mathbf{R}}\left[\|\boldsymbol{\Delta}_{\text{erase}}\boldsymbol{c}_{0}\|^{2}\right]$ serves as the filtering threshold.
+Given a predefined retain set, existing editing-based methods Gandikota et al.
+(2024); Gong et al.
+(2025) treat all non-target concepts equally when enforcing prior preservation.
+However, an overlooked fact is that parameter updates inherently induce output changes over non-target concepts, and these changes vary across different non-target concepts.
+This suggests that not all non-target concepts contribute equally to preserving prior knowledge, and weakly influenced concepts offer little benefit but introduce additional ranks that narrow the null space.
+To this end, we propose an explicit and model-consistent metric, i.e., prior shift, to quantify how much a non-target concept is affected by concept erasure.
+Specifically, we isolate the effect of erasure by solving for a closed-form update $\bm{\Delta}_{\text{erase}}$ that minimizes only the erasure error $e_{1}$ while discarding the preservation term $e_{0}$ from Eq.
+1:
+$\bm{\Delta}_{\text{erase}}=\mathop{\text{arg\ min}}\limits_{\bm{\Delta}}\underbrace{\|(\mathbf{W}+\bm{\Delta})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}}_{e_{1}}+\underbrace{\|\bm{\Delta}\|^{2}}_{\text{regularization}}=\mathbf{W}\left(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\right)\left(\mathbf{I}+\mathbf{C}_{1}\mathbf{C}_{1}\right)^{-1}.$ (5)
+where $\|\bm{\Delta}\|^{2}$ is introduced for convergence.
+Then, for each non-target concept embedding $\bm{c}$, we define its prior shift as: $\|\bm{\Delta}_{\text{erase}}\bm{c}\|^{2}$.
+This value offers a faithful reflection of how parameter updates perturb a non-target concept in the feature space with closed-form computation, and can naturally generalize to assessing multi-concept erasure effects.
+Based on this, we filter the original retain set $\mathbf{R}$ to focus only on highly influenced concepts:
+$\mathbf{R}_{f}:\mathbf{R}\mapsto\{\bm{c}_{0}\in\mathbf{R}\mid\|\bm{\Delta}_{\text{erase}}\bm{c}_{0}\|^{2}>\mu\},$ (6)
+where the mean value $\mu=\mathbb{E}_{\bm{c}_{0}\sim\mathbf{R}}\left[\|\bm{\Delta}_{\text{erase}}\bm{c}_{0}\|^{2}\right]$ serves as a filtering threshold.
 4.2 Directed Prior Augmentation (DPA)
-To prevent IPF from generating an overly small retain set, which would compromise prior coverage, an intuitive strategy is to augment the prior knowledge by perturbing the non-target embedding $\boldsymbol{c}_{0}$ with random noise [37].
-However, this strategy would introduce meaningless embeddings that fail to generate semantically coherent images (e.g., noise image), resulting in excessive preservation with increasing ranks.
-To search for more semantically consistent concepts, we introduce directed noise by projecting the random noise $\boldsymbol{\epsilon}$ onto the direction in which the model parameters $\mathbf{W}$ exhibit minimal variation.
-This operation ensures that the perturbed embeddings express closer semantics to the original concept after being mapped by $\mathbf{W}$, as shown in Fig.
-4.
-Specifically, we first derive a projection matrix $\mathbf{P}_{\text{min}}$:
-$\left\{\mathbf{U}_{\mathbf{W}},\mathbf{\Lambda}_{\mathbf{W}},\mathbf{U}_{\mathbf{W}}^{\top}\right\}=\text{SVD}\left(\mathbf{W}\right),\quad\mathbf{P}_{\text{min}}=\mathbf{U}_{\text{min}}\mathbf{U}_{\text{min}}^{\top},$ (6)
-where $\mathbf{U}_{\text{min}}={\mathbf{U}}_{\mathbf{W}}[:,-r:]$ denotes the eigenvectors w.r.t.
-the smallest $r$ singular values, which represent the $r$ least-changing directions of $\mathbf{W}$.
-Then the directed noise $\boldsymbol{\epsilon}\cdot\mathbf{P}_{\text{min}}$ is used to perturb the original embedding via:
-$\boldsymbol{c}^{\prime}_{0}=\boldsymbol{c}_{0}+\boldsymbol{\epsilon}\cdot\mathbf{P}_{\text{min}},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$ (7)
-Given a retain set $\mathbf{R}$, the augmentation process can be formulated as follows:
-$\mathbf{R}^{\text{aug}}:\mathbf{R}\mapsto\bigcup_{\boldsymbol{c}_{0}\in\mathbf{R}}\left\{\boldsymbol{c}_{0,k}^{\prime}\mid k=1,\dots,N_{A}\right\},$ (8)
-where $N_{A}$ denotes the augmentation times and $\boldsymbol{c}_{0,k}^{\prime}$ represents the $k$-th augmented embedding given $\boldsymbol{c}_{0}\in\mathbf{R}$ using Eq.
-7.
-In implementation, we first filter the original retain set $\mathbf{R}$ to obtain $\mathbf{R}_{f}$ using IPF.
-Subsequently, further augmentation and filtering are applied to $\mathbf{R}_{f}$ using DPA and IPF to obtain $\left(\mathbf{R}_{f}\right)^{\text{aug}}_{f}$, and the two filtered retain sets are then combined together to serve as the final refined retain set $\mathbf{R}_{\text{refine}}=\mathbf{R}_{f}\cup\left(\mathbf{R}_{f}\right)^{\text{aug}}_{f}$.
 Figure 3: t-SNE distribution of perturbing the original concept with random noise and our directed noise.
@@
 (b) Our directed noise preserves semantic similarity to the original concept with closer distances in the space mapped by $\mathbf{W}$.
-Figure 5: Qualitative comparison of the few-concept erasure in erasing single and multiple instances.
-The erased and preserved generations are highlighted with red and green boxes, respectively.
-Our method demonstrates superior prior preservation compared to the SOTA method, UCE [18], in single-concept erasure, and this superiority is amplified as the target concepts increase.
-For example, in the Hello Kitty case, our generation exhibits less alteration after erasing three concepts, whereas UCE’s generations are gradually deteriorating.
+To enhance prior preservation with broader retain coverage, an intuitive strategy is to augment the retain set by perturbing non-target embedding $\bm{c}_{0}$ with random noise Lyu et al.
+(2024).
+However, this strategy would introduce meaningless embeddings that fail to generate semantically coherent images (e.g., noise image), resulting in excessive preservation with increasing ranks.
+To search for more semantically consistent concepts, we introduce directed noise by projecting the random noise $\bm{\epsilon}$ onto the direction in which the model parameters $\mathbf{W}$ exhibit minimal variation.
+This operation ensures the perturbed embeddings express closer semantics to the original concept after being mapped by $\mathbf{W}$ in Fig.
+3.
+Specifically, we first derive a projection matrix $\mathbf{P}_{\text{min}}$:
+$\left\{\mathbf{U}_{\mathbf{W}},\mathbf{\Lambda}_{\mathbf{W}},\mathbf{U}_{\mathbf{W}}^{\top}\right\}=\text{SVD}\left(\mathbf{W}\right),\quad\mathbf{P}_{\text{min}}=\mathbf{U}_{\text{min}}\mathbf{U}_{\text{min}}^{\top},$ (7)
+where $\mathbf{U}_{\text{min}}={\mathbf{U}}_{\mathbf{W}}[:,-r:]$ denotes the singular vectors w.r.t.
+the smallest $r$ singular vectors [^3] Empirically, the model parameter matrix $\mathbf{W}$ is usually full rank, thus its all singular values are non-zero., which represent the $r$ least-changing directions of $\mathbf{W}$ and constrain the rank of the augmented embeddings to a maximum of $r$.
+Then the directed noise $\bm{\epsilon}\cdot\mathbf{P}_{\text{min}}$ is used to perturb the original embedding via:
+$\bm{c}^{\prime}_{0}=\bm{c}_{0}+\bm{\epsilon}\cdot\mathbf{P}_{\text{min}},\quad\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$ (8)
+Given a retain set $\mathbf{R}$, the augmentation process can be formulated as follows:
+$\mathbf{R}^{\text{aug}}:\mathbf{R}\mapsto\bigcup_{\bm{c}_{0}\in\mathbf{R}}\left\{\bm{c}_{0,k}^{\prime}\mid k=1,\dots,N_{A}\right\},$ (9)
+where $N_{A}$ denotes the augmentation times and $\bm{c}_{0,k}^{\prime}$ represents the $k$-th augmented embedding given $\bm{c}_{0}\in\mathbf{R}$ using Eq.
+8.
+In implementation, we first filter the original retain set $\mathbf{R}$ to obtain $\mathbf{R}_{f}$ using IPF.
+Subsequently, further augmentation and filtering are applied to $\mathbf{R}_{f}$ using DPA and IPF to obtain $\left(\mathbf{R}_{f}\right)^{\text{aug}}_{f}$.
+Finally, we combine them to serve as the final refined retain set $\mathbf{R}_{\text{refine}}=\mathbf{R}_{f}\cup\left(\mathbf{R}_{f}\right)^{\text{aug}}_{f}$.
 4.3 Invariant Equality Constraints (IEC)
-In parallel, we also identify certain invariants during the T2I generation process, i.e., intermediate variables that remain unchanged with sampling prompts.
+In parallel, we identify certain invariants during the T2I generation process, i.e., intermediate variables that remain unchanged with varying sampling prompts.
 One such invariant is the CLIP-encoded [SOT] token.
-Since the encoding process is masked by causal attention and all prompts are prefixed with the fixed [SOT] token during tokenization, the corresponding embedding remains unchanged.
-Another invariant is the null-text embedding, as it corresponds to the unconditional generation under the classifier-free guidance [23], which also remains unchanged despite prompt variations.
-Given the invariance of these embeddings, we consider additional protection measures for these invariant embeddings to ensure their outputs remain unchanged during concept erasure.
-Specifically, we introduce explicit equality constraints for invariants based on Eq.
-3.2:
-$\begin{gathered}\mathop{\text{min}}\limits_{\boldsymbol{\Delta}}\underbrace{\|(\mathbf{W}+\boldsymbol{\Delta}\mathbf{P})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}}_{e_{1}}+\underbrace{\|\boldsymbol{\Delta}\mathbf{P}\|^{2}}_{\text{regularization}},\\ \quad\text{s.t.}\,\underbrace{(\boldsymbol{\Delta}\mathbf{P})\mathbf{C}_{2}=\mathbf{0}}_{\text{equality constraints}},\end{gathered}$ (9)
-where $\mathbf{C}_{2}$ denotes the stacked invariant embedding matrix of [SOT] and null-textSince the null-text embeddings are only composed of [EOT] tokens (excluding [SOT]), we use the k-means algorithm [38] to select k centroids to reduce redundancy..
-With the projection matrix $\mathbf{P}$ derived from $\mathbf{R}_{\text{refine}}$, we can compute the closed-form solution of Eq.
+Since the encoding process is masked by causal attention and all prompts are prefixed with the fixed [SOT] token during tokenization, its embedding consistently remains unchanged during T2I process.
+Another invariant is the null-text embedding, as it corresponds to the unconditional generation under the classifier-free guidance Ho & Salimans (2022), which also remains unchanged despite prompt variations.
+Given the invariance of these embeddings, we consider additional protection measures to ensure their outputs remain unchanged during concept erasure.
+Specifically, we introduce explicit equality constraints over invariants based on Eq.
+3:
+$\mathop{\text{min}}\limits_{\bm{\Delta}}\underbrace{\|(\mathbf{W}+\bm{\Delta}\mathbf{P})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}}_{e_{1}}+\underbrace{\|\bm{\Delta}\mathbf{P}\|^{2}}_{\text{regularization}},\quad\text{s.t.}\,\underbrace{(\bm{\Delta}\mathbf{P})\mathbf{C}_{2}=\mathbf{0}}_{\text{equality constraints}},$ (10)
+where $\mathbf{C}_{2}$ denotes the stacked invariant embedding matrix of [SOT] and null-text.
+Derive the projection matrix $\mathbf{P}$ from $\mathbf{R}_{\text{refine}}$, we can compute the closed-form solution of Eq.
 10 using Lagrange Multipliers from Appx.
 B.3:
-$\left(\boldsymbol{\Delta}\mathbf{P}\right)_{\text{SPEED}}=\mathbf{W}\left(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\right)\mathbf{P}\mathbf{Q}\mathbf{M},$ (10)
+$\left(\bm{\Delta}\mathbf{P}\right)_{\text{Ours}}=\mathbf{W}\left(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\right)\mathbf{P}\mathbf{Q}\mathbf{M},$ (11)
 where
-$\begin{gathered}\mathbf{M}=\left(\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\mathbf{P}+\mathbf{I}\right)^{-1},\\ \mathbf{Q}=\mathbf{I}-\mathbf{M}\mathbf{C}_{2}\left(\mathbf{C}_{2}^{\top}\mathbf{P}\mathbf{M}\mathbf{C}_{2}\right)^{-1}\mathbf{C}_{2}^{\top}\mathbf{P}.\end{gathered}$ (11)
+$\mathbf{M}=\left(\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\mathbf{P}+\mathbf{I}\right)^{-1},\mathbf{Q}=\mathbf{I}-\mathbf{M}\mathbf{C}_{2}\left(\mathbf{C}_{2}^{\top}\mathbf{P}\mathbf{M}\mathbf{C}_{2}\right)^{-1}\mathbf{C}_{2}^{\top}\mathbf{P}.$ (12)
 This closed-form solution enforces the equality constraints by projecting the parameter update onto the subspace orthogonal to the invariant embeddings.
 Since image generation inevitably depends on these invariant embeddings, such constraints inherently preserve prior knowledge.
+Figure 4: Qualitative comparison of the few-concept erasure in erasing instances.
+The erased and preserved generations are highlighted with red and green boxes, respectively.
+Our method exhibits consistent prior preservation with less semantic degradation for non-target concepts.
+For example, the middle column better retains details such as Mickey’s hat and button count, and the right column demonstrates more consistent Hello Kitty generations along with three concepts erased.
 5 Experiments
-In this section, we conduct extensive experiments on three representative erasure tasks, including few-concept erasure, multi-concept erasure, and implicit concept erasure, validating our superior prior preservation.
-The compared baselines include ConAbl [29], MACE [35], RECE [19], and UCE [18], which have achieved SOTA performance across various concept erasure tasks.
-In implementation, we conduct all experiments on SDv1.4 [1] and generate each image using DPM-solver sampler [34] over 20 sampling steps with classifier-free guidance [23] of 7.5.
-More implementation details and compared baselines (e.g., SPM [37]) can be found in Appx.
+In this section, we conduct extensive experiments on three representative erasure tasks, including few-concept erasure, multi-concept erasure, and implicit concept erasure (Appx.
+D.4), validating our superior prior preservation.
+The compared baselines include ConAbl Kumari et al.
+(2023), MACE Lu et al.
+(2024a), RECE Gong et al.
+(2025), and UCE Gandikota et al.
+(2024), which have achieved SOTA performance across various concept erasure tasks.
+In implementation, we conduct all experiments on SDv1.4 AI (2022) and generate each image using DPM-solver sampler Lu et al.
+(2022) over 20 sampling steps with classifier-free guidance Ho & Salimans (2022) of 7.5.
+More implementation details and compared baselines can be found in Appx.
 C and Appx.
-D.2.
+D.3.
 5.1 On Few-Concept Erasure
 Evaluation setup.
-To compare the few-concept erasure performance with baseline methods, we conduct experiments on instance erasure and artistic style erasure following [37], where all methods are evaluated based on 80 instance templates and 30 artistic style templates, generating 10 images per template per concept.
-We use two metrics for evaluation: CLIP Score (CS) [45] measuring the text-image similarity and Fréchet Inception Distance (FID) [22] assessing the distributional distance before and after erasure.
-Following [37], we select non-target concepts with similar semantics to the target concept for comparison and report CS for targets and FID for non-targets in the main paper.
-Full comparisons are presented in Appx.
-D.1.
-We further compare the generations on MS-COCO captions [33], where we generate images with the first 1,000 captions, and report the CS and FID as the measure of general knowledge preservation.
+We evaluate few-concept erasure on instance erasure and artistic style erasure following Lyu et al.
+(2024), using 80 instance templates and 30 artistic style templates with 10 images per template per concept.
+We use two metrics for evaluation: CLIP Score (CS) Radford et al.
+(2021) for the text-image similarity and Fréchet Inception Distance (FID) Heusel et al.
+(2017) for the distributional distance before and after erasure.
+Following Lyu et al.
+(2024), we select non-target concepts with similar semantics to the target concept for comparison and report CS for targets and FID for non-targets in the main paper.
+Complete comparisons are presented in Appx.
+D.2.
+We further compare the generations on MS-COCO captions Lin et al.
+(2014), where we generate images with the first 1,000 captions, and report CS and FID to measure general knowledge preservation.
+Table 1: Quantitative comparison of the few-concept erasure on instances (left) and artistic styles (right) following Lyu et al.
+(2024).
+Arrows indicate the preferred direction for each metric, and the best results are highlighted in bold.
+Our method consistently improves prior preservation for non-target and general concepts from MS-COCO (shaded in pink) while achieving effective concept erasure.
+While our CS is not the lowest for target concpet, Appx.
+D.1 and Fig.
+7 show our method is sufficient for erasure, and lower CS may further compromise prior preservation.
 Concept Snoopy Mickey Spongebob Pikachu Hello Kitty MS-COCO
@@
 Ours 23.69 23.93 21.39 21.40 26.22 26.51 24.99
-Table 1: Quantitative comparison of the few-concept erasure in erasing instances, where the erased target concept(s) are shaded in pink.
-Arrows on the headers indicate the preferred direction for each metric, and the best results are highlighted in bold.
-Our method, which does not achieve the lowest CS but has been proven sufficient in Fig.
-5, consistently demonstrates significant improvements in prior preservation for non-target concepts.
-Analysis and discussion.
-Tables 1 and 2 compare the results of erasing various instance concepts and artistic styles.
-Our method consistently achieves the lowest FIDs across all non-target concepts, demonstrating superior prior preservation with minimal alteration to the original content.
-Moreover, we emphasize that our erasure is sufficiently effective, even without achieving the lowest CS, as shown in Fig.
-5.
-In contrast, lower CS values typically indicate over-erasure, which results in excessive degradation of prior knowledge.
-Notably, with the number of target concepts increasing from 1 to 3, our FID in Pikachu rises from 16.81 to 21.40 (4.59 ↑), while UCE increases from 19.06 to 29.20 (10.14 ↑).
-A similar pattern is observed in Hello Kitty (Our 4.48 ↑ v.s.
-UCE’s 10.29 ↑), showing the robustness of our method in erasing increasing multiple concepts.
-5.2 On Multi-Concept Erasure
-Evaluation setup.
-Another more realistic erasure scenario is multi-concept erasure, where massive concepts are required to be erased at once.
-Herein, we follow the experiment setup in [35] for erasing multiple celebrities, where we experiment with erasing 10, 50, and 100 celebrities and collect another 100 celebrities as non-target concepts.
-We prepare 5 prompt templates for each celebrity concept.
-For non-target concepts, we generate 1 image per template for each of the 100 concepts, totaling 500 images.
-For target concepts, we adjust the per-concept quantity to maintain a total of 500 images (e.g., erasing 10 celebrities involves generating 10 images with 5 templates per concept).
-In evaluation, we adopt GIPHY Celebrity Detector (GCD) [20] and measure the top-1 GCD accuracy, indicated by $\text{Acc}_{e}$ for erased target concepts and $\text{Acc}_{r}$ for retained non-target concepts.
-Meanwhile, the harmonic mean $H_{o}$ is adopted to assess the overall erasure performance:
-$H_{o}=\frac{2}{\left(1-\text{Acc}_{e}\right)^{-1}+\left(\text{Acc}_{r}\right)^{-1}}.$ (12)
-Additionally, we report the results on MS-COCO to demonstrate the prior preservation of general concepts.
 Concept Van Gogh Picasso Monet Paul Gauguin Caravaggio MS-COCO
@@
 Ours 28.78 41.21 25.06 27.85 55.20 26.48 20.87
-Table 2: Quantitative comparison of the few-concept erasure in erasing single artistic style.
-Erase 10 Celebrities MS-COCO Erase 50 Celebrities MS-COCO Erase 100 Celebrities MS-COCO
-$\text{Acc}_{e}$ $\text{Acc}_{r}$ $H_{o}$ CS ↑ FID ↓ $\text{Acc}_{e}$ $\text{Acc}_{r}$ $H_{o}$ CS ↑ FID ↓ $\text{Acc}_{e}$ $\text{Acc}_{r}$ $H_{o}$ CS ↑ FID ↓
-SD v1.4 91.99 89.66 14.70 26.53 - 93.08 89.66 12.85 26.53 - 90.18 89.66 17.70 26.53 -
-UCE 0.20 71.19 83.10 24.07 83.81 0.00 31.94 48.41 13.45 209.93 0.00 20.92 34.60 13.49 185.46
-RECE 0.34 67.43 80.44 16.75 170.65 1.03 19.77 32.95 13.49 213.39 2.43 23.71 38.16 12.09 177.57
-MACE 1.62 87.73 92.75 26.36 37.25 3.41 84.31 90.03 25.45 45.31 4.80 80.20 87.06 24.80 50.41
-Ours 1.81 89.09 93.42 26.47 30.02 3.46 88.48 92.34 26.46 39.23 5.87 85.54 89.63 26.22 44.97
+Analysis and discussion.
+Table 1 compares the results of erasing various instance concepts and artistic styles.
+Our method consistently achieves the lowest FIDs across all non-target concepts, demonstrating superior prior preservation with minimal alteration to the original content.
+Moreover, we emphasize that our erasure is sufficiently effective, even without achieving the lowest CS, as shown in Fig.
+4 and Appx.
+D.1.
+In contrast, lower CS values typically indicate over-erasure, which results in excessive degradation of prior knowledge.
+Notably, with the number of target concepts increasing from 1 to 3, our FID in Pikachu rises from 16.81 to 21.40 (4.59 ↑), while UCE increases from 19.06 to 29.20 (10.14 ↑).
+A similar pattern is observed in Hello Kitty (Our 4.48 ↑ v.s.
+UCE’s 10.29 ↑), showing our robustness in erasing increasing target concepts.
+5.2 On Multi-Concept Erasure
+Evaluation setup.
+Another more realistic erasure scenario is multi-concept erasure, where massive concepts are required to be erased at once.
+We follow the experiment setup in Lu et al.
+(2024a) for erasing multiple celebrities, where we experiment with erasing 10, 50, and 100 celebrities and collect another 100 celebrities as non-target concepts.
+Specifically, we prepare 5 prompt templates for each celebrity concept.
+For non-target concepts, we generate 1 image per template for each of the 100 concepts, totaling 500 images.
+For target concepts, we adjust the per-concept quantity to maintain a total of 500 images (e.g., erasing 10 celebrities involves generating 10 images with 5 templates per concept).
+In evaluation, we adopt GIPHY Celebrity Detector (GCD) Hasty et al.
+(2019) and measure the top-1 GCD accuracy, indicated by $\text{Acc}_{e}$ for erased target concepts and $\text{Acc}_{r}$ for retained non-target concepts.
+Meanwhile, the harmonic mean $H_{o}=\frac{2}{\left(1-\text{Acc}_{e}\right)^{-1}+\left(\text{Acc}_{r}\right)^{-1}}$ is adopted to assess the overall erasure performance.
+Additionally, we report the results on MS-COCO to demonstrate the prior preservation of general concepts.
+Figure 5: Quantitative comparison of the multi-concept erasure in erasing celebrities (celeb).
+The erased and preserved generations are marked with red and green boxes.
+Our method precisely erases 100 celebrities while preserving generations of other non-target concepts.
 Table 2: Quantitative comparison of the multi-concept erasure in erasing 10, 50, and 100 celebrities.
 The best results are highlighted in bold.
-Our method is capable of erasing up to 100 celebrities at once with low $\text{Acc}_{e}$ (%) and preserving other non-target celebrities with less appearance alteration with high $\text{Acc}_{r}$ (%), resulting in the best overall erasure performance $H_{o}$ (shaded in pink).
-NudeNet Detection Results on I2P MS-COCO
-Armpits Belly Buttocks Feet Breasts (F) Genitalia (F) Breasts (M) Genitalia (M) Total ↓ CS ↑ FID ↓
-SD v1.4 123 134 19 14 258 9 16 3 576 26.53 -
-ConAbl 24 43 5 6 68 1 6 4 157 26.14 39.26
-MACE 28 19 1 20 37 3 6 5 119 24.06 52.78
-RECE 17 29 3 7 14 1 8 1 80 25.98 40.37
-UCE 29 42 2 11 36 3 9 7 139 26.24 38.60
-Ours 20 42 7 3 29 2 5 5 113 26.29 37.82
-Table 4: Evaluation of implicit concept erasure on I2P benchmark.
-We report the number of nude body parts (F: Female, M: Male) detected by the NudeNet with threshold $=$ 0.6.
-The best and second-best results are marked in bold and underlined.
-(Left) Our method effectively removes nude content, even though nudity is not explicitly mentioned in prompts from I2P, achieving the second-best total count.
-(Right) Our method also consistently achieves superior prior preservation for non-target concepts to other methods on MS-COCO.
+Our method can effectively erase up to 100 celebrities simultaneously, achieving low $\text{Acc}_{e}$ (%) and high $\text{Acc}_{r}$ (%) that preserve non-target celebrities with minimal appearance changes.
+This yields the best overall erasure performance $H_{o}$ and competitive runtime (s) on one A100 GPU, successfully erasing 100 concepts in just 5 seconds.
+Erase 10 Celebrities MS-COCO Erase 50 Celebrities MS-COCO Erase 100 Celebrities MS-COCO
+$\text{Acc}_{e}$ $\text{Acc}_{r}$ $H_{o}$ Time ↓ CS ↑ FID ↓ $\text{Acc}_{e}$ $\text{Acc}_{r}$ $H_{o}$ Time ↓ CS ↑ FID ↓ $\text{Acc}_{e}$ $\text{Acc}_{r}$ $H_{o}$ Time ↓ CS ↑ FID ↓
+SD v1.4 91.99 89.66 14.70 - 26.53 - 93.08 89.66 12.85 - 26.53 - 90.18 89.66 17.70 - 26.53 -
+ConAbl 60.76 77.89 52.19 900 25.60 42.12 64.00 75.44 48.74 4,500 14.30 255.36 42.86 58.82 57.97 9,000 14.93 235.27
+UCE 0.20 71.19 83.10 1.5 24.07 83.81 0.00 31.94 48.41 1.8 13.45 209.93 0.00 20.92 34.60 2.1 13.49 185.46
+RECE 0.34 67.43 80.44 2.5 16.75 170.65 1.03 19.77 32.95 6.3 13.49 213.39 2.43 23.71 38.16 11.0 12.09 177.57
+MACE 1.62 87.73 92.75 207 26.36 37.25 3.41 84.31 90.03 936 25.45 45.31 4.80 80.20 87.06 1736 24.80 50.41
+Ours 1.81 89.09 93.42 3.8 26.47 30.02 3.46 88.48 92.34 4.2 26.46 39.23 5.87 85.54 89.63 5.0 26.22 44.97
 Analysis and discussion.
 Table 2 showcases a notable improvement of our method on multi-concept erasure, particularly in prior preservation with the highest $\text{Acc}_{r}$.
-In comparison with the SOTA method, MACE [35], our method achieves superior prior preservation with better $\text{Acc}_{r}$, while maintaining comparable erasure efficacy, as reflected in similar $\text{Acc}_{e}$, resulting in the best overall erasure performance indicated by the highest $H_{o}$.
+In comparison with the SOTA method, MACE Lu et al.
+(2024a), our method achieves superior prior preservation with better $\text{Acc}_{r}$, while maintaining comparable erasure efficacy, as reflected in similar $\text{Acc}_{e}$, resulting in the best overall erasure performance indicated by the highest $H_{o}$.
 Meanwhile, our method attains the lowest FID across all methods on MS-COCO.
-The other methods, UCE [18] and RECE [19], although achieving considerable balance in few-concept erasure, fail to maintain this balance as the number of target concepts increases as shown in Fig.
-6, with catastrophic prior damage evidenced by MS-COCO as well.
-Notably, our method can erase up to 100 celebrities in 5 seconds, whereas MACE requires around 30 minutes (×$\times 350$ time).
+The other methods, UCE Gandikota et al.
+(2024) and RECE Gong et al.
+(2025), although achieving considerable balance in few-concept erasure, fail to maintain this balance as the number of target concepts increases as shown in Fig.
+5, with catastrophic prior damage evidenced by MS-COCO as well.
+Notably, our method can erase up to 100 celebrities in 5 seconds, whereas MACE requires around 30 minutes ($\times 350$ time).
 In real-world scenarios, this efficiency underscores our potential for the instant erasure of massive concepts.
-5.3 On Implicit Concept Erasure
-Evaluation setup.
-We further evaluate the erasure efficacy on implicit concepts, where the target concept does not explicitly appear in the text prompt.
-We conduct experiments on the Inappropriate Image Prompt (I2P) benchmark [51], which consists of various implicit inappropriate prompts involving violence, sexual content, and nudity.
-We follow the same setting in [19] to erase nudity $\to$ ‘ ’.
-Specifically, we generate images using all 4,703 text prompts in I2P and use NudeNet [6] to identify if the nude content is successfully erased with the threshold of 0.6.
-Additionally, we report the results on MS-COCO to demonstrate the prior preservation of general concepts.
-Analysis and discussion.
-As shown in Table 4, our method can effectively erase the implicit concept, i.e., nudity, with the second-best number of detected nude body parts.
-The SOTA method, RECE [19], achieves the best total number by extending the erasure set with more target concepts, but this comes at the cost of sacrificing prior preservation on MS-COCO.
-In contrast, our method achieves the best prior preservation, demonstrating effective erasure while maintaining strong prior preservation, striking a favorable balance between erasure and preservation.
-Figure 6: Quantitative comparison of multi-concept erasure in erasing different numbers of celebrities (celeb).
-The erased and preserved generations are highlighted with red and green boxes, respectively.
-Our method can precisely erase 100 celebrities while maintaining the proper generation of other non-target concepts.
+Figure 6: More applications across various T2I diffusion models.
+(a) We conduct composite concept erasure for “Snoopy + Van Gogh” on DreamShaper Lykon (2023) (1st row) and RealisticVision SG161222 (2023) (2nd row).
+(b) Our method also enables model knowledge editing by specifying the anchor concept on SDXL Podell et al.
+(2023).
+(c) Our method can seamlessly transfer to novel DiT-based T2I models, e.g., SDv3 Esser et al.
+(2024).
 5.3 Further Analysis
-Duration comparison.
-Table 5 presents the duration comparison in erasing 1, 10, and 100 concepts across different methods.
-It is obvious that training-based methods necessitate significantly higher computational costs than editing-based ones.
-In contrast, our method achieves precise multi-concept erasure in a remarkably short time, demonstrating superior efficiency while maintaining erasure performance, as evidenced in Table 3.
-Component ablation.
-In Table 6, we compare the individual impact of our components on prior preservation and draw the following conclusions: (1) Impact of IEC (Ablation 1 v.s.
-2): IEC reduces the non-target FID and the MS-COCO FID, demonstrating its effectiveness by preserving invariant embeddings with equality constraints.
-(2) Impact of IPF (Ablation 2 v.s.
-3): Incorporating IPF results in a significant improvement in both FIDs, underscoring its critical role in filtering out less-influenced concepts in the retain set under rank saturation.
-(3) Impact of DPA (Ablation 4 v.s.
-Ours): DPA improves RPA with directed noise and leads to a substantial improvement in non-target and MS-COCO FIDs, highlighting its advantage by introducing semantically similar concepts into the refined retain set.
-To conclude, the proposed three components (i.e., IEC, IPF, and DPA) improve the prior preservation from different perspectives and contribute to our method with the best prior preservation under null space constraints.
-Training-based Editing-based
-ConAbl MACE UCE RECE Ours
-Data Prepare $n\times 1000t_{1}$ $n\times(8t_{1}+8t_{2})$ 0 0 0
-1 concept $1\times 90$ 55.1 1.2 1.5 3.6
-10 concepts $10\times 90$ 207.0 1.5 2.5 3.8
-100 concepts $100\times 90$ 1750.9 2.1 11.0 5.0
-Table 5: Duration comparison (s) in erasing multiple celebrities on one A100 GPU, where $n$ is the number of target concepts.
-During data preparation, ConAbl requires pre-sampling 1,000 images ($t_{1}$ per image) for each target concept while MACE needs 8 pre-sampled images along with 8 corresponding segmentation masks ($t_{2}$ per mask) using SAM [27].
+More applications on other T2I models.
+To validate the transferability of our method across versatile applications, we conduct further experiments on various T2I models with different weights and architectures, including: (1) Composite concept erasure on DreamShaper Lykon (2023) and RealisticVision SG161222 (2023) from Fig 6 (a): Our method can precisely erase the target concept(s) while preserving other non-target elements within the prompt, such as the Van Gogh-style background (2nd column) and the Snoopy character (3rd column).
+(2) Knowledge editing on SDXL Podell et al.
+(2023) from Fig 6 (b): The arbitrary nature of anchor concepts allows us to edit the pre-trained model knowledge.
+Herein, our method effectively edits the model knowledge while maintaining the overall layout and semantics of the generated images.
+(3) Instance erasure on SDv3 Esser et al.
+(2024) from Fig 6 (c): To accommodate the diffusion transformer (DiT) Peebles & Xie (2023) architecture in T2I models, we adapt our method to a DiT-based model, demonstrating a well-balanced trade-off between erasure (1st row) and preservation (2nd row) as well.
+Table 3: Ablation study on proposed components in erasing Van Gogh, with the non-target FID averaged over the other four artistic styles from Table 1.
+Ablation 1 corresponds to the original objective from Fang et al.
+(2024) in Eq.
+3.
+The ablated components include: IEC (Invariant Equality Constraints), IPF (Influence-based Prior Filtering), RPA (Random Prior Augmentation), and DPA (Directed Prior Augmentation).
 Ablation Components Target Non-Target MS-COCO
@@
 SD v1.4 - - - - 28.75 - 26.53 -
-Table 6: Ablation study on the proposed components in erasing Van Gogh, with the non-target FID averaged over other four styles (i.e., Picasso, Monet, Paul Gauguin, Caravaggio) as in Table 2.
-Ablation 1 corresponds to the original objective from [14] in Eq.
-3.2.
-The ablated components include: IEC (Invariant Equality Constraints), IPF (Influence-based Prior Filtering), RPA (Random Prior Augmentation), and DPA (Directed Prior Augmentation).
-Figure 7: More applications across various T2I diffusion models.
-(a) We conduct both single and composite concept erasure for “Snoopy + Van Gogh” on two community T2I models [3, 4].
-(b) Our method also enables model knowledge editing by specifying the anchor concept, such as “Wonder WomanWoman” and “SupermanBatman” on SDXL [44].
-(c) Our method can seamlessly transfer to novel DiT-based T2I models (e.g., SDv3 [13]).
-More applications on other T2I models.
-To validate the transferability of our method across versatile applications, we conduct further experiments on various T2I models with different weights and architectures, including: (1) Composite concept erasure on DreamShaper [3] and RealisticVision [4] from Fig 7 (a): Our method can precisely erase the target concept(s) while preserving other non-target elements within the prompt, such as the Van Gogh-style background (2nd column) and the Snoopy character (3rd column).
-(2) Knowledge editing on SDXL [44] from Fig 7 (b): The arbitrary nature of anchor concepts allows us to edit the pre-trained model knowledge.
-Herein, our method effectively edits the model knowledge while maintaining the overall layout and semantics of the generated images.
-(3) Instance erasure on SDv3 [13] from Fig 7 (c): To accommodate the diffusion transformer (DiT) [43] architecture in T2I models, we adapt our method to a DiT-based model, demonstrating a well-balanced trade-off between erasure (1st row) and preservation (2nd row) as well.
+Component ablation.
+From Table 3, we compare the individual impact of our components on prior preservation and draw the following conclusions: (1) Impact of IEC (Ablation 1 v.s.
+2): IEC reduces the non-target FID and the MS-COCO FID, demonstrating its effectiveness by preserving invariant embeddings with equality constraints.
+(2) Impact of IPF (Ablation 2 v.s.
+3): Incorporating IPF results in a significant improvement in both FIDs, underscoring its critical role in filtering out less-influenced concepts in the retain set to mitigate semantic degradation.
+(3) Impact of DPA (Ablation 4 v.s.
+Ours): DPA improves RPA with directed noise and leads to a substantial improvement in non-target and MS-COCO FIDs, highlighting its advantage by introducing semantically similar concepts into the refined retain set.
+To conclude, the proposed three components (i.e., IEC, IPF, and DPA) improve the prior preservation from different perspectives and contribute to our method with the best prior preservation under null space constraints.
+More ablations are presented in Appx.
+D.5.
 6 Conclusion
@@
 Critically, SPEED overcomes the inefficacy of editing-based methods in multi-concept erasure while circumventing the prohibitive computational costs associated with training-based approaches.
-With our proposed Prior Knowledge Refinement involving three complementary techniques, SPEED not only ensures superior prior preservation but also achieves a 350×350$\times$350 × acceleration in multi-concept erasure, establishing itself as a scalable and practical solution for real-world applications.
+With our proposed Prior Knowledge Refinement involving three complementary techniques, SPEED not only ensures superior prior preservation but also achieves a $350\times$ acceleration in multi-concept erasure, establishing itself as a scalable and practical solution for real-world applications.
+Ethics Statement
+This work introduces a method for concept erasure in text-to-image diffusion models to address ethical concerns such as copyright infringement, privacy violations, and the generation of offensive content.
+By precisely removing specific target concepts while preserving the quality and semantics of non-target outputs, the proposed approach enhances the safety, reliability, and controllability of generative models.
+The method operates through parameter-space editing without requiring access to private data or involving human subjects, ensuring ethical integrity throughout the research process and promoting responsible deployment of generative AI technologies.
+Reproducibility Statement
+We have made extensive efforts to ensure reproducibility of our work.
+The proposed method, SPEED, is thoroughly described in the main paper (Secs.
+3 and 4), with complete theoretical derivations provided in Appx.
+B.
+Implementation details, including experimental setup details and erasure configurations, are given in Appx.
+C.
+The experimental setups for all three erasure tasks (few-concept, multi-concept, and implicit concept erasure) are described in detail, with complete quantitative results and ablation studies reported in Sec.
+5 and Appx.
+D.
 Appendix A Preliminaries
 T2I diffusion models.
-T2I generation has seen significant advancements with diffusion models, particularly Latent Diffusion Models (LDMs) [49].
+T2I generation has seen significant advancements with diffusion models, particularly Latent Diffusion Models (LDMs) Rombach et al.
+(2022).
 Unlike pixel-space diffusion, LDMs operate in the latent space of a pretrained autoencoder, reducing computational costs while maintaining high-quality synthesis.
-LDMs consist of a vector-quantized autoencoder [59, 12] and a diffusion model [11, 24, 55, 26, 57].
-The autoencoder encodes an image $\boldsymbol{x}$ into a latent representation $\boldsymbol{z}=\mathcal{E}(\boldsymbol{x})$ and reconstructs it via $\boldsymbol{x}\approx\mathcal{D}(\boldsymbol{z})$.
+LDMs consist of a vector-quantized autoencoder Van Den Oord et al.
+(2017); Esser et al.
+(2021) and a diffusion model Dhariwal & Nichol (2021); Ho et al.
+(2020); Sohl-Dickstein et al.
+(2015); Kingma et al.
+(2021); Song et al.
+(2020b).
+The autoencoder encodes an image $\bm{x}$ into a latent representation $\bm{z}=\mathcal{E}(\bm{x})$ and reconstructs it via $\bm{x}\approx\mathcal{D}(\bm{z})$.
 The diffusion model learns to generate latent codes through a denoising process.
-The training objective is given by [24, 49]:
-$\mathcal{L}_{\text{LDM}}=\mathbb{E}_{\boldsymbol{z}\sim\mathcal{E}(\boldsymbol{x}),\boldsymbol{c},\boldsymbol{\epsilon}\sim\mathcal{N}(0,1),t}\left[\left\|\boldsymbol{\epsilon}-\boldsymbol{\epsilon}_{\boldsymbol{\theta}}(\boldsymbol{z}_{t},t,\boldsymbol{c})\right\|^{2}_{2}\right],$ (13)
-where $\boldsymbol{z}_{t}$ is the noisy latent at timestep $t$, $\boldsymbol{\epsilon}$ is Gaussian noise, $\boldsymbol{\epsilon}_{\boldsymbol{\theta}}$ is the denoising network, and $\boldsymbol{c}$ is conditioning information from text, class labels, or segmentation masks [49].
-During inference, a latent $\boldsymbol{z}_{T}$ is sampled from a Gaussian prior and progressively denoised to obtain $\boldsymbol{z}_{0}$, which is then decoded into an image via $\boldsymbol{x}_{0}\approx\mathcal{D}(\boldsymbol{z}_{0})$.
+The training objective is given by Ho et al.
+(2020); Rombach et al.
+(2022):
+$\mathcal{L}_{\text{LDM}}=\mathbb{E}_{\bm{z}\sim\mathcal{E}(\bm{x}),\bm{c},\bm{\epsilon}\sim\mathcal{N}(0,1),t}\left[\left\|\bm{\epsilon}-\bm{\epsilon}_{\bm{\theta}}(\bm{z}_{t},t,\bm{c})\right\|^{2}_{2}\right],$ (13)
+where $\bm{z}_{t}$ is the noisy latent at timestep $t$, $\bm{\epsilon}$ is Gaussian noise, $\bm{\epsilon}_{\bm{\theta}}$ is the denoising network, and $\bm{c}$ is conditioning information from text, class labels, or segmentation masks Rombach et al.
+(2022).
+During inference, a latent $\bm{z}_{T}$ is sampled from a Gaussian prior and progressively denoised to obtain $\bm{z}_{0}$, which is then decoded into an image via $\bm{x}_{0}\approx\mathcal{D}(\bm{z}_{0})$.
 Cross-attention mechanisms.
-Current T2I diffusion models usually leverage a generative framework to synthesize images conditioned on textual descriptions in latent space [49].
+Current T2I diffusion models usually leverage a generative framework to synthesize images conditioned on textual descriptions in the latent space Rombach et al.
+(2022).
 The conditioning mechanism is implemented through cross-attention (CA) layers.
-Specifically, textual descriptions are first tokenized into $n$ tokens and embedded into a sequence of vectors $\boldsymbol{e}\in\mathbb{R}^{d_{0}\times n}$ via a pre-trained CLIP model [45].
+Specifically, textual descriptions are first tokenized into $n$ tokens and embedded into a sequence of vectors $\bm{e}\in\mathbb{R}^{d_{0}\times n}$ via a pre-trained CLIP model Radford et al.
+(2021).
 These text embeddings serve as the key $\mathbf{K}\in\mathbb{R}^{n\times d_{k}}$ and value $\mathbf{V}\in\mathbb{R}^{n\times d_{v}}$ inputs using parametric projection matrices $\mathbf{W_{K}}\in\mathbb{R}^{d_{k}\times d_{0}}$ and $\mathbf{W_{V}}\in\mathbb{R}^{d_{v}\times d_{0}}$, while the intermediate image representations act as the query $\mathbf{Q}\in\mathbb{R}^{m\times d_{k}}$.
@@
 1, we are tasked with minimizing the following editing objective, where the hyperparameters $\alpha$ and $\beta$ correspond to the weights of the erasure error $e_{1}$ and the preservation error $e_{0}$, respectively:
-$\mathop{\text{min}}\limits_{\boldsymbol{\Delta}}\left[\alpha\|(\mathbf{W}+\boldsymbol{\Delta})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}+\beta\|\boldsymbol{\Delta}\mathbf{C}_{0}\|^{2}\right].$ (15)
-To derive the closed-form solution, we begin by computing the gradient of the objective function with respect to $\boldsymbol{\Delta}$. The gradient is given by:
-$\alpha\left(\mathbf{W}\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}+\boldsymbol{\Delta}\mathbf{C}_{1}\right)\mathbf{C}_{1}^{\top}+\beta\boldsymbol{\Delta}\mathbf{C}_{0}\mathbf{C}_{0}^{\top}=0.$ (16)
-Solving the resulting equation yields the closed-form solution for $\boldsymbol{\Delta}_{\text{UCE}}$:
-$\boldsymbol{\Delta}_{\text{UCE}}=\alpha\mathbf{W}\left(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\right)\left(\alpha\mathbf{C}_{1}\mathbf{C}_{1}^{\top}+\beta\mathbf{C}_{0}\mathbf{C}_{0}^{\top}\right)^{-1}.$ (17)
+$\mathop{\text{min}}\limits_{\bm{\Delta}}\left[\alpha\|(\mathbf{W}+\bm{\Delta})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}+\beta\|\bm{\Delta}\mathbf{C}_{0}\|^{2}\right].$ (15)
+To derive the closed-form solution, we begin by computing the gradient of the objective function with respect to $\bm{\Delta}$. The gradient is given by:
+$\alpha\left(\mathbf{W}\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}+\bm{\Delta}\mathbf{C}_{1}\right)\mathbf{C}_{1}^{\top}+\beta\bm{\Delta}\mathbf{C}_{0}\mathbf{C}_{0}^{\top}=0.$ (16)
+Solving the resulting equation yields the closed-form solution for $\bm{\Delta}_{\text{UCE}}$:
+$\bm{\Delta}_{\text{UCE}}=\alpha\mathbf{W}\left(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\right)\left(\alpha\mathbf{C}_{1}\mathbf{C}_{1}^{\top}+\beta\mathbf{C}_{0}\mathbf{C}_{0}^{\top}\right)^{-1}.$ (17)
 In practice, an additional identity matrix $\mathbf{I}$ with hyperparameter $\lambda$ is added to $\left(\alpha\mathbf{C}_{1}\mathbf{C}_{1}^{\top}+\beta\mathbf{C}_{0}\mathbf{C}_{0}^{\top}\right)^{-1}$ to ensure its invertibility.
 This modification results in the following closed-form solution for UCE:
-$\boldsymbol{\Delta}_{\text{UCE}}=\alpha\mathbf{W}\left(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\right)\left(\alpha\mathbf{C}_{1}\mathbf{C}_{1}^{\top}+\beta\mathbf{C}_{0}\mathbf{C}_{0}^{\top}+\lambda\mathbf{I}\right)^{-1}.$ (18)
+$\bm{\Delta}_{\text{UCE}}=\alpha\mathbf{W}\left(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\right)\left(\alpha\mathbf{C}_{1}\mathbf{C}_{1}^{\top}+\beta\mathbf{C}_{0}\mathbf{C}_{0}^{\top}+\lambda\mathbf{I}\right)^{-1}.$ (18)
 B.2 Proof of the Lower Bound of $e_{0}$ for UCE
 Herein, we aim to establish the existence of a strictly positive constant $c>0$ such that
-$e_{0}=\|\boldsymbol{\Delta}_{\text{UCE}}\mathbf{C}_{0}\|^{2}=\|\alpha\mathbf{W}\left(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\right)(\alpha\mathbf{C}_{1}\mathbf{C}_{1}^{\top}+\beta\mathbf{C}_{0}\mathbf{C}_{0}^{\top}+\lambda\mathbf{I})^{-1}\mathbf{C}_{0}\|^{2}\geq c>0.$ (19)
+$e_{0}=\|\bm{\Delta}_{\text{UCE}}\mathbf{C}_{0}\|^{2}=\|\alpha\mathbf{W}\left(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\right)(\alpha\mathbf{C}_{1}\mathbf{C}_{1}^{\top}+\beta\mathbf{C}_{0}\mathbf{C}_{0}^{\top}+\lambda\mathbf{I})^{-1}\mathbf{C}_{0}\|^{2}\geq c>0.$ (19)
 Assumption B.1.
@@
 $(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top})\mathbf{B}=\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{\top}\mathbf{B}.$ (26)
-Define the projection of $\mathbf{B}$ onto the subspace spanned by the first $r$ right singular vectors (corresponding to the nonzero singular values) as
-$\mathbf{B}_{\text{proj}}=\mathbf{V}_{r}^{\top}\mathbf{B},$ (27)
-where $\mathbf{V}_{r}$ consists of the first $r$ columns of $\mathbf{V}$. Then, we can rewrite the expression as
+Define the projection of $\mathbf{B}$ onto the subspace spanned by the right singular vectors as
+$\mathbf{B}_{\text{proj}}=\mathbf{V}^{\top}\mathbf{B}.$ (27)
+Then, we can rewrite the expression as
 $(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top})\mathbf{B}=\mathbf{U}\mathbf{\Sigma}\mathbf{B}_{\text{proj}}.$ (28)
@@
 $\|\mathbf{\Sigma}\mathbf{B}_{\text{proj}}\|\geq\sigma_{r}\|\mathbf{B}_{\text{proj}}\|.$ (30)
-Next, we establish a lower bound for $\|\mathbf{B}_{\text{proj}}\|$.
-Since $\mathbf{B}$ is not entirely in the null space of $\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}$ (by assumption that the matrix is nonzero), there exists a constant $c_{1}>0$ such that
+Next, we establish a lower bound for $\|\mathbf{B}_{\text{proj}}\|$. Given that $\mathbf{V}$ is composed of right singular vectors, there exists a smallest non-zero singular value $c_{1}$ such that:
 $\|\mathbf{B}_{\text{proj}}\|\geq c_{1}\|\mathbf{B}\|.$ (31)
@@
 This completes the proof. ∎
-B.3 Deriving the Closed-Form Solution for Ours
+B.3 Deriving the Closed-Form Solution for SPEED
 From Eq. 10, we are tasked with minimizing the following editing objective:
-$\mathop{\text{min}}\limits_{\boldsymbol{\Delta}}\|(\mathbf{W}+\boldsymbol{\Delta}\mathbf{P})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}+\|\boldsymbol{\Delta}\mathbf{P}\|^{2},\quad\text{s.t.}\,(\boldsymbol{\Delta}\mathbf{P})\mathbf{C}_{2}=\mathbf{0}.$ (41)
-This is a weighted least squares problem subject to an equality constraint. To solve it, we first formulate the Lagrangian function, where $\boldsymbol{\Lambda}$ is the Lagrange multiplier:
-$\mathcal{L}(\boldsymbol{\Delta},\boldsymbol{\Lambda})=\|(\mathbf{W}+\boldsymbol{\Delta}\mathbf{P})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}+\|\boldsymbol{\Delta}\mathbf{P}\|^{2}+\boldsymbol{\Lambda}^{\top}\left((\boldsymbol{\Delta}\mathbf{P})\mathbf{C}_{2}\right).$ (42)
-We compute the gradient of the Lagrangian function in Eq. 42 with respect to $\boldsymbol{\Delta}$ and set it to zero, yielding the following equation for $\boldsymbol{\Delta}$:
-$\frac{\partial\mathcal{L}(\boldsymbol{\Delta},\boldsymbol{\Lambda})}{\partial\boldsymbol{\Delta}}=2\left((\mathbf{W}+\boldsymbol{\Delta}\mathbf{P})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\right)\mathbf{C}_{1}^{\top}\mathbf{P}^{\top}+2\boldsymbol{\Delta}\mathbf{P}\mathbf{P}^{\top}+\boldsymbol{\Lambda}\mathbf{C}_{2}^{\top}\mathbf{P}^{\top}=\mathbf{0}.$ (43)
+$\mathop{\text{min}}\limits_{\bm{\Delta}}\|(\mathbf{W}+\bm{\Delta}\mathbf{P})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}+\|\bm{\Delta}\mathbf{P}\|^{2},\quad\text{s.t.}\,(\bm{\Delta}\mathbf{P})\mathbf{C}_{2}=\mathbf{0}.$ (41)
+This is a weighted least squares problem subject to an equality constraint. To solve it, we first formulate the Lagrangian function, where $\bm{\Lambda}$ is the Lagrange multiplier:
+$\mathcal{L}(\bm{\Delta},\bm{\Lambda})=\|(\mathbf{W}+\bm{\Delta}\mathbf{P})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}+\|\bm{\Delta}\mathbf{P}\|^{2}+\bm{\Lambda}^{\top}\left((\bm{\Delta}\mathbf{P})\mathbf{C}_{2}\right).$ (42)
+We compute the gradient of the Lagrangian function in Eq. 42 with respect to $\bm{\Delta}$ and set it to zero, yielding the following equation for $\bm{\Delta}$:
+$\frac{\partial\mathcal{L}(\bm{\Delta},\bm{\Lambda})}{\partial\bm{\Delta}}=2\left((\mathbf{W}+\bm{\Delta}\mathbf{P})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\right)\mathbf{C}_{1}^{\top}\mathbf{P}^{\top}+2\bm{\Delta}\mathbf{P}\mathbf{P}^{\top}+\bm{\Lambda}\mathbf{C}_{2}^{\top}\mathbf{P}^{\top}=\mathbf{0}.$ (43)
 Given that the projection matrix $\mathbf{P}$ is derived from $R_{\text{refine}}$ using Eq.
 2, $\mathbf{P}$ is a symmetric matrix (i.e., $\mathbf{P}=\mathbf{P}^{\top}$) and an idempotent matrix (i.e., $\mathbf{P}^{2}=\mathbf{P}$), the above formulation can be simplified to:
-$\frac{\partial\mathcal{L}(\boldsymbol{\Delta},\boldsymbol{\Lambda})}{\partial\boldsymbol{\Delta}}=2\left((\mathbf{W}+\boldsymbol{\Delta}\mathbf{P})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\right)\mathbf{C}_{1}^{\top}\mathbf{P}+2\boldsymbol{\Delta}\mathbf{P}+\boldsymbol{\Lambda}\mathbf{C}_{2}^{\top}\mathbf{P}=\mathbf{0}.$ (44)
-Therefore, we can obtain the closed-form solution for $\boldsymbol{\Delta}\mathbf{P}$ from this equation:
-$\boldsymbol{\Delta}\mathbf{P}=(\mathbf{W}\mathbf{C}_{*}\mathbf{C}_{1}^{\top}\mathbf{P}-\mathbf{W}\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\mathbf{P}-\frac{1}{2}\boldsymbol{\Lambda}\mathbf{C}_{2}^{\top}\mathbf{P})(\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\mathbf{P}+\mathbf{I})^{-1}.$ (45)
-Next, we differentiate the Lagrangian function in Eq. 42 with respect to $\boldsymbol{\Lambda}$ and set it to zero:
-$\frac{\partial\mathcal{L}(\boldsymbol{\Delta},\boldsymbol{\Lambda})}{\partial\boldsymbol{\Lambda}}=(\boldsymbol{\Delta}\mathbf{P})\mathbf{C}_{2}=\mathbf{0}.$ (46)
+$\frac{\partial\mathcal{L}(\bm{\Delta},\bm{\Lambda})}{\partial\bm{\Delta}}=2\left((\mathbf{W}+\bm{\Delta}\mathbf{P})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\right)\mathbf{C}_{1}^{\top}\mathbf{P}+2\bm{\Delta}\mathbf{P}+\bm{\Lambda}\mathbf{C}_{2}^{\top}\mathbf{P}=\mathbf{0}.$ (44)
+Therefore, we can obtain the closed-form solution for $\bm{\Delta}\mathbf{P}$ from this equation:
+$\bm{\Delta}\mathbf{P}=(\mathbf{W}\mathbf{C}_{*}\mathbf{C}_{1}^{\top}\mathbf{P}-\mathbf{W}\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\mathbf{P}-\frac{1}{2}\bm{\Lambda}\mathbf{C}_{2}^{\top}\mathbf{P})(\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\mathbf{P}+\mathbf{I})^{-1}.$ (45)
+Next, we differentiate the Lagrangian function in Eq. 42 with respect to $\bm{\Lambda}$ and set it to zero:
+$\frac{\partial\mathcal{L}(\bm{\Delta},\bm{\Lambda})}{\partial\bm{\Lambda}}=(\bm{\Delta}\mathbf{P})\mathbf{C}_{2}=\mathbf{0}.$ (46)
 For simplicity, we define $\mathbf{M}=(\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\mathbf{P}+\mathbf{I})^{-1}$. Then, we substitute the result of Eq.45 into Eq.46 and obtain:
-$(\mathbf{W}\mathbf{C}_{*}\mathbf{C}_{1}^{\top}\mathbf{P}-\mathbf{W}\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\mathbf{P}-\frac{1}{2}\boldsymbol{\Lambda}\mathbf{C}_{2}^{\top}\mathbf{P})\mathbf{M}\mathbf{C}_{2}=\mathbf{0}.$ (47)
+$(\mathbf{W}\mathbf{C}_{*}\mathbf{C}_{1}^{\top}\mathbf{P}-\mathbf{W}\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\mathbf{P}-\frac{1}{2}\bm{\Lambda}\mathbf{C}_{2}^{\top}\mathbf{P})\mathbf{M}\mathbf{C}_{2}=\mathbf{0}.$ (47)
 Solving this equation leads to:
-$\frac{1}{2}\boldsymbol{\Lambda}=\mathbf{W}(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top})\mathbf{P}\mathbf{M}\mathbf{C}_{2}(\mathbf{C}_{2}^{\top}\mathbf{P}\mathbf{M}\mathbf{C}_{2})^{-1}.$ (48)
+$\frac{1}{2}\bm{\Lambda}=\mathbf{W}(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top})\mathbf{P}\mathbf{M}\mathbf{C}_{2}(\mathbf{C}_{2}^{\top}\mathbf{P}\mathbf{M}\mathbf{C}_{2})^{-1}.$ (48)
 Substituting Eq.48 back into Eq.45, we have the closed-form solution of our objective:
-$(\boldsymbol{\Delta}\mathbf{P})_{\text{SPEED}}=\mathbf{W}(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top})\mathbf{P}\mathbf{Q}\mathbf{M},$ (49)
+$(\bm{\Delta}\mathbf{P})_{\text{SPEED}}=\mathbf{W}(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top})\mathbf{P}\mathbf{Q}\mathbf{M},$ (49)
 where $\mathbf{Q}=\mathbf{I}-\mathbf{M}\mathbf{C}_{2}(\mathbf{C}_{2}^{\top}\mathbf{P}\mathbf{M}\mathbf{C}_{2})^{-1}\mathbf{C}_{2}^{\top}\mathbf{P}$ and $\mathbf{M}=(\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\mathbf{P}+\mathbf{I})^{-1}$.
 Appendix C Implementation Details
+Table 4: Evaluation setup for multi-concept erasure.
+This dataset contains an erasure set with 100 celebrities and a retain set with another 100 celebrities.
+We experiment with erasing 10, 50, and 100 celebrities with the predefined target concepts and the entire retain set is utilized in all cases.
 Group Number Anchor Concept Celebrity
@@
 Retain Set 10, 50, and 100 - ‘Aaron Paul’, ‘Alec Baldwin’, ‘Amanda Seyfried’, ‘Amy Poehler’, ‘Amy Schumer’, ‘Amy Winehouse’, ‘Andy Samberg’, ‘Aretha Franklin’, ‘Avril Lavigne’, ‘Aziz Ansari’, ‘Barry Manilow’, ‘Ben Affleck’, ‘Ben Stiller’, ‘Benicio Del Toro’, ‘Bette Midler’, ‘Betty White’, ‘Bill Murray’, ‘Bill Nye’, ‘Britney Spears’, ‘Brittany Snow’, ‘Bruce Lee’, ‘Burt Reynolds’, ‘Charles Manson’, ‘Christie Brinkley’, ‘Christina Hendricks’, ‘Clint Eastwood’, ‘Countess Vaughn’, ‘Dakota Johnson’, ‘Dane Dehaan’, ‘David Bowie’, ‘David Tennant’, ‘Denise Richards’, ‘Doris Day’, ‘Dr Dre’, ‘Elizabeth Taylor’, ‘Emma Roberts’, ‘Fred Rogers’, ‘Gal Gadot’, ‘George Bush’, ‘George Takei’, ‘Gillian Anderson’, ‘Gordon Ramsey’, ‘Halle Berry’, ‘Harry Dean Stanton’, ‘Harry Styles’, ‘Hayley Atwell’, ‘Heath Ledger’, ‘Henry Cavill’, ‘Jackie Chan’, ‘Jada Pinkett Smith’, ‘James Garner’, ‘Jason Statham’, ‘Jeff Bridges’, ‘Jennifer Connelly’, ‘Jensen Ackles’, ‘Jim Morrison’, ‘Jimmy Carter’, ‘Joan Rivers’, ‘John Lennon’, ‘Johnny Cash’, ‘Jon Hamm’, ‘Judy Garland’, ‘Julianne Moore’, ‘Justin Bieber’, ‘Kaley Cuoco’, ‘Kate Upton’, ‘Keanu Reeves’, ‘Kim Jong Un’, ‘Kirsten Dunst’, ‘Kristen Stewart’, ‘Krysten Ritter’, ‘Lana Del Rey’, ‘Leslie Jones’, ‘Lily Collins’, ‘Lindsay Lohan’, ‘Liv Tyler’, ‘Lizzy Caplan’, ‘Maggie Gyllenhaal’, ‘Matt Damon’, ‘Matt Smith’, ‘Matthew Mcconaughey’, ‘Maya Angelou’, ‘Megan Fox’, ‘Mel Gibson’, ‘Melanie Griffith’, ‘Michael Cera’, ‘Michael Ealy’, ‘Natalie Portman’, ‘Neil Degrasse Tyson’, ‘Niall Horan’, ‘Patrick Stewart’, ‘Paul Rudd’, ‘Paul Wesley’, ‘Pierce Brosnan’, ‘Prince’, ‘Queen Elizabeth’, ‘Rachel Dratch’, ‘Rachel Mcadams’, ‘Reba Mcentire’, ‘Robert De Niro’
-Table 7: The evaluation setup for multi-concept erasure.
-This celebrity dataset contains an erasure set with 100 celebrities and a retain set with another 100 celebrities.
-We experiment with erasing 10, 50, and 100 celebrities with the pre-defined target concepts and the entire retain set is utilized in all cases.
+Figure 7: Qualitative demonstration of our erasure performance across (a) instance erasure, (b) artistic style erasure, (c) celebrity erasure, and (d) implicit concept erasure.
+Our method achieves precise erasure efficacy across various scenarios while exhibiting superior prior preservation.
+The corresponding CS is highlighted in blue, indicating that successful erasure can be achieved without pushing CS much lower, as our results demonstrate sufficient erasure at a moderate level.
 C.1 Experimental Setup Details
-[…]
+Few-concept erasure.
+We first compare methods on few-concept erasure, a fundamental concept erasure task, including both instance erasure and artistic style erasure following Lyu et al.
+(2024).
+For instance erasure, we prepare 80 instance templates proposed in CLIP Radford et al.
+(2021), such as “a photo of the {Instance}, “a drawing of the {Instance}, and “a painting of the {Instance}.
+For artistic style erasure, we use ChatGPT OpenAI (2022); Achiam et al.
+(2023) to generate 30 artistic style templates, including {Artistic} style painting of the night sky with bold strokes”, {Artistic} style landscape of rolling hills with dramatic brushwork”, and “Sunrise scene in {Artistic} style, capturing the beauty of dawn”.
+Following Lyu et al.
+(2024), we handpick the representative target and anchor concepts as the erasure set (i.e., Snoopy, Mickey, SpongeBob $\to$ ‘ ’ in instance erasure and Van Gogh, Picasso, Monet $\to$ ‘art’ in artistic style erasure) and non-target concepts for evaluation (i.e., Pikachu and Hello Kitty in instance erasure and Paul Gauguin and Caravaggio in artistic style erasure).
+In terms of the retain set, for instance erasure, we use a scraping script to crawl Wikipedia category pages to extract fictional character names and their page view counts with a threshold of 500,000 views from 2020.01.01 to 2023.12.31, resulting in 1,352 instances.
+For artistic style erasure, we use the 1,734 artistic styles collected from UCE Gandikota et al.
+(2024).
+In evaluation, we generate 10 images per template per concept, resulting in 800 and 300 images for each concept in instance erasure and artistic style erasure, respectively.
+Moreover, we introduce the MS-COCO captions Lin et al.
+(2014) to serve as general prior knowledge.
+In implementation, we use the first 1,000 captions to generate a total of 1000 images to compare CS and FID before and after erasure.
 Multi-concept erasure.
 We then compare methods on multi-concept erasure, a more challenging and realistic concept erasure task.
-Following the experiment setup from [35], we introduce a dataset consisting of 200 celebrities, where their portraits generated by SDv1.4 [1] can be recognizable with exceptional accuracy by the GIPHY Celebrity Detector (GCD) [20].
+Following the experiment setup from Lu et al.
+(2024a), we introduce a dataset consisting of 200 celebrities, where their portraits generated by SDv1.4 AI (2022) can be recognizable with exceptional accuracy by the GIPHY Celebrity Detector (GCD) Hasty et al.
+(2019).
 This dataset is divided into two groups: an erasure set with 10, 50, and 100 celebrities and a retain set with 100 other celebrities.
 The full list for both sets is presented in Table 4.
-We experiment with erasing 10, 50, and 100 celebrities with the pre-defined target concepts and the entire retain set is utilized in all cases.
+We experiment with erasing 10, 50, and 100 celebrities with the predefined target concepts and the entire retain set is utilized in all cases.
 In evaluation, we prepare five celebrity templates, (i.e., “a portrait of {Celebrity}”, “a sketch of {Celebrity}”, “an oil painting of {Celebrity}”, “{Celebrity} in an official photo”, and “an image capturing {Celebrity} at a public event”) and generate 500 images for both sets.
@@
 For target concepts, we adjust the per-concept quantity to maintain a total of 500 images (e.g., erasing 10 celebrities involves generating 10 images with 5 templates).
-Implicit concept erasure.
-We adopt the same setting in [19] to erase nudity $\to$ ‘ ’ as the erasure set and ‘ ’ as the retain set.
-In evaluation, we generate images using all 4,703 prompts in I2P and use NudeNet [6] to identify nude content with the threshold of 0.6.
+C.2 Erasure Configurations
+Implementation of previous works.
+In our series of three concept erasure tasks, we mainly compare against four methods: ConAbl [^4] https://github.com/nupurkmr9/concept-ablation Kumari et al.
+(2023), MACE [^5] https://github.com/Shilin-LU/MACE Lu et al.
+(2024a), RECE [^6] https://github.com/CharlesGong12/RECE Gong et al.
+(2025), and UCE [^7] https://github.com/rohitgandikota/unified-concept-editing Gandikota et al.
+(2024), as they achieve SOTA performance across different concept erasure tasks.
+All the compared methods are implemented using their default configurations from the corresponding official repositories.
+One exception is that for MACE when erasing 50 celebrities, since it doesn’t provide an official configuration and the preserve weight varies with the number of target celebrities, we set it to $1.2\times 10^{5}$ to ensure a consistent balance between erasure and preservation.
+Implementation of SPEED.
+In line with previous methods Kumari et al.
+(2023); Lu et al.
+(2024a); Gong et al.
+(2025); Gandikota et al.
+(2024), we edit the cross-attention (CA) layers within the diffusion model due to their role in text-image alignment Hertz et al.
+(2022).
+In contrast, we only edit the value matrices in the CA layers, as suggested by Wang et al.
+(2024b).
+This choice is grounded in the observation that the keys in CA layers typically govern the layout and compositional structure of the attention map, while the values control the content and visual appearance of the images Tewel et al.
+(2023).
+In the context of concept erasure, our goal is to effectively remove the semantics of the target concept, and we find that only editing the value matrices is sufficient as shown in Fig.
+4 and 5 (further ablation comparison is provided in Appx.
+D.5).
+The augmentation times $N_{A}$ in Eq.
+9 is set to 10 and the augmentation ranks $r$ in Eq.
+7 is set to 1 as ablated in Appx.
+D.5.
+Meanwhile, given that eigenvalues are rarely strictly zero in practical applications when determining the null space, we select the singular vectors corresponding to the singular values below $10^{-1}$ on few-concept and implicit concept erasure and $10^{-4}$ on multi-concept erasure following Fang et al.
+(2024).
+Moreover, since the retain set only includes ‘ ’ in implicit concept erasure, we add an identity matrix $\mathbf{I}$ with weight $\lambda=0.5$ to the term $(\mathbf{C}_{2}^{\top}\mathbf{P}\mathbf{M}\mathbf{C}_{2})^{-1}$ in Eq.
+12 to ensure invertibility.
+Table 5: Complete quantitative comparison of the few-concept erasure in erasing instances from Table 1 (left).
+The best results are highlighted in bold, and grey columns are indirect indicators for measuring erasure efficacy on target concepts or prior preservation on non-target concepts.
 Snoopy Mickey Spongebob Pikachu Hello Kitty MS-COCO
@@
 UCE 23.19 102.86 26.64 24.87 27.29 29.86 27.47 19.06 27.75 27.86 26.46 22.18
+SPM 23.72 116.26 26.62 31.21 27.21 31.96 27.41 19.82 27.80 30.95 26.47 20.71
+SPM w/o FT 23.72 116.26 26.55 43.03 26.84 42.96 27.38 25.95 27.71 42.53 26.48 20.86
 Ours 23.50 108.51 26.67 23.41 27.31 24.64 27.48 16.81 27.82 21.74 26.48 19.95
@@
 UCE 23.60 99.30 24.79 86.32 27.32 30.58 27.38 23.51 27.74 31.76 26.38 26.06
+SPM 23.18 122.17 22.71 117.30 26.92 38.35 27.35 27.13 27.76 39.61 26.45 24.59
+SPM w/o FT 22.45 127.95 21.77 127.57 25.96 61.52 27.39 42.63 27.14 68.75 26.43 23.82
 Ours 23.58 103.62 23.62 83.70 27.34 29.67 27.39 22.51 27.78 28.23 26.47 23.66
@@
 UCE 23.29 101.40 24.63 88.11 19.08 140.40 27.45 29.20 27.82 38.15 26.30 28.71
+SPM 22.86 125.66 22.08 123.20 20.92 153.36 27.45 37.51 27.63 46.63 26.48 25.47
+SPM w/o FT 21.80 137.98 20.86 139.48 20.19 163.21 26.68 66.15 26.24 85.35 26.33 25.05
 Ours 23.69 103.33 23.93 86.55 21.39 109.28 27.47 21.40 27.76 26.22 26.51 24.99
-Table 8: Full quantitative comparison of the few-concept erasure in erasing instances from Table 1.
-The erased target concept(s) are shaded in pink, the best results are highlighted in bold, and the grey columns are indirect indicators for measuring erasure efficacy on target concepts or prior preservation on non-target concepts.
+Table 6: Complete quantitative comparison of the few-concept erasure in erasing artistic styles from Table 1 (right).
+The best results are highlighted in bold, and grey columns are indirect indicators for measuring erasure efficacy on target concepts or prior preservation on non-target concepts.
 Van Gogh Picasso Monet Paul Gauguin Caravaggio MS-COCO
@@
 UCE 28.10 133.87 27.70 43.02 28.92 40.49 29.62 32.62 26.23 61.72 26.54 19.63
+ESD-X 27.04 200.05 26.50 111.07 28.14 90.35 29.45 106.70 25.70 107.85 26.10 33.19
+ESD-U 26.24 205.06 26.28 153.10 27.79 105.78 29.59 164.83 26.14 124.41 26.35 38.08
+RACE 23.03 233.25 25.54 127.28 26.44 94.49 27.78 106.43 25.08 114.94 25.92 41.52
+Receler 21.53 245.40 24.88 134.35 23.61 143.17 25.02 194.58 24.52 133.94 25.95 37.00
 Ours 26.29 131.02 27.96 35.86 28.94 16.85 29.71 24.94 26.24 39.75 26.55 20.36
@@
 UCE 28.72 37.58 26.99 102.21 28.92 16.72 29.71 32.48 26.22 59.27 26.50 20.33
+ESD-X 28.58 104.48 26.07 178.18 28.32 62.79 29.31 96.70 25.84 100.54 26.15 34.12
+ESD-U 28.69 109.39 26.47 156.35 28.64 67.69 29.64 95.39 26.04 105.76 26.35 35.78
+RACE 28.12 112.29 24.84 185.78 27.88 72.79 28.91 93.19 25.81 110.23 25.77 42.01
+Receler 25.92 199.56 23.10 243.28 26.92 94.89 26.51 208.01 25.34 135.35 25.88 37.20
 Ours 28.76 19.18 26.22 117.71 28.88 19.87 29.75 24.73 26.24 43.63 26.51 19.98
@@
 UCE 28.65 42.25 27.91 38.73 27.12 98.37 29.58 33.00 26.16 56.49 26.51 21.58
+ESD-X 28.15 115.51 26.56 92.69 25.97 124.90 28.85 89.07 25.92 102.53 25.98 35.79
+ESD-U 28.73 134.10 26.87 114.64 25.15 134.02 29.44 135.64 25.72 131.90 26.21 38.16
+RACE 27.13 132.42 25.99 106.70 23.08 149.16 27.52 98.71 24.96 110.34 25.81 41.96
+Receler 24.94 169.55 26.16 105.24 21.06 182.34 24.81 199.23 25.03 122.42 25.99 36.39
 Ours 28.76 28.78 27.93 41.21 25.06 134.11 29.66 27.85 26.22 55.20 26.48 20.87
-Table 9: Full quantitative comparison of the few-concept erasure in erasing styles from Table 2.
-The erased target concept(s) are shaded in pink, the best results are highlighted in bold, and the grey columns are indirect indicators for measuring erasure efficacy on target concepts or prior preservation on non-target concepts.
-Concept Snoopy Mickey Spongebob Pikachu Hello Kitty
-CS FID CS FID CS FID CS FID CS FID
-SD v1.4 28.51 - 26.62 - 27.30 - 27.44 - 27.77 -
-Erase Snoopy
-CS ↓ FID ↑ CS ↑ FID ↓ CS ↑ FID ↓ CS ↑ FID ↓ CS ↑ FID ↓
-SPM 23.72 116.26 26.62 31.21 27.21 31.96 27.41 19.82 27.80 30.95
-SPM w/o FT 23.72 116.26 26.55 43.03 26.84 42.96 27.38 25.95 27.71 42.53
-Ours 23.50 108.51 26.67 23.41 27.31 24.64 27.42 16.81 27.82 21.74
-Erase Snoopy and Mickey
-CS ↓ FID ↑ CS ↓ FID ↑ CS ↑ FID ↓ CS ↑ FID ↓ CS ↑ FID ↓
-SPM 23.18 122.17 22.71 117.30 26.92 38.35 27.35 27.13 27.76 39.61
-SPM w/o FT 26.66 83.11 25.64 75.46 26.73 42.30 27.50 28.19 27.75 41.19
-Ours 23.58 103.62 23.62 83.70 27.24 29.67 27.39 22.51 27.78 28.23
-Erase Snoopy and Mickey and Spongebob
-CS ↓ FID ↑ CS ↓ FID ↑ CS ↓ FID ↑ CS ↑ FID ↓ CS ↑ FID ↓
-SPM 22.86 125.66 22.08 123.20 20.92 153.36 27.50 37.51 27.63 46.63
-SPM w/o FT 27.75 68.89 26.24 63.30 24.82 77.38 27.44 30.00 27.74 38.89
-Ours 23.69 103.33 23.93 86.55 21.39 109.28 27.47 21.40 27.76 26.22
-Table 10: Quantitative comparison with SPM and SPM w/o FT (Facilitated Transport).
-The erased target concept(s) are shaded in pink, the best results are highlighted in bold, and the grey columns are indirect indicators for measuring erasure efficacy on target concepts or prior preservation on non-target concepts.
-C.2 Erasure Configurations
-Implementation of previous works.
-In our series of three concept erasure tasks, we mainly compare against four methods: ConAblhttps://github.com/nupurkmr9/concept-ablation [29], MACE§§§https://github.com/Shilin-LU/MACE [35], RECEhttps://github.com/CharlesGong12/RECE [19], and UCEhttps://github.com/rohitgandikota/unified-concept-editing [18], as they achieve SOTA performance across different concept erasure tasks.
-All the compared methods are implemented using their default configurations from the corresponding official repositories.
-One exception is that for MACE when erasing 50 celebrities, since it doesn’t provide an official configuration and the preserve weight varies with the number of target celebrities, we set it to $1.2\times 10^{5}$ to ensure a consistent balance between erasure and preservation.
-Implementation of SPEED.
-In line with previous methods [29, 35, 19, 18], we edit the cross-attention (CA) layers within the diffusion model due to their role in text-image alignment [21].
-In contrast, we only edit the value matrices in the CA layers, as suggested by [62].
-This choice is grounded in the observation that the keys in CA layers typically govern the layout and compositional structure of the attention map, while the values control the content and visual appearance of the images [58].
-In the context of concept erasure, our goal is to effectively remove the semantics of the target concept, and we find that only editing the value matrices is sufficient as shown in Fig.
-5 and 6 (further ablation comparison is provided in Appx.
-D.3).
-The augmentation times $N_{A}$ in Eq.
-8 is set to 10 and the augmentation ranks $r$ in Eq.
-6 is set to 1 as ablated in Appx.
-D.3.
-Meanwhile, given that eigenvalues are rarely strictly zero in practical applications when determining the null space, we select the eigenvectors corresponding to the eigenvalues below $10^{-1}$ on few-concept and implicit concept erasure and $10^{-4}$ on multi-concept erasure following [14].
-Moreover, since the retain set only includes ‘ ’ in implicit concept erasure, we add an identity matrix $\mathbf{I}$ with weight $\lambda=0.5$ to the term $(\mathbf{C}_{2}^{\top}\mathbf{P}\mathbf{M}\mathbf{C}_{2})^{-1}$ in Eq.
-11 to ensure invertibility following [18].
+Table 7: Quantitative comparison with SPM and SPM w/o FT in multi-concept erasure.
+The best results are highlighted in bold.
+Our method is capable of erasing up to 100 celebrities at once with low $\text{Acc}_{e}$ (%) and preserving other non-target celebrities with less appearance alteration with high $\text{Acc}_{r}$ (%), resulting in the best overall erasure performance $H_{o}$ (shaded in pink).
+FAIL indicates that the model collapses with noisy generations ($\text{Acc}_{e}=\text{Acc}_{r}=0.00\%$).
+Erase 10 Celebrities MS-COCO Erase 50 Celebrities MS-COCO Erase 100 Celebrities MS-COCO
+$\text{Acc}_{e}$ $\text{Acc}_{r}$ $H_{o}$ CS ↑ FID ↓ $\text{Acc}_{e}$ $\text{Acc}_{r}$ $H_{o}$ CS ↑ FID ↓ $\text{Acc}_{e}$ $\text{Acc}_{r}$ $H_{o}$ CS ↑ FID ↓
+SD v1.4 91.99 89.66 14.70 26.53 - 93.08 89.66 12.85 26.53 - 90.18 89.66 17.70 26.53 -
+SPM 0.00 51.79 68.24 26.42 48.44 0.00 0.00 FAIL 26.32 52.61 0.00 0.00 FAIL 25.15 63.20
+SPM w/o FT 0.00 5.08 9.68 26.38 52.23 0.00 0.00 FAIL 16.22 170.68 0.00 0.00 FAIL 14.34 245.92
+Ours 1.81 89.09 93.42 26.47 30.02 3.46 88.48 92.34 26.46 39.23 5.87 85.54 89.63 26.22 44.97
 Appendix D Additional Experiments
-D.1 Full Comparison on Few-Concept Erasure
-We present full quantitative comparisons of few-concept erasure, including both CS and FID, in Table 8 and Table 9.
+D.1 More Demonstrations
+We further provide qualitative visualizations of the erasure results in Fig.7, illustrating the effectiveness of our method in performing precise and targeted concept erasure across diverse scenarios.
+Specifically, we showcase: (a) instance erasure from Table 1 (left); (b) artistic style erasure from Table 1 (right); (c) celebrity erasure from Table 2; and (d) implicit concept erasure (e.g., nudity) from Table 8.
+In all cases, our method successfully removes the intended concept while preserving unrelated content, demonstrating its universal erasure applications.
+We also evaluate the CLIP score (CS) before and after concept erasure to assess the erasure efficacy.
+As shown in Fig.
+8, our method achieves successful erasure of specific concepts such as Snoopy and Mickey while maintaining moderate CS values (24.18 and 23.44, respectively).
+This indicates that effective erasure does not require minimizing CS to an extreme.
+In contrast, RECE obtains the lowest CS (19.79 and 18.75), but this is achieved at the cost of overly aggressive erasure.
+For example, transforming Snoopy into an unrecognizable image and replacing Mickey with a generic human figure.
+While such strategies may enhance erasure efficacy, they also risk compromising prior knowledge.
+This trade-off is reflected in higher non-target FIDs, as shown in Tables 1 and 2.
+D.2 Complete Results on Few-Concept Erasure
+We present complete quantitative comparisons of few-concept erasure, including both CS and FID, in Table 5 and Table 6.
 Our results demonstrate that our method consistently achieves superior prior preservation, as indicated by higher CS and lower FID across the majority of non-target concepts.
-D.2 More Baselines
-In this section, we compare against more methods because of the page limit in our main paper.
-Since our method focuses on improving prior preservation and multi-concept erasure performance, we mainly compare it with similar methods, other methods like ESD [17], FMN [66], and SLD [51] are omitted, as they fail to achieve satisfactory prior preservation proved by previous comaprisons [37, 35, 62].
-The remaining comparable method is SPM******https://github.com/Con6924/SPM [37], which is proposed to improve prior preservation and can scale to multi-concept erasure tasks.
-Notable, SPM not only fine-tunes the model weights using LoRA [25] but also intervenes in the image generation process through Facilitated Transport.
+Figure 8: Comparison of CLIP scores (CS) across different erasure methods.
+We compare the results in erasing Snoopy and Mickey, and highlight the corresponding CS in blue.
+Our method achieves successful concept erasure with moderate CS values.
+In contrast, RECE achieves the lowest CS by enabling more aggressive erasure.
+For example, removing Snoopy to the extent of producing a semantically void image, and changing Mickey into a generic person.
+We argue that such over-erasure unnecessarily compromises prior preservation as evidenced by Tables 1 and 2.
+Figure 9: Qualitative comparison with SPM and SPM w/o FT in erasing single and multiple instances.
+The erased and preserved generations are highlighted with red and green boxes, respectively.
+Our method demonstrates superior prior preservation compared to both baselines.
+Meanwhile, without Facilitated Transport, SPM w/o FT shows poorer prior preservation in multi-concept erasure (e.g., marked by Refer to caption) with significant semantic changes compared to original generations.
+D.3 Comparison on More Baselines
+In this section, we compare against more methods because of the page limit in our main paper, including ESD [^8] https://github.com/rohitgandikota/erasing Gandikota et al.
+(2023), RACE [^9] https://github.com/chkimmmmm/R.A.C.E.
+Kim et al.
+(2024), Receler [^10] https://github.com/jasper0314-huang/Receler Huang et al.
+(2024), and SPM [^11] https://github.com/Con6924/SPM Lyu et al.
+(2024).
+While the first three methods are training-based, focusing solely on modifying model parameters, SPM not only fine-tunes the model weights using LoRA Hu et al.
+(2021) but also intervenes in the image generation process through Facilitated Transport.
 Specifically, this module dynamically adjusts the LoRA scale based on the similarity between the sampling prompt and the target concept.
@@
 In the latter version, the LoRA scale is set to 1 by default.
-The quantitative comparative results are shown in Table 10.
-It can be seen that our method consistently achieves the best prior preservation compared to both SPM and SPM w/o FT.
-Even equipped with Facilitated Transport, our method achieves the lowest non-target FID (e.g., on Pikachu and Hello Kitty).
-This superiority amplifies as the number of target concepts increases.
+The quantitative comparative results are shown in Tables 5 and 6, where our method consistently achieves the best prior preservation compared to all compared baselines.
+Even equipped with Facilitated Transport (i.e., SPM w/ FT), our method achieves the lowest non-target FID (e.g., on Pikachu and Hello Kitty).
+This superiority amplifies as the number of target concepts increases as shown in Table 7.
 For example, with the number of target concepts increasing from 1 to 3, our FID in Pikachu rises from 16.81 to 21.40 (4.59 ↑), while SPM increases from 19.82 to 37.51 (17.69 ↑), where a similar pattern is observed in Hello Kitty (Our 4.48 ↑ v.s.
 SPM’s 15.68 ↑).
-Once removing the Facilitated Transport module, SPM w/o FT shows poor erasure efficacy in multi-concept erasure (highlighted by red in Table 10 and illustrated in Fig.
-8).
-This indicates that the success of SPM in multi-concept erasure relies on the Facilitated Transport module, which dynamically allocates the LoRA scales by calculating the similarity between the sampling prompt and each target concept.
-For example, when erasing Snoopy + Mickey + SpongeBob, if the sampling prompt is “a photo of Snoopy”, SPM will allocate a larger scale to Snoopy’s LoRA according to the text similarity.
-On the contrary, if the sampling prompt is “a photo of Pikachu” with the non-target concept, all three LoRA scales will be assigned lower values, thereby preserving the prior knowledge.
-We argue that this strategy of dynamically tuning the LoRA scales based on the sampling prompt similarity is vulnerable to attacks and easily bypassed, especially in white-box attack scenarios, where an attacker can reconstruct the erased concepts by simply modifying the code with extremely low attack costs, e.g., open-source T2I models like Stable Diffusion [1, 2].
-Figure 8: Qualitative comparison with SPM and SPM w/o FT in erasing single and multiple instances.
-The erased and preserved generations are highlighted with red and green boxes, respectively.
-Our method demonstrates superior prior preservation compared to both SPM and SPM w/o FT.
-Meanwhile, without the Facilitated Transport module, SPM w/o FT shows poor erasure efficacy in multi-concept erasure (e.g., the 1st case in Snoopy, 2nd case in erasing Mickey, and 3rd case in SpongeBob) with relatively high target CS in Table 10.
-D.3 Ablation Studies
-Augmentation times.
-We ablate the augmentation times $N_{A}$ proposed in the Directed Prior Augmentation (DPA) module in Sec.
-4.2, which controls the balance between rank saturation and prior coverage along with the Influence-based Prior Filtering (IPF) module.
-It can be observed from Fig.
-9 (a) that: (1) As $N_{A}$ increases, the non-target FID exhibits a trend of first decreasing and then increasing.
-This suggests that when $N_{A}$ is small (i.e., $1\to 10$), augmenting existing non-target concepts with semantically similar concepts facilitates a more comprehensive prior coverage, thereby improving prior preservation.
-However, when $N_{A}$ exceeds a certain threshold (i.e., $10\to 20$), further augmentation of non-target concepts leads to narrowing the null-space derivation due to rank saturation, ultimately degrading prior preservation.
-(2) Target CS generally shows a declining trend, indicating that the proposed Prior Knowledge Refinement strategy not only improves prior preservation but also exerts a positive impact on erasure efficacy.
-Augmentation ranks.
-Another hyperparameter to be ablated is the augmentation ranks $r$.
-From Eq.
-6, we introduce the number of the smallest singular values, i.e., augmentation ranks $r$ in deriving $\mathbf{P}_{\text{min}}=\mathbf{U}_{\text{min}}\mathbf{U}_{\text{min}}^{\top}$ with $\mathbf{U}_{\text{min}}={\mathbf{U}}_{\mathbf{W}}[:,-r:]$.
-Mathematically, $r$ represents the directions in which the DPA module can augment in the concept embedding space and constrains the rank of the augmented embeddings to a maximum of $r$.
-As shown in Fig.
-9 (b), as $r$ increases, the non-target FID exhibits an overall upward trend, indicating that introducing more ranks does not benefit prior preservation, as it narrows the null space.
-At the same time, as shown in Table 6, such augmentation by DPA also remains necessary, as it enables more comprehensive coverage of non-target knowledge with semantically similar concepts, leading to improved prior preservation.
-Figure 9: Ablation study on two parameters, i.e., augmentation times $N_{A}$ and augmentation ranks $r$ of the DPA module.
-We report the target CS of erasing Van Gogh and the non-target FID averaged over other four styles (i.e., Picasso, Monet, Paul Gauguin, Caravaggio).
+Once removing the Facilitated Transport module, SPM w/o FT shows poorer prior preservation with rapidly increasing FIDs (highlighted in red in Table 5).
+More qualitative results are shown in Fig.
+9.
+D.4 On Implicit Concept Erasure
+Table 8: Evaluation of implicit concept erasure in erasing nudity on four benchmarks.
+We report the Attack Success Rate (ASR) detected by NudeNet with a threshold of 0.6.
+✓ and $\times$ indicate whether the method can defend against white-box attacks, respectively.
+The best and second-best results are marked in bold and underlined.
+I2P MMA Ring-A-Bell UnlearnDiff Time (s) ↓ MS-COCO White-Box Attack
+CS FID
+MACE Lu et al. (2024a) 0.21 0.04 0.05 0.67 55 (×15) 24.06 52.78
+CPE Lee et al. (2025b) 0.07 0.01 0.00 - 500 (×138) 26.32 48.23 $\times$
+AdvUnlearn Zhang et al. (2024b) 0.04 0.00 0.00 0.21 15860 (×4400) 24.05 57.22
+UCE Gandikota et al. (2024) 0.24 0.38 0.39 0.80 1.2 (×0.33) 26.24 38.60
+RECE Gong et al. (2025) 0.14 0.20 0.18 0.65 1.5 (×0.41) 25.98 40.37
+RACE Kim et al. (2024) 0.23 0.29 0.21 0.47 2910 (×800) 25.54 42.73
+Receler Huang et al. (2024) 0.13 0.07 0.01 - 5560 (×1500) 25.93 40.29 $\times$
+Ours w/o AT 0.20 0.24 0.20 0.75 3.6 (×1) 26.29 37.82
+Ours w/ AT 0.10 0.01 0.00 0.45 4.5 (×1.25) 26.03 39.51
+Evaluation setup.
+We evaluate the erasure efficacy on implicit concepts (e.g., nudity), where the target concept does not explicitly appear in the text prompt.
+We conduct experiments on the Inappropriate Image Prompt (I2P) benchmark Schramowski et al.
+(2023), which consists of various implicit inappropriate prompts involving violence, sexual content, and nudity.
+To evaluate adversarial robustness, we further introduce three adversarial attack benchmarks, including two black-box benchmarks (MMA Yang et al.
+(2024b) and Ring-A-Bell Tsai et al.
+(2023)) and one white-box benchmark (UnlearnDiff Zhang et al.
+(2024c)).
+For concept erasure, we follow the setting in Gong et al.
+(2025) to erase nudity $\to$ ‘ ’.
+During evaluation, we use NudeNet Bedapudi (2019) with a threshold of 0.6 to detect nude content and report the Attack Success Rate (ASR).
+Analysis and discussion.
+In addition to the aforementioned methods, we introduce additional adversarial training-based methods for comparison, including CPE Lee et al.
+(2025b), AdvUnlearn Zhang et al.
+(2024b), RACE Kim et al.
+(2024), and Receler Huang et al.
+(2024).
+These methods enhance robustness against adversarial attacks by explicitly incorporating adversarial training objectives.
+We also adapt our method with adversarial training/editing (denoted as Ours w/ AT) following the setting in RECE Gong et al.
+(2025) to provide a fair comparison.
+As shown in Table 8, we observe that adversarial training-based methods such as CPE and AdvUnlearn achieve strong erasure efficacy but incur extremely high computational costs.
+Editing-based approaches like UCE and RECE are more efficient yet less robust under both black-box and white-box attacks.
+Notably, CPE and Receler rely on additional modules for concept erasure, which makes them particularly vulnerable in the white-box setting since attackers can directly exploit these components to bypass erasure.
+In contrast, our method without adversarial training (Ours w/o AT) already offers a favorable balance between efficiency and prior preservation, and extending it with adversarial training/editing (Ours w/ AT) further improves robustness, reducing ASR across all benchmarks and lowering the white-box UnlearnDiff score from 0.75 to 0.45 while maintaining competitive runtime and prior knowledge preservation.
+D.5 Ablation Studies
+Table 9: Ablation study on the edited parameters.
+Our scheme on only editing the value matrices achieves a superior balance between erasure efficacy (e.g., target CS of 26.29) and prior preservation (e.g., the lowest FIDs across all non-target concepts).
 Ablation Parameters Van Gogh Picasso Monet Paul Gauguin Caravaggio MS-COCO
@@
 Ours $\times$ 26.29 35.86 16.85 24.94 39.75 26.55 20.36
-Table 11: Ablation study on the edited parameters.
-Our scheme on only editing the value matrices achieves a superior balance between erasure efficacy (e.g., target CS of 26.29) and prior preservation (e.g., the lowest FIDs across all non-target concepts).
-Metirc Van Gogh Picasso Monet Paul Gauguin Caravaggio MS-COCO
-CS ↓ FID ↓ FID ↓ FID ↓ FID ↓ CS ↑ FID ↓
-Text Similarity 26.35 36.87 19.69 25.18 41.44 26.52 20.78
-Prior Shift (Ours) 26.29 35.86 16.85 24.94 39.75 26.55 20.36
-Table 12: Ablation study on the importance metrics in IPF.
 Edited parameters.
@@
 Compared to Ablation 2, the editing of key matrices leads to excessive erasure, which is unnecessary in concept erasure.
-Importance metrics in IPF.
-In Sec.
-4.1, we propose Importance-based Prior Filtering (IPF) in Eq.
-5 and evaluate this importance with the metric prior shift $=\|\boldsymbol{\Delta}_{\text{erase}}\boldsymbol{c}\|^{2}$.
-Another intuitive and plausible metric is based on text similarity, e.g., the cosine similarity between each non-target embedding $\boldsymbol{c}_{0}$ and each target concept embedding $\boldsymbol{c}_{1}$, i.e., $\cos(\boldsymbol{c}_{0},\boldsymbol{c}_{1})$.
-Herein, we conduct an additional ablation study in terms of the metric selection in Table 12.
-It can be seen that text similarity can also serve as an effective metric for evaluating importance with improved non-target FID while the prior shift provides better prior preservation.
-This may be because text similarity is implicitly related to importance, while prior shift explicitly reflects the impact of erasure on different concepts from the model updates $\mathbf{\Delta}$.
-Moreover, our method can be directly scaled up to multi-concept erasure scenarios, whereas text similarity calculates $n$ similarities for $n$ target concepts, requiring additional fusion or selection strategies, introducing accumulated errors during fusion or selection.
+((a))
+((b))
+((c))
+Figure 10: Ablation study on hyperparameters. We report target CS of erasing Van Gogh and non-target FID averaged over other four styles (i.e., Picasso, Monet, Paul Gauguin, Caravaggio).
+Filtering scale.
+We ablate the filtering threshold scale in the Influence-based Prior Filtering (IPF) module in Sec.
+4.1 by scaling the impact scores $\mu$ with a factor $\alpha$, which controls the strength of filtering influential priors.
+As shown in Fig.
+10 (a), varying $\alpha$ directly affects the trade-off between erasure efficacy and prior preservation.
+When $\alpha$ is small (i.e., close to 0), more weakly affected priors are included in the retain set, increasing its rank and overly shrinking the null space.
+This leads to worse erasure efficacy (higher CS) and poor preservation (higher FID).
+Conversely, a higher $\alpha$ yields better erasure performance due to fewer retain concepts, but still increases the FID because of non-comprehensive prior coverage.
+The best balance is observed at moderate thresholds (e.g., $\alpha=1$ in our setup), achieving both effective erasure and competitive prior preservation.
+Augmentation times.
+We ablate the augmentation times $N_{A}$ proposed in the Directed Prior Augmentation (DPA) module in Sec.
+4.2, which controls the balance between semantic degradation and retain coverage along with the Influence-based Prior Filtering (IPF) module.
+It can be observed from Fig.
+10 (b) that: (1) As $N_{A}$ increases, the non-target FID exhibits a trend of first decreasing and then increasing.
+This suggests that when $N_{A}$ is small (i.e., $1\to 10$), augmenting existing non-target concepts with semantically similar concepts facilitates a more comprehensive retain coverage, thereby improving prior preservation.
+However, when $N_{A}$ exceeds a certain threshold (i.e., $10\to 20$), further augmentation of non-target concepts leads to narrowing the null-space derivation with semantic degradation, ultimately degrading prior preservation.
+(2) Target CS generally shows a declining trend, indicating that the proposed Prior Knowledge Refinement strategy not only improves prior preservation but also exerts a positive impact on erasure efficacy.
+Augmentation ranks.
+Another hyperparameter to be ablated is the augmentation ranks $r$.
+From Eq.
+7, we introduce the number of the smallest singular values, i.e., augmentation ranks $r$ in deriving $\mathbf{P}_{\text{min}}=\mathbf{U}_{\text{min}}\mathbf{U}_{\text{min}}^{\top}$ with $\mathbf{U}_{\text{min}}={\mathbf{U}}_{\mathbf{W}}[:,-r:]$.
+Mathematically, $r$ represents the directions in which the DPA module can augment in the concept embedding space and constrains the rank of the augmented embeddings to a maximum of $r$.
+As shown in Fig.
+10 (c), as $r$ increases, the non-target FID exhibits an overall upward trend, indicating that introducing more ranks does not benefit prior preservation, as it narrows the null space.
+At the same time, as shown in Table 3, such augmentation by DPA also remains necessary, as it enables more comprehensive coverage of non-target knowledge with semantically similar concepts, leading to improved prior preservation.
 Appendix E Limitation
-While SPEED demonstrates superior prior preservation, its erasure efficacy may not be as strong as some adversarial training/editing-based methods (e.g., RECE [19]), which explicitly optimize for robust concept removal.
-This trade-off arises from SPEED’s emphasis on maintaining non-target knowledge, potentially leading to residual traces of erased concepts in extreme cases.
-However, due to its efficiency and scalability in multi-concept erasure, an interesting direction for future work is to explore simultaneous erasure of adversarial examples.
-Given that null-space constraints inherently minimize the impact on prior knowledge, even with the addition of extra target concepts, SPEED is expected to achieve better prior preservation compared to existing methods while effectively handling adversarial concept erasure.
+Despite the promising results, our work still has limitations.
+SPEED relies on linear null-space projections, which may not fully capture the nonlinear interactions between concepts in large diffusion models.
+In practice, this can lead to imperfect preservation when erasing highly entangled or stylistically subtle concepts.
+In addition, our evaluation mainly covers benchmarks with explicit or moderately implicit concepts; the effectiveness on more abstract (e.g., freedom), compositional (e.g., a blue cat), or culturally nuanced (e.g., Día de los Muertos) concepts remains less explored.
+Finally, although our method scales efficiently to 100 concepts, extending it to even larger-scale or continual erasure settings may require additional mechanisms.

### v4

@@
 (2020); Song et al.
-(2020a; b); Nichol & Dhariwal (2021); Rombach et al.
-(2022); Ho & Salimans (2022) have facilitated significant breakthroughs in generating highly realistic and contextually consistent images simply from textual descriptions Dhariwal & Nichol (2021); Ramesh et al.
+(2020a; b); Nichol and Dhariwal (2021); Rombach et al.
+(2022); Ho and Salimans (2022) have facilitated significant breakthroughs in generating highly realistic and contextually consistent images simply from textual descriptions Dhariwal and Nichol (2021); Ramesh et al.
 (2021); Gal et al.
@@
 (2023); Shan et al.
+(2023); Yan et al.
 (2024), offensive content Schramowski et al.
@@
 2.
-In this light, we introduce Prior Knowledge Refinement, a suite of techniques that strategically and selectively refine the retain set to mitigate the semantic degradation in searching for the null space.
+In this light, we introduce Prior Knowledge Refinement, a suite of techniques designed to strategically and selectively refine the retain set so as to mitigate the semantic degradation in searching for the null space.
 Particularly, we propose Influence-based Prior Filtering (IPF), which first quantifies the influence of concept erasure on each non-target concept.
@@
 Subsequently, to further enhance prior preservation over the resulting retain set, we propose Directed Prior Augmentation (DPA), which expands the retain set with directed, semantically consistent perturbations to improve retain coverage.
-In addition, we incorporate Invariant Equality Constraints (IEC) to preserve specific representations, such as the [SOT] token, that should remain unchanged during editing.
-IEC enforces equality constraints on such invariants to regularize the retaining of essential generation properties.
+In addition, we incorporate Invariant Equality Constraints (IEC), which enforce equality constraints on invariant representations (e.g., [SOT] token), to ensure they remain unchanged during editing.
 We evaluate SPEED on three representative concept erasure tasks, i.e., few-concept, multi-concept, and implicit concept erasure, where it consistently exhibits superior prior preservation across all erasure tasks.
-Overall, our contributions can be summarized as follows:
+Our contributions can be summarized as follows:
   • We propose SPEED, a scalable, precise, and efficient concept erasure method with null-space constrained model editing, capable of erasing 100 concepts in 5 seconds.
@@
 (2024b); Zhao et al.
-(2024a) and additional regularization terms Kumari et al.
-(2023); Lyu et al.
-(2024); Lu et al.
-(2024a).
-In contrast, the editing-based paradigm edits model parameters using a closed-form solution to facilitate efficiency in concept erasure Orgad et al.
+(2024a); Wu et al.
+(2025a) and additional regularization In contrast, the editing-based paradigm edits model parameters using a closed-form solution to facilitate efficiency in concept erasure Orgad et al.
 (2023); Gandikota et al.
@@
 1:
-$\bm{\Delta}_{\text{erase}}=\mathop{\text{arg\ min}}\limits_{\bm{\Delta}}\underbrace{\|(\mathbf{W}+\bm{\Delta})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}}_{e_{1}}+\underbrace{\|\bm{\Delta}\|^{2}}_{\text{regularization}}=\mathbf{W}\left(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\right)\left(\mathbf{I}+\mathbf{C}_{1}\mathbf{C}_{1}\right)^{-1}.$ (5)
+$\bm{\Delta}_{\text{erase}}=\mathop{\text{arg\ min}}\limits_{\bm{\Delta}}\underbrace{\|(\mathbf{W}+\bm{\Delta})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}}_{e_{1}}+\underbrace{\|\bm{\Delta}\|^{2}}_{\text{regularization}}=\mathbf{W}\left(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\right)\left(\mathbf{I}+\mathbf{C}_{1}\mathbf{C}_{1}\right)^{-1},$ (5)
 where $\|\bm{\Delta}\|^{2}$ is introduced for convergence.
@@
 Finally, we combine them to serve as the final refined retain set $\mathbf{R}_{\text{refine}}=\mathbf{R}_{f}\cup\left(\mathbf{R}_{f}\right)^{\text{aug}}_{f}$.
+Figure 4: Qualitative comparison of the few-concept erasure in erasing instances.
+The erased and preserved generations are highlighted with red and green boxes, respectively.
+Our method exhibits consistent prior preservation with less semantic degradation for non-target concepts.
+For example, the middle column better retains details such as Mickey’s hat and button count, and the right column demonstrates more consistent Hello Kitty generations along with three concepts erased.
 4.3 Invariant Equality Constraints (IEC)
@@
 Since the encoding process is masked by causal attention and all prompts are prefixed with the fixed [SOT] token during tokenization, its embedding consistently remains unchanged during T2I process.
-Another invariant is the null-text embedding, as it corresponds to the unconditional generation under the classifier-free guidance Ho & Salimans (2022), which also remains unchanged despite prompt variations.
+Another invariant is the null-text embedding, as it corresponds to the unconditional generation under the classifier-free guidance Ho and Salimans (2022), which also remains unchanged despite prompt variations.
 Given the invariance of these embeddings, we consider additional protection measures to ensure their outputs remain unchanged during concept erasure.
@@
 Since image generation inevitably depends on these invariant embeddings, such constraints inherently preserve prior knowledge.
-Figure 4: Qualitative comparison of the few-concept erasure in erasing instances.
-The erased and preserved generations are highlighted with red and green boxes, respectively.
-Our method exhibits consistent prior preservation with less semantic degradation for non-target concepts.
-For example, the middle column better retains details such as Mickey’s hat and button count, and the right column demonstrates more consistent Hello Kitty generations along with three concepts erased.
 5 Experiments
@@
 (2024), which have achieved SOTA performance across various concept erasure tasks.
-In implementation, we conduct all experiments on SDv1.4 AI (2022) and generate each image using DPM-solver sampler Lu et al.
-(2022) over 20 sampling steps with classifier-free guidance Ho & Salimans (2022) of 7.5.
+In implementation, we conduct all experiments on SDv1.4 Stability AI (2022) and generate each image using DPM-solver sampler Lu et al.
+(2022) over 20 sampling steps with classifier-free guidance Ho and Salimans (2022) of 7.5.
 More implementation details and compared baselines can be found in Appx.
@@
 (2014), where we generate images with the first 1,000 captions, and report CS and FID to measure general knowledge preservation.
-Table 1: Quantitative comparison of the few-concept erasure on instances (left) and artistic styles (right) following Lyu et al.
-(2024).
+Table 1: Quantitative comparison of the few-concept erasure on instances (left) and artistic styles (right), where target concepts are highlighted in pink.
 Arrows indicate the preferred direction for each metric, and the best results are highlighted in bold.
-Our method consistently improves prior preservation for non-target and general concepts from MS-COCO (shaded in pink) while achieving effective concept erasure.
-While our CS is not the lowest for target concpet, Appx.
-D.1 and Fig.
-7 show our method is sufficient for erasure, and lower CS may further compromise prior preservation.
+Our method consistently improves prior preservation on non-target and general concepts from MS-COCO while achieving effective concept erasure.
+Although our CS is not always the lowest for target concepts, Appx.
+D.1 shows our method is sufficient for erasure, and lower CS may further compromise prior preservation.
 Concept Snoopy Mickey Spongebob Pikachu Hello Kitty MS-COCO
@@
 Our method consistently achieves the lowest FIDs across all non-target concepts, demonstrating superior prior preservation with minimal alteration to the original content.
-Moreover, we emphasize that our erasure is sufficiently effective, even without achieving the lowest CS, as shown in Fig.
-4 and Appx.
+Moreover, we emphasize that our erasure is sufficiently effective, even without achieving the lowest CS, as shown in Figs.
+4 and 7.
+On this basis, lower CS values typically indicate “over-erasure” of the target concept, since further reductions in CS after successful erasure usually come at the cost of prior preservation, as detailed in Appx.
 D.1.
-In contrast, lower CS values typically indicate over-erasure, which results in excessive degradation of prior knowledge.
 Notably, with the number of target concepts increasing from 1 to 3, our FID in Pikachu rises from 16.81 to 21.40 (4.59 ↑), while UCE increases from 19.06 to 29.20 (10.14 ↑).
 A similar pattern is observed in Hello Kitty (Our 4.48 ↑ v.s.
-UCE’s 10.29 ↑), showing our robustness in erasing increasing target concepts.
+UCE’s 10.29 ↑), showing our superiority of prior preservation in erasing increasing target concepts.
 5.2 On Multi-Concept Erasure
@@
 Our method can effectively erase up to 100 celebrities simultaneously, achieving low $\text{Acc}_{e}$ (%) and high $\text{Acc}_{r}$ (%) that preserve non-target celebrities with minimal appearance changes.
-This yields the best overall erasure performance $H_{o}$ and competitive runtime (s) on one A100 GPU, successfully erasing 100 concepts in just 5 seconds.
+This yields the best overall erasure performance $H_{o}$ and competitive runtime (s), successfully erasing 100 concepts in just 5 seconds.
 Erase 10 Celebrities MS-COCO Erase 50 Celebrities MS-COCO Erase 100 Celebrities MS-COCO
@@
 (3) Instance erasure on SDv3 Esser et al.
-(2024) from Fig 6 (c): To accommodate the diffusion transformer (DiT) Peebles & Xie (2023) architecture in T2I models, we adapt our method to a DiT-based model, demonstrating a well-balanced trade-off between erasure (1st row) and preservation (2nd row) as well.
+(2024) from Fig 6 (c): To accommodate the diffusion transformer (DiT) Peebles and Xie (2023) architecture in T2I models, we adapt our method to a DiT-based model, demonstrating a well-balanced trade-off between erasure (1st row) and preservation (2nd row) as well.
 Table 3: Ablation study on proposed components in erasing Van Gogh, with the non-target FID averaged over the other four artistic styles from Table 1.
@@
 Reproducibility Statement
-We have made extensive efforts to ensure reproducibility of our work.
+We have made extensive efforts to ensure the reproducibility of our work.
 The proposed method, SPEED, is thoroughly described in the main paper (Secs.
@@
 (2017); Esser et al.
-(2021) and a diffusion model Dhariwal & Nichol (2021); Ho et al.
+(2021) and a diffusion model Dhariwal and Nichol (2021); Ho et al.
 (2020); Sohl-Dickstein et al.
@@
 (2021); Song et al.
-(2020b).
+(2020b); Wang et al.
+(2024c); Li et al.
+(2025).
 The autoencoder encodes an image $\bm{x}$ into a latent representation $\bm{z}=\mathcal{E}(\bm{x})$ and reconstructs it via $\bm{x}\approx\mathcal{D}(\bm{z})$.
@@
 Following the experiment setup from Lu et al.
-(2024a), we introduce a dataset consisting of 200 celebrities, where their portraits generated by SDv1.4 AI (2022) can be recognizable with exceptional accuracy by the GIPHY Celebrity Detector (GCD) Hasty et al.
+(2024a), we introduce a dataset consisting of 200 celebrities, where their portraits generated by SDv1.4 Stability AI (2022) can be recognizable with exceptional accuracy by the GIPHY Celebrity Detector (GCD) Hasty et al.
 (2019).
@@
 (2023); Lu et al.
-(2024a); Gong et al.
+(2024a); Wu et al.
+(2025b); Gong et al.
 (2025); Gandikota et al.
@@
 12 to ensure invertibility.
+Figure 8: Comparison of CS values across different erasure methods.
+We compare the results in erasing Snoopy and Mickey, and highlight the corresponding CS in blue.
+Our method achieves successful concept erasure with moderate CS values.
+In contrast, RECE achieves the lowest CS by enabling more aggressive erasure.
+For example, removing Snoopy into a landscape without a subject, and changing Mickey into a generic person.
+We argue that such over-erasure unnecessarily compromises prior preservation as evidenced by Tables 1 and 2.
+Appendix D Additional Experiments
+D.1 More Demonstrations
+We further provide qualitative visualizations of the erasure results in Fig.7, illustrating the effectiveness of our method in performing precise and targeted concept erasure across diverse scenarios.
+Specifically, we showcase: (a) instance erasure from Table 1 (left); (b) artistic style erasure from Table 1 (right); (c) celebrity erasure from Table 2; and (d) implicit concept erasure (e.g., nudity) from Table 9.
+In all cases, our method successfully removes the intended concept while preserving unrelated content, demonstrating its universal erasure applications.
+Table 5: Human study of erasure efficacy in erasing Snoopy, Mickey, and Spongebob from Table 1.
+Snoopy Mickey Spongebob Average
+RECE 98.93% 98.27% 99.07% 98.76%
+Ours 98.20% 98.40% 98.80% 98.47%
+We also evaluate the CS value before and after concept erasure to assess the erasure efficacy.
+As shown in Fig.
+8, our method achieves successful erasure of specific concepts such as Snoopy and Mickey while maintaining moderate CS values (24.18 and 23.44, respectively).
+This indicates that effective erasure does not require minimizing CS to an extreme.
+In contrast, RECE obtains the lowest CS (19.79 and 18.75), but this is achieved at the cost of overly aggressive erasure.
+For example, transforming Snoopy into an unrecognizable image and replacing Mickey with a generic human figure.
+While such strategies may enhance erasure efficacy, they also risk compromising prior knowledge.
+This trade-off is reflected in higher non-target FIDs, as shown in Tables 1 and 2.
+To further demonstrate that our current erasure is adequate, we additionally conduct a human study.
+For our method and RECE, we randomly sample 50 generated images per method to erase Snoopy, Mickey, and Spongebob.
+We then recruit 30 human participants through Amazon Mechanical Turk Amazon to vote “yes” or “no” on whether the target concept is visually erased or not.
+As shown in Table 5, the overall results (RECE’s 98.76% v.s.
+Our 98.47%) indicate that our method achieves successful erasure on par with RECE from the human perspective.
+D.2 Complete Results on Few-Concept Erasure
+We present complete quantitative comparisons of few-concept erasure, including both CS and FID, in Table 6 and Table 7.
+Our results demonstrate that our method consistently achieves superior prior preservation, as indicated by higher CS and lower FID across the majority of non-target concepts.
 Table 6: Complete quantitative comparison of the few-concept erasure in erasing instances from Table 1 (left).
@@
 Ours 1.81 89.09 93.42 26.47 30.02 3.46 88.48 92.34 26.46 39.23 5.87 85.54 89.63 26.22 44.97
-Appendix D Additional Experiments
-D.1 More Demonstrations
-We further provide qualitative visualizations of the erasure results in Fig.7, illustrating the effectiveness of our method in performing precise and targeted concept erasure across diverse scenarios.
-Specifically, we showcase: (a) instance erasure from Table 1 (left); (b) artistic style erasure from Table 1 (right); (c) celebrity erasure from Table 2; and (d) implicit concept erasure (e.g., nudity) from Table 8.
-In all cases, our method successfully removes the intended concept while preserving unrelated content, demonstrating its universal erasure applications.
-We also evaluate the CLIP score (CS) before and after concept erasure to assess the erasure efficacy.
-As shown in Fig.
-8, our method achieves successful erasure of specific concepts such as Snoopy and Mickey while maintaining moderate CS values (24.18 and 23.44, respectively).
-This indicates that effective erasure does not require minimizing CS to an extreme.
-In contrast, RECE obtains the lowest CS (19.79 and 18.75), but this is achieved at the cost of overly aggressive erasure.
-For example, transforming Snoopy into an unrecognizable image and replacing Mickey with a generic human figure.
-While such strategies may enhance erasure efficacy, they also risk compromising prior knowledge.
-This trade-off is reflected in higher non-target FIDs, as shown in Tables 1 and 2.
-D.2 Complete Results on Few-Concept Erasure
-We present complete quantitative comparisons of few-concept erasure, including both CS and FID, in Table 5 and Table 6.
-Our results demonstrate that our method consistently achieves superior prior preservation, as indicated by higher CS and lower FID across the majority of non-target concepts.
-Figure 8: Comparison of CLIP scores (CS) across different erasure methods.
-We compare the results in erasing Snoopy and Mickey, and highlight the corresponding CS in blue.
-Our method achieves successful concept erasure with moderate CS values.
-In contrast, RECE achieves the lowest CS by enabling more aggressive erasure.
-For example, removing Snoopy to the extent of producing a semantically void image, and changing Mickey into a generic person.
-We argue that such over-erasure unnecessarily compromises prior preservation as evidenced by Tables 1 and 2.
 Figure 9: Qualitative comparison with SPM and SPM w/o FT in erasing single and multiple instances.
@@
 In the latter version, the LoRA scale is set to 1 by default.
-The quantitative comparative results are shown in Tables 5 and 6, where our method consistently achieves the best prior preservation compared to all compared baselines.
+The quantitative comparative results are shown in Tables 6 and 7, where our method consistently achieves the best prior preservation compared to all compared baselines.
 Even equipped with Facilitated Transport (i.e., SPM w/ FT), our method achieves the lowest non-target FID (e.g., on Pikachu and Hello Kitty).
@@
 Ours w/ AT 0.10 0.01 0.00 0.45 4.5 (×1.25) 26.03 39.51
+Table 10: Ablation study on the edited parameters.
+Our scheme on only editing the value matrices achieves a superior balance between erasure efficacy (e.g., target CS of 26.29) and prior preservation (e.g., the lowest FIDs across all non-target concepts).
+Ablation Parameters Van Gogh Picasso Monet Paul Gauguin Caravaggio MS-COCO
+Key Value CS ↓ FID ↓ FID ↓ FID ↓ FID ↓ CS ↑ FID ↓
+1 $\times$ 27.67 42.11 26.09 28.08 52.44 26.55 18.72
+2 26.24 48.41 28.65 33.79 57.23 26.53 23.20
+Ours $\times$ 26.29 35.86 16.85 24.94 39.75 26.55 20.36
 Evaluation setup.
@@
 D.5 Ablation Studies
-Table 9: Ablation study on the edited parameters.
-Our scheme on only editing the value matrices achieves a superior balance between erasure efficacy (e.g., target CS of 26.29) and prior preservation (e.g., the lowest FIDs across all non-target concepts).
-Ablation Parameters Van Gogh Picasso Monet Paul Gauguin Caravaggio MS-COCO
-Key Value CS ↓ FID ↓ FID ↓ FID ↓ FID ↓ CS ↑ FID ↓
-1 $\times$ 27.67 42.11 26.09 28.08 52.44 26.55 18.72
-2 26.24 48.41 28.65 33.79 57.23 26.53 23.20
-Ours $\times$ 26.29 35.86 16.85 24.94 39.75 26.55 20.36
 Edited parameters.
@@
 At the same time, as shown in Table 3, such augmentation by DPA also remains necessary, as it enables more comprehensive coverage of non-target knowledge with semantically similar concepts, leading to improved prior preservation.
-Appendix E Limitation
-Despite the promising results, our work still has limitations.
-SPEED relies on linear null-space projections, which may not fully capture the nonlinear interactions between concepts in large diffusion models.
+Appendix E LLM Usage Statement
+We use large language models (LLMs) as an auxiliary tool during preparing this work.
+The LLM is employed to generate artistic style templates for the artistic style erasure task (see Appx.
+C.1) and to refine the clarity and readability of certain parts of the manuscript, such as polishing grammar, improving fluency, and standardizing terminology.
+In addition, LLMs are occasionally used to suggest alternative phrasings when writing sections like the introduction and related work, but the final narrative, arguments, and presentation choices are made solely by the authors.
+All methodological ideas, theoretical derivations, experiment designs, and analyses are developed independently by the authors without assistance from LLMs.
+We do not rely on LLMs for generating novel research ideas, conducting experiments, interpreting results, or writing technical content.
+The role of LLMs is purely supportive and limited to stylistic refinement and auxiliary text generation, and thus they are not regarded as scientific contributors to this paper.
+Appendix F Limitation
+Despite the promising results, SPEED is designed with linear null-space projections, which may not fully capture the nonlinear interactions between concepts in large diffusion models.
 In practice, this can lead to imperfect preservation when erasing highly entangled or stylistically subtle concepts.
-In addition, our evaluation mainly covers benchmarks with explicit or moderately implicit concepts; the effectiveness on more abstract (e.g., freedom), compositional (e.g., a blue cat), or culturally nuanced (e.g., Día de los Muertos) concepts remains less explored.
-Finally, although our method scales efficiently to 100 concepts, extending it to even larger-scale or continual erasure settings may require additional mechanisms.
+In addition, our evaluation mainly covers benchmarks with explicit or implicit concepts; the effectiveness on more abstract (e.g., freedom), compositional (e.g., a blue cat), or cultural (e.g., Día de los Muertos) concepts remains less explored.
+Finally, although our method scales efficiently to 100 concepts, extending it to even larger-scale or continual erasure may require additional mechanisms.